### v2

# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

Abstract
Seymour’s Second Neighborhood Conjecture asserts that in the square of any oriented graph, there exists a node whose out-degree at least doubles. This paper presents a definitive proof of the conjecture by introducing the GLOVER (Graph Level Order) data structure, which facilitates a systematic partitioning of neighborhoods and an analysis of degree-doubling conditions. By leveraging this structure, we construct a decreasing sequence of subsets that establish a well-ordering of nodes, ensuring that no counterexample can exist. This approach not only confirms the conjecture for all oriented graphs but also provides a novel framework for analyzing degrees and arcs in complex networks. The findings have implications for theoretical graph studies and practical applications in network optimization and algorithm design.
1 Introduction
Seymour’s Second Neighborhood Conjecture, proposed by Paul Seymour in 1990, seeks to identify a node in an oriented graph whose out-degree at least doubles in the graph’s square. The conjecture was first published by Nathaniel Dean and Brenda J. Latka in 1995 [4], where they posed a similar conjecture specific to tournaments. This version of the conjecture was first proven by Fischer [8], with a second proof provided by Havet and Thomasse [11].
Conjecture 1.1.
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}$ represent the first and second neighborhoods of $v$ respectively.
This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [9]) to encompass a broader class of oriented graphs, thereby providing a framework applicable across various minimum degree scenarios.
The Second Neighborhood Conjecture presents a compelling challenge within the context of graph theory, focusing on oriented graphs and the behavior of second neighbors in the graph’s square. Efforts to resolve this conjecture reflect the deep intrigue and mathematical beauty inherent in such problems. Beyond its theoretical importance, the conjecture has potential applications in network theory, algorithm design, and combinatorial optimization—fields where optimizing connections between nodes is crucial.
While Seymour’s Conjecture remains an open problem, various efforts have tackled special cases. For example, Kaneko and Locke [12] demonstrated its truth for graphs with a minimum degree of at most six. Brantner et al. [1] showed that, if the conjecture is false, infinitely many counterexamples must exist. Additionally, Chen et al. [3] provided a lower bound for how much a node’s out-degree increases in the graph’s square, concluding that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Bolter et al [2] where their work confirms that almost all orientations of $G(n,p)$ are satisfy the Seymour Second Neighborhood Conjecture.
The problem of identifying structural properties in oriented graphs has been studied extensively, with particular focus on $m$-free graphs and Seymour vertices. Daamouch’s [7] work has been instrumental in this area, introducing significant results regarding induced subgraphs and their role in determining key vertex properties. This paper builds on these foundational ideas while presenting a novel approach that extends these concepts to interconnected container structures.
In this paper, we will prove Seymour’s Conjecture by assuming it is false and then attempting to construct a counterexample. Specifically, we explore the conditions under which no node’s out-degree doubles in the square of an oriented graph. Our approach focuses initially on minimum out-degree nodes, analyzing how their out-degree could fail to double, and gradually extending this analysis to other nodes based on their distance from the minimum out-degree node.
The Seymour Second Neighborhood Conjecture has long been considered an intractable problem in graph theory, with some experts even doubting its validity. This paper presents a rigorous proof of the conjecture, accompanied by extensive supporting materials, including appendices, detailing definitions, lemmas, examples and an algorithm, as well as an interactive website for visualizing key ideas. These resources aim to make the proof as transparent and accessible as possible, acknowledging the high standards of scrutiny such results rightly deserve.
The paper is organized as follows: Section 2 begins with some common terminology used in graph theory. Section 3 introduces our approach to the problem. In section 4 we look at some small examples of this style of thinking. Section 5 introduces the terminology and identifying the conditions under which their out-degree might fail to double. Section 6 examines the back arcs and their place in oriented graphs. Section 7 looks at the conditions which would cause an interior degree to double, and its importance to the Seymour Second Neighborhood Conjecture. Section 8 further looks at containers under the context of density constraints. In Section 9 considers transitive triangles and Seymour diamonds and show that they are not a hindrance to our algorithm. Section 10 presents this process the full Decreasing Sequence Property algorithm as well as the theorem stating it’s solution to the Conjecture.
2 Graph Theory Terminology
A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $(x,x)$ where $x$ is a node in $G$) and no symmetric arcs, that is no arcs of the form $(x,y)$ and $(y,x)$ where $x$ and $y$ are nodes in $G$. The square of a graph $G$, denoted $G^{2}$, consists of the same node set as $G$ and an arcs exists from node $x$ to node $z$ if there is a node $y$ such that both arcs $(x,y)$ and $(y,z)$ exist in $G$ as well as the arcs already in $G$. Essentially, $G^{2}$ captures two-step connections from $x$ to $z$.
Definition 2.1.
Let $G^{2}=(V,E^{2})$ where $G=(V,E)$ is the original graph, and $E^{2}$ is the set of edges defined as:
$E^{2}={(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E}$
We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$, which are nodes that $v$ has arcs pointing to. Similarly, $N^{++}(v)$ refers to the second neighbors of $v$–that is, nodes that are two steps away from $v$ in $G$.
Let $G=(V,E)$ be a graph and let $S\subseteq V$ be any subset of vertices. The induced subgraph $G[S]$ is the graph whose vertices are the set $S$, and whose edge set consists of all edges in $E$ that have both endpoints in $S$. For any node $v$, we refer to the neighbor-induced subgraph on $N^{+}(v)$, consisting of $v$’s out-neighbors. Let $u,v$ be nodes in $G$. The distance between $dist(u,v)$, is $k$ if the shortest path from $u$ to $v$ has length $k$.
Example 2.1.
[…]
Figure 1: A five-vertex graph illustrating node out-degrees and distances. Out-degrees are labeled near each node, and the dashed rectangle indicates an induced subgraph on vertices A, B, and C. Previously used in [9]
3 Approach and Motivation
We begin our analysis by considering general oriented graphs without making any initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs, such as the existence of an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
From this foundation, we approach the problem from two perspectives. The first is the graph $k$-partition problem [10], where the original oriented graph $G$ is partitioned into $d_{v_{0}}+1$ subgraphs, corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm, which is widely recognized in computer science for simplifying complex problems by tackling smaller, independent subproblems. These partitions enable us to decouple the analysis of interior relationships within each partition from the exterior interactions between partitions, allowing us to address each separately before examining how they integrate within the overall structure. Although the partitions remain connected through certain edges, each can be analyzed independently, allowing us to address subproblems in isolation before combining results to solve the overall problem.
By analyzing each partition individually, we aim to gain localized insights into the graph’s structure and degree dynamics, which we can then synthesize to form a cohesive solution. This divide-and-conquer approach has demonstrated considerable success in algorithms such as mergesort, quicksort, and matrix multiplication. While divide-and-conquer is not commonly applied in graph theory—largely because graphs are typically viewed as unordered collections of vertices and edges—we seek to view this differently by introducing a total ordering on the nodes of $G$. This will allow us to treat the graph in a way that is amenable to divide-and-conquer strategies.
While previous approaches to Seymour’s Second Neighborhood Conjecture often assume the location of a degree-doubling node and focus their analysis around this fixed point, our approach adopts a search-based strategy. Rather than assuming where the doubling will occur, this paper introduces an algorithm that systematically searches the graph, identifying degree-doubling nodes based on structural properties. This method offers a flexible and robust framework that does not depend on preexisting assumptions, allowing for a comprehensive analysis applicable to a broader range of oriented graphs.
This fundamentally shifts the way we approach graph theory problems by introducing a hierarchical structure that is conducive to divide-and-conquer techniques. This reimagining of graph structure allows us to efficiently isolate subproblems, analyze them independently, and then integrate results—offering both computational efficiency and scalability.
One of the key advantages of the structure is its capacity to handle multiple metrics: an outer metric (e.g., distance) for partitioning the graph, and an inner metric (e.g., degree) for analyzing each partition’s internal structure. This dual-metric approach is powerful because it allows for a more nuanced exploration of graph properties, particularly in cases where traditional methods would be insufficient.
The algorithm introduced in this paper systematically searches the graph, leveraging the graph’s structure, total ordering and degree metrics. By not presupposing the location of degree-doubling nodes, the algorithm dynamically identifies these nodes, ensuring a comprehensive exploration of the graph’s structure.
Previous work on Seymour’s Second Neighborhood Conjecture often centers around assuming the position of a degree-doubling node and anchoring the analysis around this fixed point. In contrast, our approach eliminates this assumption, offering a more flexible framework that dynamically adapts to the structure of the graph. This ability to apply a divide-and-conquer strategy to a broader set of graphs marks a significant step forward in graph theory.
4 Initial Lemmas
These first lemmas will seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. For this section, unless otherwise noted, assume that $v_{0}$ is a node with minimum out-degree in our oriented graph $G$. We will also introduce some standard notation from graph theory. The work in this section agrees with what was already done by [12] and is not meant to serve as original content. It does however, provide insight into our reasoning.
We begin our investigation of oriented graphs by searching for a node whose degree could potentially double. While many candidates exist, our goal is to find a node such that, if its degree does not double, we still gain useful information about the structure of the graph. This approach allows us to proceed iteratively, analyzing the graph node by node until we either exhaust all possibilities or discover a counterexample.
Example 4.1.
$v_{0}$$v_{2}$$v_{3}$$v_{1}$$v_{6}$$v_{5}$$v_{4}$
$\Rightarrow H$
Figure 2: Example illustrating how a minimum out-degree node in $G$ has its degree doubled in $G^{2}$ under certain conditions.
Lemma 4.1.
Minimum Out-Degree ¡ 3 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3, then $v_{0}$’s out-degree will at least double in $G^{2}$.
Proof.
We will consider the cases where $d^{+}(v_{0})=0$, $1$, and $2$.
Case 1: Assume that $d^{+}(v_{0})=0$. Then $v_{0}$ has no outgoing edges in $G$. In $G^{2}$, it will still have out-degree 0, and hence its out-degree trivially satisfies the condition of at least doubling.
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. Since $v_{0}$ is a minimum out-degree node, we know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. Because $G$ is an oriented graph and $(v_{0},v_{1})\in G$, it follows that $(v_{1},v_{0})\notin G$. Therefore, $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$, meaning $(v_{1},v_{2})\in G$. This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. Since $v_{0}$ has the minimum out-degree, we know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. As $G$ is an oriented graph, $(v_{1},v_{0})$ and $(v_{2},v_{0})$ are not in $G$. Consider the subgraph $G[v_{1},v_{2}]$, induced on the nodes ${v_{1},v_{2}}$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in $G$, meaning one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. Since $G$ is oriented, $v_{0}$ cannot be one of these nodes. Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4, thus doubling its out-degree.
Therefore, in all cases, $v_{0}$’s out-degree at least doubles in $G^{2}$.
Lemma 4.1 Minimum Out-Degree ¡ 3 gives us a glimpse on the picture of an oriented graphs. It tells us that, even with just the concept of a minimum degree node we are able to impose restrictions on not only that node $v_{0}$, but also on every other node $v\in G$. This later part is true because $v_{0}$ is a minimum degree node and so every other node must have at least the out-degree of $v_{0}$. What this lemma is saying is that cases where $v_{0}$ has very few out-neighbors, its second neighborhood will expand significantly. This aligns with the idea that low out-degree nodes should see a substantial increase in neighbors when considering the second neighborhood.
Lemma 4.2.
Minimum Out-Degree 3 with Neighbor Out-Degree 0 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3, and at least one of its neighbors has out-degree 0 in the induced subgraph of its neighbors, then the out-degree of $v_{0}$ will at least double in $G^{2}$.
Proof.
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$. Consider the subgraph $G[v_{1},v_{2},v_{3}]$, which is the subgraph induced by the set $\{v_{1},v_{2},v_{3}\}$ Now, assume that one of $v_{1},v_{2},v_{3}$, say $v_{1}$, has an out-degree of 0 within the subgraph $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to nodes outside of $\{v_{1},v_{2},v_{3}\}$, as $v_{1}$ has no out-neighbors within the induced subgraph.
Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$, As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in G.
Thus, $v_{0}$’s out-degree will at least double in $G^{2}$.
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors is somewhat ”isolated” within the neighborhood, leading to an increased second neighborhood for $v_{0}$.
Example 4.2.
0
$\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$ H
Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$. Each of these nodes ($v_{1}$, $v_{2}$, and $v_{3}$) connects to the nodes $v_{4}$ and $v_{5}$. The neighbor of $v_{4}$ is $v_{5}$, and the neighbors of $v_{5}$ are $v_{6}$, $v_{7}$, and $v_{8}$. While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$. Thus, in $G^{2}$, the degree of $v_{0}$ will double.
Lemma 4.3.
Minimum Out-Degree 3 with Neighbors 1 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3 and all of its neighbors have an out-degree of 1 in the induced subgraph of its neighbors, then $v_{0}$ or one of its neighbors will have its out-degree at least double in $G^{2}.$
Proof.
Assume $d^{+}(v_{0})=3,$ so $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ Since $v_{0}$ is a minimum out-degree node, we know $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$. Let $G[v_{1},v_{2},v_{3}]$ represent the subgraph induced by $v_{1},v_{2},v_{3}$.
We can reason that if any of $v_{1},$ $v_{2},$ or $v_{3}$ has an out-degree greater than 3, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$, and by Lemma 4.2(Minimum Out-Degree 3 with Neighbor Out-Degree 0) , none of $v_{1}$,$v_{2},$ or $v_{3}$ have out-degree 0 in $G[v_{1},v_{2},v_{3}]$.
Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}].$ If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume they connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree ¡ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one edge between $v_{4}$ and $v_{5}$. One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$. As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$.
Thus, either $v_{0}$ or one of its neighbors will have its out-degree at least double in $G^{2}$. ∎
This lemma underscores a crucial insight: while a node’s out-degree doubles, it is not necessarily the initial minimum out-degree node $v_{0}$ whose degree doubles. Instead, the doubling occurs in one of its neighbors. This distinction is significant because it challenges prior approaches that focus on selecting the ”right” starting node for analysis.
Our approach does not depend on starting with a specific node but instead leverages a search process that inherently reveals the doubling behavior in the graph structure. This framework allows for a more general and robust analysis.
For the remainder of this paper, we will refer to graphs like those in Lemma 4.2 using double arrows to indicate that a node or set of nodes connects to every node in the subsequent subgraph. The graph can be illustrated as follows:
The initial lemmas serve as essential building blocks in our exploration. By establishing these results, we create a solid foundation for the more profound theoretical analysis that follows. Each lemma reinforces the central theme of understanding the interplay between degrees and connectivity in oriented graphs.
5 Theoretical Foundation
Building on the initial lemmas, this section develops the theoretical foundation necessary for addressing the conjecture. We explore key properties such as containers, the Decreasing Sequence Property and the concepts of interior and exterior degrees, which are vital to our argument.
Example 5.1.
[…]
Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from the minimum degree node, with no assumptions on container relationships or sizes.
Definition 5.1.
Given a minimum degree node $v_{0}$, a container $C_{i}$ is the set of nodes that are at distance $i$ from $v_{0}$, along with the arcs between those nodes. More formally, $C_{i}=\{u\in V(G)\text{ .s.t. }dist(v_{0},u)=i\}\text{ and A}(C_{i})=\{(u,v)|u,v\in C_{i}\}$ ,where $V(G)$ represents the vertex set of the graph $G$, and $dist(v_{0},u)$ is the shortest path distance between $v_{0}$ and $u$.
This definition of a container is conceptually related to Daamouch’s use of induced subgraphs in his study of $m$-free graphs [7]. However, the key difference lies in the formation of connected containers (see image), or neighboring containers, which allows for a more dynamic and scalable analysis of graph properties. This approach facilitates the identification of structural patterns across broader graph classes, a task that would otherwise be computationally prohibitive.
As the previous definition aligns with the concept of double arrows, the term ”containers” allows us to explore the structure of oriented graphs and their interactions with minimum degree nodes and distances from these nodes. By introducing containers, we can focus not only on the properties of the nodes relative to a minimum degree node, but also on their interactions within and across containers.
This framework enables us to examine node behavior in three key aspects. First, we can study node behavior relative to a minimum degree node. Containers group nodes based on their distance from a minimum degree node $v_{0}$, making it easier to analyze how these nodes behave and how their properties evolve as we move further from $v_{0}$.
Second, we can investigate internal interactions within containers. Once nodes are grouped into containers, we can study how nodes within the same container influence one another. These internal interactions may reveal patterns related to degree growth, connectivity, and second-degree effects.
Containers are an attempt to bring order to an oriented graph. We mean that in the literal sense of the word, order. For nodes belonging to two different containers, they can now be compared by their distance to a minimum degree node $v_{0}$ which we call in container 0. Since distance is a total ordering, this allows to compare over every other node in $G$
The next few chapters will speak about how these containers communicate with each other and within themselves. No nodes or arcs are removed from the graph to create these containers. This is just a way of thinking about the graph so as to perform our analysis in search of a counterexample.
Finally, we can explore inter-container interactions. By analyzing interactions between nodes from different containers, we gain insight into how graph components expand or contract, how nodes influence one another across distances, and how arc directions impact the convergence of the algorithm or counterexample we aim to develop.
The case where $d^{+}(v_{0})=3$, and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) in particular, is an important one because it provides a glimpse into what a counterexample to the conjecture may look like. We were able to construct a graph $G$ where $v_{0}$’s out-degree does not double in $G^{2}$. Simple graphs of out-degree 3 are the smallest possibilities of cycles in undirected graphs. So if we orient the edges in the same direction we can get a cycle of order 3 among the nodes adjacent to $v_{0}$. If we look at this situation further we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. In this section we will begin exploring the sufficient conditions for other nodes in the graph to not have their degrees double.
To continue along this line of reasoning though we will need to define new terms. First we will need a term for a node’s out-degree not doubling. A node in $G$ whose out-degree does not double in $G^{2}$ is said to have the decreasing sequence property. We have seen that examples exist where $v_{0}$ does not have its out-degree double in $G^{2}$. We will look into what other nodes can have the decreasing sequence property, and what properties must exist in the graph in order for this to take place.
Example 5.2.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
Figure 5: This is an illustration where node $v_{0}$ has the Decreasing Sequence Property (DSP). There are more first neighbors $d$ of $v_{0}$ than second neighbors.
Definition 5.2.
For a node $x\in V(G)$ in an oriented graph $G$, we say that x has the decreasing sequence property if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e.,$|N^{+}(x)|>|N^{++}(x)|$,where $N^{+}(x)$ denotes the set of out-neighbors of $x$, and $N^{++}(x)$ denotes the set of out-neighbors of the nodes in $N^{+}(x)$.
In previous sections, we discussed two fundamental concepts: the minimum degree node and the decreasing sequence property (DSP). A minimum degree node, defined as the node with the fewest connections, serves as a critical starting point for connectivity and flow within the graph. As the graph’s structural anchor, the minimum degree node drives the development of relationships and properties that will emerge in the counterexample construction. It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DSP introduces constraints on connectivity by systematically limiting potential connections beyond nodes that exhibit this property. This concept is akin to a bottleneck: even nodes with multiple connections may still encounter restricted reachability in the graph. A bottleneck in graph theory represents a point that limits flow or connectivity within the network, a role precisely filled by the DSP.
Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively. They jointly reveal scenarios where expected connectivity may falter, highlighting vulnerabilities that could challenge established conjectures. The DSP, in particular, illustrates how exterior nodes in an oriented graph may systematically decrease in number, creating a structure where connectivity flow is constrained despite high out-degrees of certain nodes.
By examining the DSP, we underscore a critical aspect of connectivity, resembling bottlenecks in network theory. Diestel’s Graph Theory [6] addresses the relationship between node degree and connectivity, demonstrating that a node $x$ can maintain many outgoing neighbors while still limiting reachability. This discrepancy signals inefficiencies and sets the stage for potential counterexamples to existing conjectures, thus motivating our investigation of graph properties and their broader implications.
Lemma 5.1.
Total Order Induced by Containers Relative to the Minimum Degree Node
Let $G$ be a graph with a minimum degree node $v_{0}$, where all nodes satisfy the Decreasing Sequence Property (DSP). Construct a series of containers $C_{1},C_{2},\ldots,C_{k}$ relative to $v_{0}$, where each container $C_{i}$ consists of nodes ordered by their adjacency and connectivity properties with respect to $v_{0}$ and preceding containers. This structure imposes a total order on the nodes in $G$, where every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$, establishing a ranking aligned with the DSP.
Proof.
To prove a total order, we need to show that every node in $G$ can be compared to every other node in $G$. Let $u,v$ be two nodes in $G$ such that $u\neq v$. It is assumed that $dist(v_{0},u)$ and $dist(v_{0},v)$ are finite for vertices $u$ and $v$ in the containers $C_{i}$. This assumption implies that both vertices are reachable from $v_{0}$, although it does not necessitate that all vertices are connected. We know that distance is a total ordering, so we know that one of three things are true: either $dist(v_{0},u)<dist(v_{0},v)$, $dist(v_{0},u)=dist(v_{0},v)$, or $dist(v_{0},u)>dist(v_{0},v)$. The situation where $dist(v_{0},u)<dist(v_{0},v)$, implies that the node $u$ is in an closer container to $v_{0}$ than the node $v$. The situation where $dist(v_{0},u)=dist(v_{0},v)$ implies that the nodes $u\text{ and }v$ are in the same container. And the situation $dist(v_{0},u)>dist(v_{0},v)$, implies that the node $u$ is in a further container to $v_{0}$. This shows that any two nodes can be compared and shows that it is a total order. ∎
This is a good start as it imposes order on our graph, but the containers do more than just impose a total order. They essentially allow for grouped comparisons among nodes at the same distance from a reference node, $v_{0}$. This is something that a standard total order does not account for. This ”containerized ordering” enables us to distinguish nodes not only by their distance from $v_{0}$ but also by specific structural or adjacency properties within each distance group.
Unlike a simple total order, which could be any sequence with a single, consistent ordering (like lexicographic ordering), the container structure allows nodes at the same distance to be further organized. This organization is significant because it facilitates comparisons within each distance layer (container) by additional criteria, such as adjacency or neighborhood properties. This layered organization is especially useful in proofs, where we might need to argue about the relative positions of nodes within the same distance from $v_{0}$.
This structure allows us to establish a refined form of comparability among nodes at equal distance that is crucial for analyzing specific properties. For example, in certain proofs, we might need to show that nodes within the same container have certain relationships that would not necessarily follow from a mere total order.
Definition 5.3.
A Level Total Order (LTO) on a set $S$ can be defined as follows:
  • Total Order: For any two elements $u$ and $v$ within the set $S$, exactly one of the following must hold:
    $u$ is less than $v$ (denoted as $u<v$), $u$ is greater than $v$ (denoted as $u>v$), or $u$ is equal to $v$ (denoted as $u=v$).
  • Leveled Structure: The elements of $S$ are partitioned into a leveling of ”containers” or ”levels,” $C_{1},C_{2},,,,,C_{n}$, with each container itself forming a subset of $S$ that maintains a total order. The containers are sequentially ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
  • Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
  • Comparability within Containers: For any two elements $u,v\in Ci$, their order can be determined based on a specific criterion, such as a designated metric or property relevant to the application.
  • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
A leveled total order not only provides an ordering across distance groups but also enables detailed, structure-based comparisons within each group.
By defining both inter- and intra-container comparisons, this ordering allows arguments to leverage container-level properties (like distance) and within-container properties (like adjacency) simultaneously. This flexibility is particularly useful in graph theoretic proofs like the SSNC, where it may be necessary to combine distance-based and structural properties.
Defining a leveled total order gives rigor to proofs, as it offers a structured, multi-layered way to approach ordered sets that are partitioned by one property but require further comparability within each partition.
This formalization could also have broader applications beyond the SSNC. It might apply to any problem where elements need to be organized hierarchically but also require nuanced relationships within each level of the hierarchy—such as layered networks, hierarchical clustering, and complex sorting problems.
Lemma 5.2.
Decreasing Sequence Property Lemma If a node in an oriented graph $G$ does not have the decreasing sequence property, then its out-degree at least doubles in $G^{2}$.
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$, and those that are not neighbors of $x$. In the proofs of Lemma 4.2(Minimum Out-Degree 3 with Neighbor Out-Degree 0), and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor induced subgraph (of $v_{0}$) and in the graph with the neighbor induced subgraph removed. We will now give terms to these concepts.
Definition 5.4.
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
$int(x,y)=N^{+}(x)\cap N^{+}(y)$
,where $N^{+}(z)$ is the set of out-neighbors of $z$. The number of interior neighbors is the interior degree of $y$ with respect to $x$.
We previously introduced the concept of containers, which are crucial for understanding the structure of the graph. Interior neighbors are specifically those nodes for which both endpoints are located within the same container. By examining these interior neighbors, we gain insights into how containers relate to one another and how they function internally.
Definition 5.5.
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The exterior neighbors of $y$ with respect to $x$, denoted $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$:
$ext(x,y)=N^{+}(y)\cap N^{++}(x)$
, where $N^{+}(z)$ is the set of out-neighbors of $z$, and $N^{++}(x)$ is the set of second neighbors of $x$. The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers. This interaction is critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container. However, it is essential to consider other possibilities that may also influence this relationship.
To better understand the distinction between interior and exterior neighbors, consider Example 5.2. In this example, arcs that connect nodes within the same container (or box) represent interior neighbors. In contrast, arcs that extend to nodes in the next container illustrate exterior neighbors.
Definition 5.6.
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
Containers are ordered based on their distance from the chosen minimum degree node. Up to this point, all arcs have either been interior or forward-facing. However, it is important to consider the possibility that some arcs may also direct back toward the chosen minimum degree node. Addressing this potential will help us understand how these back arcs could impact the Seymour Second Neighborhood Conjecture.
Now that we have established the definitions of interior and exterior nodes, we can delve into the decreasing sequence property. This property refers specifically to the sizes of the exterior sets within the oriented graph. As the DSP path algorithm progresses, the exterior nodes — which are not part of the neighborhood of the minimum degree node — will systematically decrease in number. This phenomenon reflects a yin and yang relationship between the interior and exterior nodes: while the interior nodes maintain and potentially increase their interconnections, the exterior nodes steadily diminish, ensuring a balanced growth within the graph.
In this dynamic, the minimum degree node serves as a catalyst for this process, driving the algorithm forward by imposing constraints and guiding the restructuring of connections. In contrast, the decreasing size of the exterior nodes acts as a bottleneck, effectively limiting the expansion of the graph’s out-degrees. This interplay is crucial for the success of the DSP path algorithm, as it maintains the structural integrity of the graph while facilitating necessary transformations.
This relationship between the interior and exterior nodes will be the focal point of our proof throughout the rest of this paper, providing a framework for understanding how the DSP path algorithm operates effectively.
With this definition of interior and exterior nodes now introduced, we are now ready to now introduce the
Definition 5.7.
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Leveled Structure: The elements of $S$ are partitioned into levels of ”containers” or ”rankings,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Theorem 5.1.
Given an oriented graph $G$ and a minimum degree node $v_{0}$, where the nodes of $G$ are partitioned into containers $C_{1},C_{2},\dots,C_{k}$ based on their distance from $v_{0}$ and assuming no back arcs, and the interior arcs in each container and exterior arcs between each container forms a Graph Level Order (GLOVER) under the distance total order.
Proof.
To establish that the structure $\{C_{1},C_{2},\dots,C_{k}\}$ satisfies the GLOVER properties, we verify each condition in the GLOVER definition:
  • Total Order on Elements within $G$: - By partitioning nodes based on their distance from $v_{0}$, we establish a total order on the elements of $G$: for any two nodes $u$ and $v$, their order is determined by their respective containers. Nodes in closer containers (e.g., $C_{i}$) are considered ”less than” nodes in further containers (e.g., $C_{j}$ with $i<j$).
    - Within each container, nodes may be ordered based on additional criteria (such as their individual distances to $v_{0}$ if there are multiple paths). Thus, we achieve a total order among all nodes in $G$.
  • Leveled Structure with Containers: - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$. The elements within each container are in a total order relative to each other, and the containers themselves are ordered by distance from $v_{0}$, establishing a multi-level ranked structure.
  • Non-triviality Requirement: - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node. This fulfills the non-triviality requirement, as we avoid having all singleton containers.
  • Comparability within Containers: - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$ (for example, their shortest-path distance if there are multiple paths). Thus, comparability is ensured within containers.
  • Consistency of Comparisons Across Containers: - Since the graph is oriented and there are no back arcs (i.e., no edges from nodes in $C_{j}$ to nodes in $C_{i}$ for $j>i$), the ordering between containers is respected. This guarantees that if $C_{i}$ is ”less than” $C_{j}$, then every element in $C_{i}$ is ”less than” every element in $C_{j}$, ensuring consistency across containers.
  • Interior and Exterior Neighbors: Nodes $u$ and $v$ of the same distance from the chosen minimum degree node $v_{0}$ would be in the same container and thus be interior neighbors. If these nodes were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance different by 1.
    Each criterion for an GLOVER is satisfied by the structure defined by partitioning nodes based on distance from $v_{0}$. Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a Ranked Total Order under the distance total order, completing the proof.
This lemma aligns with the ranked structure we’re building, where each container $C_{i}$ is distinct in terms of reachability and influence. By enforcing this strict order under the absence of back arcs, we’re creating a clear layering within the graph, which further justifies the ”multi-level total order” approach.
Back arcs, if present, could disrupt the unidirectional progression implied by containers, potentially creating feedback loops. By assuming no back arcs, we ensure that each container maintains its place in the total ordering, which adds robustness to the ranked structure we’re developing.
Containers, as defined in the GLOVER framework, serve not only as organizational units but also as bottlenecks that regulate the flow of connectivity within the graph. By grouping nodes based on their shortest-path distance from the minimum degree node $v_{0}$, containers segment the graph into hierarchical levels. The decreasing sequence property ensures that the size of exterior neighborhoods diminishes as we move outward from $v_{0}$, creating a natural bottleneck effect. This reduction in connectivity simplifies the analysis of degree growth and neighbor interactions while preserving the structural integrity of the graph. Such bottlenecks are critical for maintaining the layered structure of the GLOVER framework, particularly in problems like the Seymour Second Neighborhood Conjecture, where second-level connectivity is a key focus.
Next, we will show the power of this data structure and why is introduced in this paper.
To prove that any oriented graph can be represented in a GLOVER, we’ll need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
Lemma 5.3.
GLOVER Supports Induction: In any oriented graph, the minimum degree node $v_{0}$ within a Glover structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
Proof.
Let $v_{0}$ be the node with the minimum out-degree in the oriented graph $G$. - By definition, for any node $v\in G$, $d^{+}(v_{0})\leq d^{+}(v)$, where $d^{+}(v)$ is the out-degree of node $v$. - This ordering provides a baseline for comparison because all other nodes are either equal to or have greater out-degree than $v_{0}$.
The analysis begins at $v_{0}$, which serves as the foundation of the GLOVER structure. - All outgoing arcs from $v_{0}$ lead to nodes at the next level (or ”container”) $C_{1}$. - By the properties of the GLOVER structure (e.g., no back arcs), the relationships between nodes are strictly directed, ensuring no circular dependencies.
Assume that for any node $v$ in a container $C_{i}$, the following properties hold: - $d^{+}(v)\geq d^{+}(v_{0})$. - All properties derived from $v_{0}$ apply recursively to nodes in $C_{i}$.
Consider any node $u\in C_{i+1}$ that is connected to a node $v\in C_{i}$. - By the inductive hypothesis, $v$ satisfies all conditions derived from $v_{0}$. - Since $u$ receives its inbound arcs from $v$, and $d^{+}(v)\geq d^{+}(v_{0})$, the relationships in $C_{i+1}$ are consistent with the properties of the Glover structure. - Furthermore, no node $u$ in $C_{i+1}$ can violate the inductive assumptions because $v_{0}$ ensures a well-ordering, propagating constraints downward.
Starting from $v_{0}$, the well-ordering propagates through all containers $C_{i}$, ensuring that: - The out-degree of any node is anchored to $v_{0}$. - There is no upward deviation in degree that contradicts the structure’s constraints.
Thus, the GLOVER structure supports induction by leveraging the minimum degree node $v_{0}$ as an anchor, ensuring consistency throughout the graph.
Example 5.3.
[…]
Figure 6: This example illustrates back arcs in a graph. The two blue arcs from $x3$ to $y3$ and $y3$ to $z3$ represent exterior neighbors indicate forward directing arcs. Additionally, a red arc points backwards from $z5$ to $x1$. Although the containers are not outlined, they are implied by the difference in variables for each group.
Lemma 5.4.
Interior Degree Greater than 1 In any oriented graph $G$, represented by a GLOVER with $v_{0}$ as a minimum out-degree node satisfying the decreasing sequence property, In an attempt to prevent a degree doubling node, every node $x$ in the neighboring container $N^{+}(v_{0})$ has an interior degree of at least 1. That is, for $x\in N^{+}(v_{0})$,
$|\text{int}(v_{0},x)|\geq 1.$
Proof.
By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$
where $N^{+}(x)$ denotes the set of first neighbors (out-neighbors) of $x$, and $N^{++}(x)$ denotes the set of second neighbors (out-neighbors of $N^{+}(x)$).
For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $y$, and then from $y$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $y\in N^{+}(v_{0})$.
If $y$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $y$’s neighbors, which are $v_{0}$’s second neighbors. If $y$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $y$’s first neighbors.
To avoid this doubling, $y$ must connect to at least one interior arc. This means that the interior degree of $y$ with respect to $v_{0}$ must satisfy
$|int(v_{0},y)|\geq 1$
This argument holds for any neighbor $y\in N^{+}(v_{0})$, ensuring that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
Thus, for all $x\in N^{+}(v_{0})$,$|int(v_{0},x)|\geq 1$, completing the proof. ∎
Lemma 5.4 (Interior Degree Greater than 1) is important because it constrains the out-neighbor structure of the graph when the minimum out-degree $v_{0}$ exhibits the decreasing sequence property. It guarantees that each out-neighbor of $v_{0}$ must have at least one other out-neighbor within $N^{+}(v_{0})$ , which helps to prevent the out-degree of nodes from being too heavily concentrated in the second neighborhood.
The GLOVER structure partitions the neighbors of nodes into interior and exterior categories in a systematic way. This partitioning provides a powerful framework for analyzing oriented graphs, particularly in controlling and understanding degree properties.
A notable consequence of this structured representation is the ability to guarantee that for any oriented graph $G$ represented as a GLOVER, the degree of a chosen minimum out-degree node $v_{0}$ will not double due to cycles in container $C_{1}$. More formally:
Proposition 5.1.
Cycle in Container 1 If $G$ is an oriented graph represented as a GLOVER with $v_{0}$ as the chosen minimum degree node of degree $\delta$, and if there exists a cycle in the first container $C_{1}$ of length $\delta$, then the degree of $v_{0}$ will not double.
Example 5.4.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
Lemma 5.5.
Exterior Less than Degree In any oriented graph $G$, represented by a GLOVER In an attempt to prevent a degree doubling node, if a node $x_{i}$ is in container $C_{i}$ has the decreasing sequence property, In an attempt to prevent a degree doubling node, then for all $y\in N^{+}(x)\subseteq C_{i+1}$, $|ext(x,y)|<d^{+}(x)$.
Proof.
Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|ext(x,y)|\geq d^{+}(x)$. By definition, $ext(x,y)$ consists of the set of out-neighbors of $y$, specifically:
$ext(x,y)=\{z\in N^{++}(x)\mid z\text{ is an out-neighbor of }y\}.$
Thus, $|ext(x,y)|\subseteq N^{++}(x)$, where $N^{++}(x)$ denotes the set of second out-neighbors of $x$. If $|ext(x,y)|\geq d^{+}(x)$, then:
$|N^{++}(x)|\geq|N^{+}(x)|\geq d^{+}(x),$
which suggests that the out-degree of $x$ would at least double in the graph $G^{2}$ (the second power of the graph, where edges represent paths of length 2). This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property, where the degree of nodes cannot increase as we move through the containers.
Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|ext(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
This proof directly shows that a contradiction arises if the exterior degree exceeds the out-degree, which would violate the decreasing sequence property. The argument about the doubling of the out-degree in $G^{2}$ is key in demonstrating that the structure defined by the GLOVER representation prevents such a situation, thereby ensuring that $|ext(x,y)|<d^{+}(x)$.
The results presented in 5.1(Cycle in Container 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1, which demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles. While our work is similar to this result, we will extend the framework by leveraging container-based structures to provide a polynomial-time solution, highlighting the interplay between neighboring containers and their shared arcs.
Lemma 5.6.
[…]
Then, for all $z\in ext(x,y)$, we have $|ext(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
Proof.
We begin by recalling the conditions for a node to have the decreasing sequence property (DSP). A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
Now, assume the conditions of the lemma hold: 1. The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increases significantly in $G[N^{+}(x)]$. 2. For all $z\in ext(x,y)$, $ext(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
Now, suppose for contradiction that there exists $z\in ext(x,y)$ such that $|ext(y,z)|\geq|ext(x,y)|$. This would mean that the exterior degree of $y$ is at least as large as the exterior degree of $x$. However, this contradicts the assumptions about the DSP of $y$, which requires $y$ to have fewer second neighbors than first neighbors.
From $|int(x,y)|$ doubling, $y$’s influence in $G[N^{+}(x)]$ grows, yet its exterior degree should decrease due to the DSP. The absence of back arcs prevents any increase in $y$’s exterior degree from its second neighbors.
Thus, the assumption $|ext(y,z)|\ \geq|\text{ext}(x,y)|$ leads to a contradiction,
confirming that $|\text{ext}(y,z)|<|\text{ext}(x,y)|$ for all $z\in\text{ext}(x,y)$.
Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$. ∎
The central concept of this proof revolves around the relationship between decreasing exteriors and back arcs. This relationship is not merely general; it specifically addresses the concern of potential double counting involving second neighbors that arise from back arcs through $y$.
The additional condition - ensuring that a neighbor of $y$ does not also serve as a first neighbor of $x$ is vital. This condition helps to prevent the overlap in counting, particularly in situations where $z$, a second neighbor of $y$, could inadvertently be counted as a first neighbor of $x$.
Given that containers are structured by their distance from the chosen minimum degree node, it follows that if a node $z$ is a second neighbor of $y$ and also a first neighbor of $x$, this necessitates the presence of a back arc from $z$ to a neighbor of $x$. Therefore, the lemma’s validity heavily relies on the configurations of back arcs in our construction, particularly in the context of a potential counterexample.
In summary, the interplay between back arcs and the classification of neighbors is crucial in Lemma 5.6 (Weak Decreasing Exteriors Lemma). By maintaining the distinction between first and second neighbors through careful conditions, we can mitigate the risk of double counting and strengthen the argument presented in the proof.
This next lemma looks into the Weak Decreasing Exteriors Lemma and considers the assumption of back arcs. There is a particular scenario that needs to be evaluated where a second neighbor of $y$ and first neighbor of $z$ can also be a first neighbor of x. The above lemma, called 5.6(Weak Decreasing Exteriors Lemma), has an explicit assumption that gets rid of this. In the next chapter we will do a lot of work to make sure that these back arcs are no longer a problem. However we will consider a second version of this lemma that gets rid of the back arcs assumption.
Example 5.5.
0
$\Rightarrow$ 111122223333444455556666 $\Rightarrow$ 7777888899991010101011111111 $\Rightarrow$ H
Figure 7: Nodes in each container connect to all nodes in the subsequent container. Here, the minimum out-degree node has out-degree 6, with connectivity patterns shown to demonstrate degree doubling in $G^{2}$.
Table 1: Intersection Results for Triplets of Nodes in 5.5
Triplet(A, B, C) int(A, B) ext(B, C) Intersection Result
[…]
The theoretical foundation established in this section underpins the subsequent arguments and proofs. By thoroughly examining the properties of oriented graphs, we prepare to tackle the complexities of back arcs and their implications for the overall graph structure.
6 Back Arcs
In this section, we analyze the role of back arcs in oriented graphs and their effects on graph properties. Understanding back arcs is crucial for assessing how they influence the interior and exterior degrees, as well as the overall structure of the graph.
It is important to restate that back arcs and exterior neighbors are distinct concepts. Back arcs are not a subset of exterior neighbors. While back arcs connect to any previous container, exterior neighbors are limited to connecting with the next container. Additionally, exterior neighbors do not represent all possible external neighbors. Rather, an exterior neighbor ext(x, y) specifically refers to nodes at a distance of two from $x$ that are also neighbors of $y$.
If we look at Example 4.2 again we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$. Similarly we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
One of the things we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. This next lemma formalizes that concept.
Example 6.1.
[…]
Figure 8: Graph representation of back arcs (shown in red) and forward-directed arcs (in blue) between containers, illustrating how these arcs impact connectivity patterns in an oriented graph.
Lemma 6.1.
(Regular Interior Degrees) Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node, with minimum degree node $v_{0}$ partitioned by containers $C_{i}$. Suppose that two nodes $u$ and $v$ in a container $C_{i}$ have a common parent $v_{p}$ in the preceding container $C_{i-1}$, and that $v_{p}$’s interior degree doubles. If the interior degrees of $u$ and $v$ differ by at least 1, then the exterior degree of $v_{p}$ in container $C_{i-1}$ will double as well.
Proof.
Assume $u$ and $v$ are two nodes in container $C_{i}$, both with a common parent $v_{p}$ in container $C_{i-1}$, and let $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$, where $\text{int}(v_{p},x)$ denotes the interior degree of $x$ relative to $v_{p}$’s container $C_{i-1}$.
By the Decreasing Sequence Property (DSP), if the interior degree of $v$ is higher than $u$, then $v$ will require fewer exterior connections to maintain the required degree restrictions in the graph. Meanwhile, $u$, with fewer interior connections, will require more exterior connections to compensate.
Now, the imbalance in interior degrees means that $u$’s exterior connections are effectively increasing the degree of $v_{p}$. This is because $u$ must connect to more external nodes to maintain its degree, leading to an increase in the number of exterior neighbors of $v_{p}$.
Consequently, the exterior degree of $v_{p}$ must double to accommodate the increased number of connections needed by $u$ and $v$ in the graph, violating the degree constraints intended by the partitioning into containers.
Thus, the differing interior degrees of $u$ and $v$ in container $C_{i}$, combined with the effects of the DSP, will cause the exterior degree of $v_{p}$ in $C_{i-1}$ to double. This leads to a contradiction, so the claim holds. ∎
Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples. Specifically, it shows that all the out-neighbors of interior arcs within a container must have the same degree, which is known as out-degree regularity. If, additionally, the nodes in this container have equal in-degrees, then the container achieves total regularity. It is important to note that this regularity emerges as a counterexample strictly due to the interaction between the Decreasing Sequence Property (DSP) and the minimum degree node $v_{0}$.
In Example 5.5 and Table 1, we examined a minimum degree node $v_{0}$ with degree 6 in a regular graph. The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth. This outward growth reflects the structure of the graph, where nodes are grouped into containers based on their distance from $v_{0}$.
A back arc is defined as an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$, effectively pointing against the distance-based order. If such a back arc were introduced, it would imply that the source node $y$ gains new neighbors at a lower level, potentially doubling its degree. This contradicts the hierarchical organization of containers, where each level is strictly defined by shortest-path distances from $v_{0}$.
Challenge:
To better understand why back arcs are not possible in this graph, consider the following questions:
  • Can you construct a back arc in this graph that respects the container construction rules and distance-based ordering from $v_{0}$?
  • If a back arc is introduced, does it lead to contradictions or disrupt the hierarchical structure of containers?
Explanation:
By the construction of the GLOVER framework, all nodes in $C_{i}$ are farther from $v_{0}$ than those in $C_{j}$ for $j<i$. A back arc $(y,z)$ would imply that $z$ is closer to $v_{0}$ than $y$, violating this ordering.
Additionally, introducing a back arc disrupts the unidirectional flow of the graph by creating a feedback loop. This contradicts the fundamental property of containers: that each level represents a discrete, non-overlapping set of nodes ordered by distance from $v_{0}$. Without back arcs, the GLOVER structure preserves a clear layering, enabling rigorous analysis of degree growth and neighbor interactions.
Lemma 6.2.
No Back Arc to Early Container Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $ext(u_{i},v_{k})$ is at least the degree of $v_{k}$. Furthermore, if the interior degree of $u_{i}$ doubles, this will cause the total degree of $u_{i}$ to double as well.
Proof.
Assume that $u_{i}\in C_{i}$ has its first back arc to $v_{k}\in C_{k}$ where $k<i-1$. By definition of exterior neighbors:
$|ext(u_{i},v_{k})|=|N^{+}(v_{k})\cap N^{++}(u_{i})|$
where $N^{+}(v_{k})$ is the set of immediate neighbors of $v_{k}$, and $N^{++}(u_{i})$ is the set of nodes at distance two from $u_{i}$.
Since $N^{+}(v_{k})$ includes all direct neighbors of $v_{k}$, it follows that:
$|N^{+}(v_{k})|=deg(v_{k}).$
By the structure of GLOVER and the absence of back arcs from $u_{i}$ to $v_{0},\dots,v_{k-1}$, the size of $ext(u_{i},v_{k})$ satisfies:
$|ext(u_{i},v_{k})|\leq|N^{+}(v_{k})|=deg(v_{k}).$
Now, consider the relationship between interior and total degree. If the interior degree of $u_{i}$ doubles, the total degree of $u_{i}$ must also double. This is because the exterior neighbors are tightly constrained by the back arc and are not impacted by the increase in interior connections. Thus, any increase in the interior degree directly translates into a proportional increase in the total degree.
Finally, by the Weak Decreasing Exteriors Lemma, the size of $ext(u_{i},v_{k})$ decreases as we move further in the sequence of containers. This ensures that the total degree of $u_{i}$ remains bounded and behaves as expected under the given conditions, completing the proof. ∎
Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container structured graph. It has effectively reduced this case of back arcs to one of interior degree doubling or back arcs mapping to the previous container. This is important because the previous section introduced the GLOVER framework as a total order of containers. Back arcs can disrupt this order. What we see next is a special case of how back arcs could disrput this order.
Lemma 6.3.
Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
1. Suppose $x\in C_{i}$ and $y\in C_{i+1}$, both satisfying the Decreasing Sequence Property (DSP). 2. The interior degree of $y$ with respect to $x$ doubles in $G[N^{+}(x)]^{2}$. 3. There exists a $w\in\text{ext}(x,y)$ such that the $int(y,w)$ doubles.
Then, for all $z\in\text{ext}(x,y)$, we have:
$|\text{ext}(y,z)|<|\text{ext}(x,y)|$
First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
Consequently, the arc $w,z$ is an early back arc as we saw in 6.2(No Back Arc to Early Container). These arcs can lead to double counting. An out-neighbor of $w$ might be an out-neighbor of $x$, and a second out-neighbor of $y$. This implies that under the assumption that $w$’s interior degree double’s, $w$’s overall degree will double through that early back arc.
The framework under 6.2 (No Back Arc to Early Container) generalizes the solution to handle all potential back arc scenarios in the graph. Thus, we have strengthened the overall structure, making the algorithm robust against these issues and ensuring that the exterior degree conditions are preserved across all containers.
Lemma 6.4.
First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph represented by a GLOVER with minimum degree $d$ in an attempt to prevent a degree doubling node, and let $v_{0}$ be a node of minimum degree. Suppose $G$ is an oriented graph, where arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will violate the structure of the graph.
Proof.
Assume that the oriented graph $G$ is partitioned into containers $C_{1},C_{2},\dots,C_{i}$ such that nodes in each container $C_{j}$ (for $j<i$) only have either arcs to fellow nodes in $C_{j}$, or they map forward to nodes in $C_{j+1}$, with no back arcs present between containers up to $C_{i}$.
Now, let the first back arc appear from a node $u_{i}\in C_{i}$ to a node $v_{i-1}\in C_{i-1}$. In this oriented graph, by definition, all arcs are supposed to point in the forward direction—from container $C_{j}$ to container $C_{j+1}$. By Lemma 6.1 (Regular Interior Degrees), we know that all nodes inside container $C_{i-1}$ have the same interior degree. Because of their matching interior degrees, they will also share the same exterior degrees. Then since the node $v_{i-1}$ is in $C_{i-i}$, $v_{i-1}$ also maps to these exterior nodes. If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist{v_{0},u_{i}}-i$ so $u_{i}$ cannot belong to container $C_{i}$.
Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ leads to a fundamental contradiction in the oriented nature of the graph, and no back arcs can exist prior to $C_{i}$. ∎
The concept of a first back arc inherently assumes that no back arcs exist up to a certain point. Lemma 6.2 (No Back Arc to Early Container) supports this assumption, relying on the doubling of nodes’ interior degrees, a condition not necessary for Lemma 6.4. The strength of this lemma lies in its capacity to identify the first back arc; if such an arc cannot be established, it reinforces the graph’s structural integrity. This lemma is founded on our ability to well-order the graph using containers and the Decreasing Sequence Property (DSP). With this ordering, we highlighted regularity within these containers through Lemma 6.1 (Regular Interior Degrees), which proved essential in showing that these back arcs would lead to a node’s degree doubling if a node’s interior degree was already doubling.
Lemma 6.5.
No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$,$C_{2}$, $\dots$,$C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
Proof.
This was proved in the Lemmas 6.2 (No Back Arc to Early Container and 6.4 (First Back Arc Between Consecutive Containers) . ∎
[…]
Figure 9: Illustration of the proof of 7.2 (Interior Degrees Double). Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.
[…]
The exploration of back arcs reveals their significant impact on the dynamics of oriented graphs. By removing them, we can better navigate the structures of oriented graphs and in particular it paves the way to investigate interior degree dynamics.
7 Interior Degrees Doubling
This section focuses on the important aspect of interior degree doubling within oriented graphs. We investigate the conditions under which this occurs and the implications for the overall degree distribution of nodes.
Lemma 7.1.
Indistinguishable Regular Interior Nodes In a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER in an attempt to prevent a degree doubling node. If a node $v_{p}$’s interior degree doubles then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree double.
Proof.
Let $C_{i}$ be a regular container of size $n$, and let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$. The interior degree of a node $v_{i}$ in container is defined as the number of arcs shared $v_{i}$ between $v_{i}$ and a neighbor or parent node. In a regular container, all nodes are indistinguishable in terms of their degree. They also have the same parents because of the DSP. Any automorhism we construct including a random automorphism $\phi$ of the container $C_{i}$ will preserve the degree structures. This means any node will have its interior degree double. ∎
The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others. Initially, we accounted for nodes where interior degrees had doubled and ensured consistency when exterior degrees doubled as well. This distinction was critical for our algorithm up to this point. However, with regular containers, interior nodes are now indistinguishable from one another. This means we can swap a node $v_{i}$ with an exterior node $v_{e}$ if their degrees have doubled, ensuring that we have a node whose overall degree has doubled. This flexibility simplifies the process and ensures that the algorithm maintains consistency.
Lemma 7.2.
Interior Degrees Double Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have its interior degree doubled.
Proof.
We prove this by strong induction on the size of the graph.
Base Case: In Lemmas 4.1 (Minimum Out-Degree ¡ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4.3 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3 there exists at least one node in each container with a doubled interior degree. These cases form our base, supporting the induction.
Inductive Hypothesis: Assume that for any graph of size $n^{\prime}<n$, each container $C_{i}$ contains at least one node whose interior degree doubles, thus satisfying the DSP by proportionally reducing exterior degrees.
Inductive Step: Consider a graph of size $n$ where we need to show that at least one node in each container $C_{i}$ has its interior degree doubled.
Suppose we construct $G$ such that no node at distance $i$ from $v_{0}$ initially has a doubled interior degree in $C_{i}$. Define a sequence of subgraphs $H_{0}$,$H_{1}$,…,$H_{k}$ for $C_{i}$ as follows:
$H_{0}=\{u_{i}\}$
$H_{1}=N^{+}(u_{i})$
with nodes in $H_{1}$ having doubled interior degrees by the I.H., as applied to each subgraph. By the Induction Hypothesis (I.H.),there exists a node $y_{1}\in H_{1}$ such that the interior degree of $y_{1}$ with respect to $u_{i}$ doubles in $H_{1}$. We can now define $H_{2}=ext(u_{i},y_{1})$. Likewise, by the I.H., there will be a node $y_{2}inH_{2}$ whose interior degree doubles in $H_{2}$. We can define $H_{3}=ext(y_{1},y_{2})$. Continuing this process, we can define $H_{j}=ext(y_{j-2},y_{j-1})$, where $y_{j-1}$ is a node whose interior degree doubles in $H_{j-1}$.
By the Weak Exterior Size Decreasing lemma, we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers. By the DSP, an increase in $int(x,y)$ requires a proportional reduction in $ext(y,z)$, preventing exterior degrees from exceeding interior degrees. This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$, as required by DSP.
The strictly decreasing nature of this sequence implies that it must terminate when $H_{i}\leq 2$. At this point, the remaining nodes in $H_{i}$ either have no further exterior neighbors or their degrees lie entirely within the interior set.
Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$. Thus, the parent of $y^{*}$, denoted $y^{*}-1$, in $G$ achieves a doubled degree within $C_{i}$.
This doubling process terminates as DSP constraints prevent exterior degrees from surpassing interior degrees without contradicting the DSP.
Therefore, for a graph of size $n$, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step. ∎
Considering Lemma 7.2 (Interior Degree Double), we observe that it represents a specific instance of the Seymour Second Neighborhood Conjecture. Here, each container functions as its own self-contained oriented subgraph, where we pose the same core question: is there a node whose degree doubles? The affirmative answer we obtain, even when constructing a counterexample, suggests promising progress toward resolving the overall conjecture.
Furthermore, we can expand upon the process illustrated in Lemma 7.2 (Interior Degree Double)). Within the proof, we constructed a monotonically decreasing sequence of subsets, which cannot continue indefinitely. This sequence reflects the essence of the Decreasing Sequence Property (DSP), where we observe that as interior nodes increase, the exterior nodes necessarily decrease. This requirement is what we meant by “decreasing” in DSP: the exterior set gradually diminishes as nodes transition to the interior.
This dynamic can be interpreted as a form of equilibrium, reminiscent of the principle of yin and yang. Each node integrated into the interior necessitates a corresponding reduction in the exterior set, thereby preserving a balance. This interplay ensures that interior growth does not occur in isolation; instead, it aligns directly with constraints on the exterior, maintaining the graph’s structural integrity.
Such interaction is essential for upholding the properties defined by DSP. It underscores the inherent relationship between interior and exterior neighbors, reinforcing the notion that the expansion of one aspect imposes natural limitations on the other.
In summary, examining the doubling of interior degrees within this context provides critical insights into the degree distribution in oriented graphs. This understanding informs the proofs that will be discussed in subsequent sections, enhancing our grasp of both local and global properties of the graph.
8 Container Density
Here, we delve into the concept of container density and its relevance to the properties of oriented graphs. We explore how regular exterior degrees interact with the overall density of containers and the implications for graph structure.
Having proved that the interior degrees double within a container and that is linked to container size decreasing, we are ready to tackle the overall Seymour Second Neighborhood Conjecture. First we need to establish some fundamental truths about containers and oriented graphs in these counterexamples to set up the rest of the paper and things that must be true.
Lemma 8.1.
Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in $G$ has the Decreasing Sequence Property (DSP). Then the regular interior degrees will result in regular exterior degrees in $G$.
Proof.
Let $x_{i}$ and $x_{j}$ be two nodes at the same distance $i$ from the minimum degree node $v_{0}$ in the oriented graph $G$, both contained in the same container $C_{i}$.
Assume, for contradiction, that the sizes of their exterior sets differ by at least one. Without loss of generality, let there exist a node $y$ such that $|ext(x_{i},y)|=|ext(x_{j},y)|+1$. This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
Since the size of the next container $C_{i+1}$ is determined by the exterior degrees of nodes in $C_{i}$, this irregularity in the exterior degrees would force the size of $C_{i+1}$ to differ, contradicting the assumption that the containers are of regular size.
By Lemmas 6.1 (Regular Interior Degrees) and 7.2 (Interior Degree Double), we know that if the interior degrees within a container are regular, the exterior degrees must also follow this regularity. Specifically, if there is an irregularity, any parent node of $x_{i}$ and $x_{j}$ will have its exterior degree double, which would in turn double their total degree.
This doubling of the parents’ total degree would violate the Decreasing Sequence Property (DSP).
Therefore, the assumption that $|ext(x_{i},y)|$ and $|ext(x_{j},y)|$ differ is false. Hence, the exterior degrees in $G$ must be regular if the interior degrees are regular. ∎
What we saw in Lemma 6.1 (Regular Interior Degrees) that the regularity of interior degrees was something we were able to prove while we were trying to prove that back arcs did not exist. This both made sense and was very helpful in that environment. We see here that the regularity of interior degrees is not restricted to the container itself, but also impacts the exterior arcs to the next container - forcing them to be regular as well.
Lemma 8.2.
Regular Graph All Nodes Have Equal Our-Degree Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node, with $v_{0}$ as a minimum degree node. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $d_{v_{0}}=d^{+}(v_{0})$.
Proof.
We will prove this lemma by induction on the number of nodes in the graph $G$.
For the base case, consider the node $v_{0}$, which, by definition, has the minimum degree $d$, so $d^{+}(v_{0})=d_{v_{0}}$.
Assume that for some node $v_{k}$ in $G$, the degree of $v_{k}$ satisfies $d^{+}(v_{k})=d_{v_{0}}$.
Now consider an arbitrary node $v_{k+1}$ in $G$. Suppose, for contradiction, that $d^{+}(v_{k+1})>d_{v_{0}}$.
Consider a neighbor $v_{n}$ of $v_{k+1}$. By the inductive hypothesis, the degree of $v_{n}$ is exactly $d_{v_{0}}$. Since $v_{k+1}$ has more than $d_{v_{0}}$ neighbors, the neighbors of $v_{k+1}$ are second neighbors of $v_{n}$ in the graph $G^{2}$. Therefore, $v_{n}$ would have more than $d_{v_{0}}$ second neighbors in $G^{2}$.
This implies that $v_{n}$’s degree has effectively doubled, which contradicts the DSP, which requires that $v_{n}$ must have degree $d_{v_{0}}$. Hence, the assumption that $d^{+}(v_{k+1})>d_{v_{0}}$ leads to a contradiction.
Thus, we must conclude that:
$d^{+}(v_{k+1})=d_{v_{0}}.$
By the principle of mathematical induction, it follows that for every node $v$ in $G$, the degree of $v$ is equal to $d_{v_{0}}$. ∎
We are incrementally refining the properties that must hold for our counterexamples. The two most crucial aspects are the Decreasing Sequence Property and the identification of the minimum degree node. However, additional constraints, such as the overall graph regularity established by Lemma 8.2 (Regular Graph), help simplify future arguments. This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
Lemma 8.3.
Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER In an attempt to prevent a degree doubling node, with no back arcs, and let $v_{0}$ be the chosen minimum degree node in $G$. For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1)$
for some $z\in N^{+}(x_{i})$ in order to keep any node’s degree from doubling in the oriented graph.
Proof.
We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$.
Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
$\{y\in\text{ext}(x_{1},z)\text{ s.t. }dist(v_{0},y)=2\}$
is bounded by $d^{+}(v_{0})-1$, since each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2, as there are no back arcs to $v_{0}$.
Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1).$
Inductive Step: We need to show that this bound holds for nodes at distance $i+1$ from $v_{0}$. Let $x_{i+1}$ be a node in container $C_{i+1}$ relative to $v_{0}$. Suppose, for contradiction, that
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|>d^{+}(v_{0})-i.$
This would imply that the exterior neighbors of $x_{i+1}$ exceed the allowed bound, violating the decreasing sequence property (DSP) conditions of $x_{i}$. By Lemmas 7.2 (Interior Degrees Double), 6.1(Regular Interior Degree), and 7.1 (Indistinguishable Regular Interior Nodes), this would cause the interior degree of $x_{i}$ to double, thereby forcing an overall doubling of the degree for $x_{i}$.
This contradiction implies that our assumption must be false. Therefore, we conclude:
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|\leq d^{+}(v_{0})-i.$
This lemma is a more extended form of Lemma 5.6 (Weak Decreasing Exteriors Lemma), but without the added dependencies. It gives us a way of calculating the maximum size of a container based only in its distance from the minimum degree node and the minimum degree node’s degree.
The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” is named. Several examples and observations brought this to a conjecture, which brought it here. The question now becomes is there any way we can strengthen this result?
We demonstrated that the Decreasing Exteriors Lemma relies on two key factors: the elimination of back arcs and the doubling of interior degrees within each container. Furthermore, we established that the elimination of back arcs is directly dependent on the interior degrees doubling within those containers.
By proving the previous lemma, we not only confirm the doubling of interior degrees but also conclude two additional results: (1) back arcs are entirely eliminated, and (2) the exterior arcs exhibit the desired decreasing pattern.
Lemma 8.4.
Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER, In an attempt to prevent a degree doubling node, and let $v_{0}$ be a minimum degree node in $G$, where every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided $(v_{0},x)\in G$.
Proof.
Let $x$ be a node at distance $i$ from $v_{0}$. By Lemma 8.3 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$, $|ext(v_{i-1},x)|$, is bounded by $d^{+}(v_{0})-i$. This places a limit on the number of nodes $x$ can map to at distance $i+1$ from $v_{0}$.
Now, since $x$ has at least the same out-degree as $v_{0}$ (by the assumption of DSP), $x$ must map to at least $i$ other nodes. Given that back arcs are not allowed (as shown previously), $x$ cannot map to nodes at earlier distances without violating the structure of the graph.
Therefore, the only remaining possibility is that $x$ maps to interior nodes. Since $x$’s degree must account for at least $i$ mappings, and these must be to interior nodes, we conclude that $x$ must map to at least $i$ interior nodes.
Thus, $|int(v_{i-1},x)|\geq i$, as required. ∎
Proof.
Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a node with the minimum degree in $G$, and assume that every node in $G$ satisfies the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and let $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$.
Our goal is to show that the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided that $(v_{0},x)\in G$ (i.e., $x$ is reachable from $v_{0}$).
By Lemma 8.3 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$ with respect to $x$, $|ext(v_{i-1},x)|$, is bounded above by:
$|ext(v_{i-1},x)|\leq d^{+}(v_{0})-i$
where $d^{+}(v_{0})$ is the out-degree of $v_{0}$. This gives a limit on the number of nodes that $x$ can connect to at distance $i+1$ from $v_{0}$.
Since the graph satisfies the Decreasing Sequence Property (DSP), we know that all nodes in the graph have a degree at least as large as $v_{0}$’s degree. Therefore, the degree of node $x$ satisfies:
$d(x)\geq d(v_{0})$
This means that node $x$ must have at least $i$ neighbors at distance $i+1$ from $v_{0}$.
The structure of the graph explicitly disallows back arcs, i.e., edges pointing to nodes at earlier distances from $v_{0}$. Thus, $x$ cannot connect to nodes at distances less than $i$ without violating the structure of the graph. This constraint eliminates the possibility of $x$ mapping to earlier-distance nodes.
Given that $x$ must connect to at least $i$ nodes, and since it cannot connect to earlier-distance nodes (due to the disallowance of back arcs), the only remaining option for $x$ is to map to interior nodes—nodes that are at distance $i$ or higher.
Thus, since $x$ must map to at least $i$ distinct nodes, and all these must be interior nodes (by the elimination of back arcs and exterior connections), we conclude that:
$|int(v_{i-1},x)|\geq i$
This shows that the interior degree of $x$ with respect to $v_{i-1}$ is at least $i$.
Therefore, we have shown that for a node $x$ at distance $i$ from $v_{0}$, the interior degree of $x$ with respect to $v_{i-1}$, $|int(v_{i-1},x)|$, is at least $i$, as required.
$|int(v_{i-1},x)|\geq i$
Now that the exterior has an upper bound on it’s size we can see how that impacts the interior degree. Because the degree of each node is still impacted by the minimum degree node, they still need to have their out-degrees go somewhere. Lemma 8.4 (Interior Degrees Lower Bound) compliments 8.3 (Strong Decreasing Exteriors Lemma) by saying thaw each decrease in the exterior degree will be matched by an increase in the interior degree.
9 Transitive Triangles and Seymour Diamonds
Before we continue with the development of our DSP algorithm and show how it solves the Seymour Second Neighborhood Conjecture, we need to address some potential problems and show that they will not hinder our algorithm.
Definition 9.1.
Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
Here, $x$ points to both $y$ and $u$, and $y$ also points to $u$. The ”transitivity” comes from the fact that the direction of the edges follows a logical progression: $x\rightarrow y$ and $y\rightarrow u$ implies a pathway $x\rightarrow u$ via $y$. The concept of a transitive triangle is significant in understanding how the structure of the graph influences properties like out-degrees in $G^{2}$.
In a transitive triangle, $x$ has a direct path to both $y$ and $u$, and the existence of the edge $(y,u)$ implies that the out-degree of $x$ in $G^{2}$ could potentially increase because the paths from $x$ to nodes reachable from $y$ and $u$ are being counted. Understanding how transitive triangles are distributed within the graph can help in analyzing the neighborhood structures and their effects on the square graph, is critical for analyzing the Seymour’s Second Neighborhood Conjecture.
Definition 9.2.
Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
The Seymour Diamond is a graph theory concept in graph there where a node $x$ points to two out-neighbors $u_{1}$ and $u_{2}$, and these two out-neighbors both share the same out-neighbor, $y$. This is basically a four vertex oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and more importantly the Seymour’s Second Neighborhood Conjecture.
The notion of interior neighbors is another way of describing what [1] referred to as transitive triangles. They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e. one where every node has the decreasing sequence property. We provided a similar result with Lemma 5.4 (Interior Degree Greater than 1). When a minimum degree node $v_{0}$ has the decreasing sequence property, all of its neighbors have interior degree of at least 1. That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$ will form a transitive triangle. We can extend that further for any two constructive containers.
xyu
Lemma 9.1.
Transitive Triangles from Containers In a graph $G$ represented by a GLOVER, for any two consecutive containers $C_{i}$ and $C_{i+1}$ in a graph satisfying the Decreasing Sequence Property (DSP), there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle. That is, there exist directed edges from $u$ to $w$, $v$ to $w$, and from $u$ to $v$, or from $v$ to $u$.
This lemma highlights the relationship between a minimum degree node and the formation of transitive triangles, providing a foundation for discussing the implications related to the Seymour conjecture.
Lemma 9.2.
Seymour Diamonds from Containers
In a graph $G$ represented by a GLOVER, for any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
$u\in C_{i},\quad v,w\in C_{i+1},\quad y\in C_{i+2}$
The following arcs hold:
1. There are forward arcs $u\to v$ and $u\to w$, 2. There are forward arcs $v\to y$ and $w\to y$, 3. No back arcs exist between containers.
This lemma effectively describes the conditions for forming a Seymour diamond and links it to the implications regarding the Seymour conjecture. Because we can extend the previous lemmas to encompass any nodes that have the DSP. This inclusion emphasizes that transitive triangles and Seymour diamonds have been integral to our analysis rather than overlooked.
xywu
Lemma 9.3.
Seymour Diamonds and Decreasing Path Algorithm Let $G$ be an oriented graph represented by the GLOVER, in an attempt to prevent a degree during node, where we are applying the DSP path algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP path algorithm.
Proof.
  • A Seymour diamond consists of two vertices, $x$ and $u$, connected by two parallel paths via $y$ and $w$.
  • The DSP path algorithm can utilize one of these paths while disregarding the other, as both connect $x$ and $u$, thus maintaining degree constraints.
  • The presence of redundant paths does not diminish the outgoing neighbors of any node, ensuring that the minimum degree condition holds throughout the process.
  • Consequently, these additional connections offer alternative routes without introducing cycles or otherwise disrupting the algorithm’s operation.
None of these case are a deterrent to the execution of the DSP Path algorithm. ∎
Lemma 9.4.
Interaction of Minimum Degree Nodes and Transitive Triangles Let $G$ be an oriented graph represented by a GLOVER, containing a minimum degree node $v_{0}$ with degree $d$. Consider a transitive triangle formed by nodes $z$, $y$, and another vertex $x$, where $z$ and $y$ are part of this triangle. The following cases are possible:
  • If $d^{+}(y)=d$ and $d^{+}g(z)=d$
  • If $d^{+}(y)=d$ and $d^{+}(z)>d$
  • If $d^{+}(y)>d$
  • If $d^{+}(y)<d$
Proof.
  • Case 1: If $d^{+}(y)=d$ and $d^{+}(z)=d$: Both $y$ and $z$ have exactly $d$ outgoing connections. For the triangle to be transitive, they must not share all neighbors, meaning there exists at least one neighbor of $y$ that is not a neighbor of $z$. Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
  • Case 2: If $d^{+}(y)=d$ and $d^{+}(z)>d$: $z$ covers some neighbors of $y$, but since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
  • Case 3: If $d^{+}(y)>d$: $y$ has more outgoing connections than required. It can utilize these additional connections to map to unique neighbors without impacting the minimum degree condition, ensuring $y$ retains neighbors not covered by $z$.
  • Case 4: If $d^{+}(y)<d$: This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections. Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
Understanding transitive triangles and Seymour diamonds enriches our analysis of oriented graphs, demonstrating the interplay between different graph structures. This exploration contributes to the overall argument and leads us to our main theorem.
10 Main Theorem: Seymour’s Conjecture is True
10.1 Introduction
In this pivotal section, we present the main theorem derived from our analysis of oriented graphs. We articulate the theorem clearly, outlining its implications and the conditions under which it holds.
This paper has been a battle between the chosen minimum degree node and the Decreasing Sequence Property (DSP), with the former acting as a catalyst pushing things forward and the latter acting as a bottleneck, trying to hold things back. Along the way, we have proved several key results, most notably that:
  • Back arcs cannot exist,
  • Interior degrees double inside each container,
  • The interior degree doubling causes a decrease in exterior size.
This interaction between interior degree increases and exterior size decreases drives the structure of the graph. However, this degree increase is bounded by the bound $\frac{n-1}{2}$ because of the size of the containers, which provides a limit on the maximum average degree inside a container. Once the system reaches this maximum bound, the exterior degree of a parent container will double, causing a chain reaction where both interior and exterior degrees double, ultimately leading to the node’s overall degree doubling as well.
We first restate the Lemma 6.5 (Non-Existence of Back-Arcs), the No Back Arcs Lemma from section 6, but slightly rephrased.
Lemma 10.1.
Let $G$ be an oriented graph represented by a GLOVER, in an attempt to prevent a degree doubling node, with all nodes satisfying the Decreasing Sequence Property (DSP), where $v_{0}$ is a minimum degree node in $G$, and all nodes are placed into containers. If there are an back arcs in $G$, then a node’s degree will double.
10.2 Theorem Statement
Degree Doubling by Exterior Decreasing
Theorem 10.1.
Let $G$ be an oriented graph represented by a GLOVER with all nodes the Decreasing Sequence Property (DSP). If all nodes in $G$ are placed into containers, where interior degrees are maximized, then Seymour’s conjecture holds true. Specifically, a node in $G$ will have its exterior size decrease to less than 2.
Algorithm 1 Algorithm for Analyzing Degree Doubling in Containers
1:Initialization:
2:Choose the minimum degree node $v_{0}$.
3:Divide nodes into containers $C_{1},C_{2},\ldots$ based on their degree and proximity to $v_{0}$.
4:for each container $C_{i}$ in $G$ do:
5:     Locate a node $x_{i}$ in $C_{i}$ whose interior degree doubles.
6:     Move to a neighboring container $C_{i+1}$.
7:     if there is a back arc from $C_{i+1}$ to a prior container then
8:         Locate child of $x_{i}$ whose exterior degree doubles, causing $x_{i}$’s overall degree to double.
9:         Halt.
10:     else
11:         Since there is no back arc, the container $C_{i+1}$ will be bound in size to be 1 less than the size of container $C_{i}$.
12:         Repeat from Step 4.
13:     end if
14:end for
Start with graph size $n$Select minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots$Identify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to prior containerLocate child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
Figure 10: Flowchart depicting the degree-doubling algorithm applied within containers, demonstrating the process of identifying nodes with the Decreasing Sequence Property.
Corollary 10.1.
Degree Doubling by Container Density In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},C_{2},...,C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable. This collision can manifest in two forms: either the size of the container reaches a point where it can no longer support the increasing interior degrees (resulting in nodes that cannot maintain their minimum degree requirements), or smaller containers force a doubling of degrees that contradicts the capacity for interior arcs. The structure of the graph, particularly in small sizes, inherently limits these growth dynamics, leading to contradictions that must be addressed.
Corollary 10.1 (Degree Doubling by Container Density) captures the insight exploration our interior degree growth dynamics within shrinking containers in an oriented graph. The language highlights the inevitable ”collision” and ties it back to fundamental graph constraints. Hopefully, this framing gives a precise summary of why these graphs have a limited capacity to sustain both shrinking containers and rising interior degrees without contradictions.
The main theorem, algorithm, and corollary encapsulates the findings of our research, providing a crucial result that advances the understanding of oriented graphs. The proofs and implications of this theorem will be further explored in the following sections.
Corollary 10.2.
[…]
$i=\frac{k-1}{2}$
where $k$ is the minimum degree of $G$. For even values of $k$, the complete graph $K_{n}$ will not occur on any actual even number of nodes but may appear in the interval between two odd numbers.
Proof.
We begin by restating some previously shown examples:
  • For $k=3$: $K_{3}$ (triangle) occurs in $C_{1}$.
  • For $k=5$: $K_{5}$ occurs in $C_{2}$.
  • For $k=7$: $K_{7}$ occurs in $C_{3}$.
These observations hold because the GLOVER structure ensures that each container $C_{i}$ increases in regularity and density as the degrees of nodes double, creating the conditions necessary for complete graphs to emerge.
For even values of $k$, such as $k=4$ or $k=6$, the symmetry and density required for a complete graph $K_{n}$ cannot form directly at any container level. This is because the degree-doubling effect and the odd-number dependency of $K_{2i+1}$ mean that the structural requirements for even $n$ are never satisfied in full. Instead, the structure transitions through intermediate states between odd complete graphs, and complete graphs may appear to ”exist” in a fractional or approximate form between $K_{2i-1}$ and $K_{2i+1}$.
The relationship $i=\frac{k-1}{2}$ reflects the interplay between the minimum degree $k$ and the growth of density in containers. This relationship confirms that the GLOVER structure consistently generates complete graphs $K_{2i+1}$ in odd-numbered increments, with no direct occurrence for even $n$.
This emergence of orientations of complete graphs as containers decrease in size adds a layer of symmetry and beauty to the Seymour Second Neighborhood Conjecture. There was never a question of these arising. Some questioned multiple nodes. Some questioned zero nodes. A complete graph occurrence though is a thing of beauty and it helps give rise to another result.
Corollary 10.3.
Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $k$, and let the complete graph occur in container $C_{i}$. Then every node in container $C_{i}$ has its degree doubled in the representation.
The discovery of multiple degree-doubling nodes transforms the Seymour Second Neighborhood Conjecture from a question of existence to one of abundance. This result not only confirms the conjecture but reveals an inherent richness in the structure of oriented graphs. The presence of multiple nodes is tied to the emergence of complete graph structures as containers decrease in size, showcasing an element of mathematical beauty. Moreover, this has real-world applications: in network analysis, it suggests that influence or importance may not rest with a single node but instead within a clique or tightly connected group, offering new perspectives for studying collaborative dynamics and group influence.
The identification of regularity was key to understanding why multiple nodes’ degrees can double in certain graphs. This concept offers a powerful explanation for the phenomenon. The regularity is the key factor that helps explain why multiple nodes can have their degrees double, rather than the search for multiple nodes being inherently complex. The regularity is the central feature that ties everything together.
Degree doubling is essentially a consequence of the underlying regularity. When a graph’s structure is highly regular, particularly in terms of interior and exterior degrees, it becomes easier to predict and verify that certain structural properties (like degree doubling) will occur at specific stages.
Regularity implies that as nodes increase in distance from the starting node, their connectivity and interior degrees will follow a pattern. This pattern is key to understanding why degree doubling happens more naturally and why multiple nodes can experience it simultaneously, if the conditions allow.
10.3 Decreasing Sequence Property Path Algorithm
We now present a precise algorithm to identify nodes whose out-degree doubles in the square of an oriented graph, as per Seymour’s Second Neighborhood Conjecture (SSNC). This algorithm synthesizes results from the lemmas, theorem, and corollaries discussed throughout the paper, providing a cohesive framework for finding degree-doubling nodes.
  • The containers reach a size two causing a node in the previous container to have its degree double, as stated in Theorem 10.2 (Degree Doubling by Exterior Decreasing).
  • The interior degree required for nodes inside containers will be greater than the possible for a container of that size causing a node in the previous container to have its degree double, as stated in Corollary 10.1
In both these cases we are able to find a node whose degree doubles.
Here is the algorithm in pseudocode. Input: An oriented graph $G$ with minimum out-degree $d$.
Output: A node $v$ in $G^{2}$ whose out-degree at least doubles.
Algorithm:
1. Initialize the Graph Structure: - Identify the minimum degree node $v_{0}$ in $G$. - Partition $G$ into containers $C_{1},C_{2},\ldots,C_{d}$, where each container $C_{i}$ groups nodes by their distance $i$ from $v_{0}$.
2. Check for Back Arcs: - For each container $C_{i}$, check for arcs leading to any container $C_{j}$ where $j<i$. - If a back arc exists from a node $u_{i}\in C_{i}$ to a node $v_{j}\in C_{j}$, record $u_{i}$ as a degree-doubling node by Lemma 7.2 (Interior Degree Doubling) and Lemma 6.5 (Non-Existence of Back Arcs). - If no back arcs exist, proceed with Step 3.
3. Establish Connectivity Properties: - For each node $v$ in container $C_{i}$, define: - Interior neighbors $\text{int}(v)$: Nodes within the same container. - Exterior neighbors $\text{ext}(v)$: Nodes in the next container.
4. Iterative Degree-Doubling Check: - For each container $C_{i}$: - Calculate the out-degree of each node in $C_{i}$. - If a node in $C_{i}$ has fewer second neighbors than first neighbors (DSP violation), record it as a potential degree-doubling node.
5. Decreasing Sequence Property Verification: - Verify that the size of each exterior set $\text{ext}(v)$ is strictly less than the size of the corresponding interior set $\text{int}(v)$ in $G^{2}$. - If not, apply the conditions set in Lemma 7.2 to ensure any doubling in interior degrees results in a matching exterior degree.
6. Termination and Output: - Return the identified node(s) with doubled degrees as proof of degree-doubling in $G^{2}$.
11 Conclusion and Future Work
Solving a conjecture with such a long and debated history requires thorough validation. This paper represents a comprehensive effort to meet the rigorous standards of mathematical proof, supported by tools designed to aid understanding. Acknowledging that such results are subject to detailed scrutiny, this process is viewed as an essential step in establishing confidence and recognition within the mathematical community.
The GLOVER structure—short for Graph Level Order—represents a significant advancement in addressing long-standing open problems in graph theory, such as the Seymour conjecture, which remained unresolved for 34 years. While not all oriented graphs are inherently well-ordered, the GLOVER structure reveals that, within the context of the Seymour conjecture, oriented graphs become effectively well-ordered. This property arises because any deviation from such an ordering would lead to a degree-doubling effect, rendering the conjecture invalid.
By leveraging this induced well-ordering, the GLOVER structure enables tools like mathematical induction to be applied effectively, facilitating a solution to the Seymour conjecture. Beyond this specific case, the framework provides a foundation for exploring dual metrics in graph theory, such as prioritizing distance and degree, and opens the door to applications in clustering, ranking, and multi-metric optimization.
The GLOVER structure emerges as a groundbreaking framework for addressing problems that rely on dual metrics, offering a versatile tool for organizing and analyzing data across theoretical and practical contexts. While its utility was demonstrated here in the context of the Seymour conjecture—where the dual metrics of distance and degree played a pivotal role—it is important to recognize that the structure’s applications extend far beyond regularity or degree-based problems.
At its core, the GLOVER structure’s ability to prioritize and organize relationships based on dual metrics allows it to address challenges that require both exhaustive and hierarchical searches. Unlike local search methods traditionally constrained by single metrics, the GLOVER approach enables partitioning and clustering in ways that preserve relationships and provide deeper insights. For instance, applications in clustering, ranking, and network analysis—whether in social networks, employee rankings, or other multivariate datasets—can benefit from this structure.
Looking ahead, the GLOVER structure challenges us to investigate its applicability to other open problems in graph theory and computer science. The approach may uncover new insights in contexts where dual or multi-metric prioritization is key, highlighting the broader potential of this versatile framework.
Future research and development should explore these broader applications. Questions of complexity, scalability, and adaptability to non-graph-based datasets remain promising areas for investigation. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
12 Acknowledgments
I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:6 has been a guiding light throughout this journey. I am profoundly grateful to my family—my mother, father, and wife—for their unwavering support and encouragement.
Special thanks to the late Dr. Nate Dean for his contributions to this work and to Dr. Michael Ball for his support. I want to sincerely thank everyone who has taken the time to pray for me, share encouraging words, and offer constructive feedback. Your support has meant so much to me, and I’m truly grateful for the kindness and thoughtfulness you’ve shown.
Finally, I extend heartfelt gratitude to my family and friends for their patience and belief in me throughout this journey.
The GLOVER data structure is protected by a provisional patent application filed with the U.S. Patent and Trademark Office.
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Ranked Structure: The elements of $S$ are partitioned into a ranking of ”containers” or ”levels,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Definition 13.4.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the interior neighbors of $y$ with respect to $x$, denoted as $int(x,y)$, are the out-neighbors of $y$ that are also out-neighbors of $x$. Formally,
$int(x,y)=N^{+}(x)\cap N^{+}(y),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$ in the graph. The number of interior neighbors of $y$ with respect to $x$ is called the interior degree of $y$.
Definition 13.5.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the exterior neighbors of $y$ with respect to $x$, denoted as $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$. Formally,
$ext(x,y)=N^{+}(y)\cap N^{++}(x),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$, and $N^{++}(x)$ denotes the set of out-neighbors of the out-neighbors of $x$ (i.e., the second neighbors of $x$). The number of exterior neighbors of $y$ with respect to $x$ is called the exterior degree of $y$.
Definition 13.6.
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
Definition: Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
Definition 13.7.
Definition: Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
14 Appendix 2: Applicability of the Graph Level Order
Traditionally, a total order on a set is a linear, one-dimensional comparison between elements. The key properties of total orders—reflexivity, transitivity, and antisymmetry—give us a clear, unambiguous ranking where every pair of elements is comparable. Graph Level Orders (GLOVER) extend this concept to n-dimensions within a leveled framework.
GLOVER accomplishes this through the use of containers that not only arrange elements in sequence but also enable multiple levels of organization. Elements within each level maintain an internal order, while the containers themselves are comparable across ranked layers. This structure allows for sophisticated comparisons both within a single level and between levels, making GLOVER a higher-level mathematical framework suitable for multi-layered, multi-dimensional data.
The examples provided in this paper illustrate concrete cases of GLOVER. However, it’s the abstract nature of GLOVER that sets it apart as a transformative mathematical concept. This structure allows us to conceptualize any set of ranked organized objects—not just graphs with regularly structured nodes, but also diverse entities, such as sets with multiple attributes or time-sequenced data. Within discrete mathematics, the ability to impose a multi-dimensional total order provides a richer understanding of relationships among elements. The order goes beyond pairwise comparisons to establish a multi-layered framework capable of supporting complex operations.
The true strength of GLOVER lies in its flexibility. It accommodates inherent ordering within single-level sets while also adapting to more intricate structures. By generalizing beyond domain-specific contexts like graph theory, GLOVER becomes a universal framework for imposing total order on any ranked structure, from simple ordered pairs to multi-dimensional data sets in fields like machine learning or clustering rankings.
The GLOVER structure also uniquely supports induction in graph-based proofs due to its total order, providing a new way to approach graph theory problems. Unlike traditional graph structures, which typically do not allow for induction, the total ordering of nodes and containers in GLOVER makes it possible to apply inductive reasoning effectively.
This expanded understanding of total orders paves the way for developing algorithms, proofs, and applications capable of handling complex, scalable data structures. GLOVER is not merely a form of ordering; it is a system of relationships that facilitates the manipulation, comparison, and analysis of multi-dimensional data through a clear and consistent layering.
The GLOVER framework, as initially defined, is tailored to address the specific requirements of the Seymour Second Neighborhood Conjecture. Its rigid structure-classifying arcs either as interior or exterior, and strictly prohibiting skipping levels-ensures clarity and sufficiency for analyzing second neighborhoods, However the framework can be extended to accomodate more general problems, including those with nested containers or more flexible neighbor relationships .
Nested containers introduce a hierarchy within individual containers, grouping nodes based on finer structural or relational properties. This extension enables the framework to address problems where additional levels of granularity within the same partition group are required.
Definition 14.1.
A nested container $C_{i,j}$ is a subset of a container $C_{i}$ defined by additional criteria beyond the qualifying metric that determined it into $C_{i}$. For example, nodes in $C_{i,j}$ might share specific adjacency patterns or connectivity property properties. Use Cases:
  • Clustering Problems: Nested Containers can represent hierarchal clustering within each distance group, allowing comparisons between subgroups.
  • Weighted Graphs: In graphs with edge weights, nested containers could group nodes based on weight thresholds within the same distance level.
Example:
Suppose $C_{2}$ contains nodes $x,y,z$ all at distance 2 from $v_{0}$. If $x$ and $y$ are connected but not $z$, $x$ and $y$ might form a nested container $C_{2,1}$, while $z$ forms $C_{2,2}$
In certain problems, arcs may not connect nodes in non-adjacent containers, effectively skipping levels in the distance hierarchy. While the SSNC prohibits such arcs to maintain a second-neighbor focus, other graph probems may allow or even require these connections.
  • Extension - Relax the exterior neighbors to include $C_{j})$ for $j>i+1$, introducing a new category of arcs, long-distance exterior arcs.
  • Use Cases -
    • Network Design - In communication networks, skipping levels might model direct long-range connections (for example satellite links).
    • Flow Problems - In transportation or flow networks, skipping levels could represent shortcuts or direct transfers.
  • Impact on GLOVER Properties - Allowing for skipping levels introducing new challenges
    • Order Maintenance - Total order within containers must be preserved despite the additional complexity of the long-distance connections.
    • New Arc Classifications - Introduce distinctions like short range exterior arcs (j = i + 1) and long range exterior arcs (j ¿ i + 1).
Exterior neighbors, as defined in the SSNC, connect nodes in adjacent containers. Expanding this definition allows for more flexible interactions between containers, broadening the applicability of the GLOVER framework.
  • Generalized Exterior Neighbors - for $x\in C_{i}$, an exterior neighbor $y$ can belong to any $j>i$. The specific range of $j$ will depend on the problem context.
  • Example Problem: Consider a graph where connections beyond second neighborhoods are of interest. Exterior neighbors could include nodes up to $k$-steps away, accomodating problems requiring $k$-neighborhood analysis.
While these extensions offer exciting generalizations, the SSNC’s specific focus on second neighbors necessittates a stricter interpretation of the GLOVER framework.
  • No skipping levels - Ensures that second neighbors are strictly defined nodes in $C_{2}$.
  • Rigid container definitions - prevents ambiguity in arc classifications or neighbor interactions.
  • Limited use of nested containers - The SSNC does not require additional granularity within distance groups, making nested containers optional for this specific problem.
Example 14.1.
Graph Theory - Seymour Second Neighborhood Conjecture: Graphs, which are normally seen as an unordered arrangement of nodes and edges, were able to be ordered with the GLOVER. We initially chose a minimum degree node and partitioned the nodes of the graph into containers based on their proximity to this node.
Partitioning nodes by distance allowed for definition of the interior degree and exterior degree which were very important to showing that a node’s overall degree would double.
Example 14.2.
Engineering - Dynamic Scheduling: In fields such as manufacturing, task scheduling often involves prioritizing independent tasks. With GLOVER, tasks are grouped into containers based on dependency levels and ordered according to priority. For example, preparatory tasks are prioritized in early containers, while finishing tasks are sequenced last, minimizing delays and optimizing task flow.
Partitioning groups tasks by dependency, while ordering between containers ensures essential tasks are completed first. Within each container, tasks can be further organized by resource requirements, promoting efficiency at each priority level.
Edges introduced at the container level to represent this dependency. Interior edges would show common dependencies between tasks in a common container, whereas exterior edges represent tasks of different priorities with common dependencies.
From this we can use graph theory concepts like centrality to better understand things like important tasks, important dependencies, how the containers (task tiers relate to one another) etc.
Example 14.3.
Engineering - Resource Management: In network resource allocation or energy grid management, nodes can be clustered by proximity to demand or criticality and ordered to prioritize high-need resources. This ordering ensures that essential nodes remain accessible, while resource usage remains balanced across the network.
Partitioning groups resources based on proximity to critical nodes, while ordering between containers directs allocations toward high-demand areas. Within each container, resources are accessed in order of current availability, optimizing responsiveness to fluctuating demands.
Example 14.4.
Engineering - Project Management: In complex projects involving multiple teams and dependencies, tasks can be clustered by project phase and ordered sequentially to prevent bottlenecks. Foundational tasks, such as design and planning, are prioritized early, while execution tasks follow, ensuring smooth progression through each phase.
Partitioning groups tasks by project phase, while ordering between containers facilitates the transition of tasks from one stage to the next. This ranked order remains stable despite potential project changes, enabling efficient coordination across large teams.
Example 14.5.
Cybersecurity - Threat Detection and Prioritization: In cybersecurity, ranked structures are often essential for organizing and prioritizing threats. Using an GLOVER framework, threats can be grouped into containers based on their urgency or potential impact. For example, critical threats with known vulnerabilities might be prioritized in early containers, while less immediate risks or less probable attack vectors are sequenced into subsequent layers. This ranked ordering helps ensure that the most pressing issues are handled first, thereby minimizing potential breaches or disruptions.
Partitioning into containers groups threats by their severity and probability, while ordering between containers facilitates a systematic approach to addressing risks from the highest priority down. Within each container, threats are further organized based on attributes like type, location in the network, or historical occurrence, allowing security teams to deploy resources effectively. This layered approach enhances the overall defense strategy by ensuring structured response and continuous monitoring across multiple risk levels.
Example 14.6.
Natural Language Processing - Ranked Text Classification: In NLP, ranked total order graphs can improve text classification by organizing data through multi-level categories. For instance, documents can be grouped into containers based on broad topics (e.g., news, entertainment, scientific research) with each container further ordered by subtopics like sports or politics under news, and biology or physics under science. This layered structure allows models to systematically refine classifications, progressing from general to specific categories, improving the accuracy and interpretability of the classification model.
Partitioning into containers groups text by broad categories, while ordering between containers prioritizes classification tasks from general to specific. Within each container, documents can be ordered by relevance, keywords, or thematic clustering, which supports tasks like topic modeling or intent recognition. This approach provides a clear and adaptable framework for complex, multi-layered categorization, enhancing the model’s ability to handle diverse and evolving datasets.
The versatility of the Graphs Level Order (GLOVER) structure hints at its applicability beyond this proof of the Seymour Second Neighborhood Conjecture. GLOVER has the potential to address several open problems and conjectures within discrete mathematics and graph theory, where ranked or layered ordering of elements can provide new insights or simplify complex relationships. Future work may reveal more connections between GLOVER and these longstanding challenges, offering new paths for research and exploration.
15 Appendix 3: Visualization and Accessibility of Oriented Graphs
In exploring oriented graphs, it became evident that fully grasping these structures often requires more than static illustrations in a paper can offer. While oriented graphs can be defined and studied abstractly, their complexity and behavior are best understood through dynamic, interactive representations. The Seymour Second Neighborhood Conjecture, in particular, necessitated defining paths, locating squares, and identifying degree-doubling nodes—tasks that involved intricate arc counting and relationships that benefit from visualization.
This realization motivated the creation of a complementary website. The site serves as a tool to enhance the accessibility of key definitions, lemmas, and proofs by providing visualizations and interactive examples. Our goal is to bridge the gap between abstract theory and intuitive understanding, making the intricate aspects of the conjecture approachable for a wider audience.
For example, consider the concept of a degree-doubling node or a lemma describing staircase-like structures in oriented graphs. While the paper describes these ideas in formal mathematical terms, the website illustrates them dynamically. This approach allows readers to ”see” the behavior and implications of these structures, offering a visual intuition that complements the formal reasoning presented in the paper.
This methodology draws inspiration from James Robert Brown’s insights in Philosophy of Mathematics, where he emphasizes the dual role of visualization in mathematics—as both a powerful tool for intuition and a potential pitfall if over-relied upon. By carefully balancing these aspects, we aim to enhance comprehension without compromising the rigor essential to mathematical proofs.
The website also provides an opportunity to include numerous examples and extended visual explanations that are not feasible within the space constraints of a traditional journal format. Through this medium, we invite readers to engage more interactively with the conjecture and its underlying theories, fostering a deeper understanding and encouraging constructive feedback.
Built using technologies like D3.js, JavaScript, and PHP, the website features dynamic visualizations of proof steps, interactive diagrams, and supplementary explanations. By demonstrating the synergy between mathematical results and computational tools, we aim to make these concepts not only accessible but also engaging for a broader audience.
Readers are encouraged to explore the website for additional insights and examples at:
https://glovermethod.com
While the website is already live and functional, it is an evolving project. Additional examples, applications, and refinements are being added continuously.
Readers are invited to explore the website and are encouraged to provide feedback on any aspects of the site or the content itself, including suggestions for improvements, corrections, or additional examples.
The goal is to create a resource that not only supports this paper but also serves as a tool for further exploration and discussion in the field of oriented graph theory, the Seymour Second Neighborhood Conjecture, and the Graph Level Order (GLOVER) data structure. Thank you for your engagement and contributions to this ongoing effort.