### v4

# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. This paper first presents a data structure GLOVER (Graph Level Order) which is a partitioning of the nodes into a total order of neighborhoods. This data structure allows us to introduce a total order on oriented graphs and construct a decreasing sequence of subsets. This decreasing sequence of subsets is what ensures that no counterexample can exist. Further, this algorithmic approach allows us to know where the node is. Later findings extend the results to multiple nodes. This is true for all oriented graphs. These findings also have practical applications in network optimization and algorithm design.
1 Introduction
Seymour’s Second Neighborhood Conjecture (SSNC) was originally proposed by Paul Seymour in 1990. It asks the question of the existence of a node in an oriented graph whose out-degree at least doubles in the graph’s square. The conjecture was first published by Nathaniel Dean and Brenda J. Latka in 1995 [4], where they posed a similar conjecture specific to tournaments. This version of the conjecture was first proven by Fischer [8], with a second proof provided by Havet and Thomasse [10].
Conjecture 1.1.
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}$ represent the first and second neighborhoods of $v$ respectively.
This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [9]). Here, we hope to tackle the entire class of oriented graphs based on the structure laid out in that paper.
The Seymour Second Neighborhood Challenge is a difficult graph theory problem. The problem itself asks us to focus on the question of oriented graphs and second neighbors. There have been several papers that have looked into special cases. Kaneko and Locke [11] showed that the conjecture was true for graphs with a minimum degree of at most six. Brantner et al. [1] showed that, if the conjecture is false, infinitely many counterexamples must exist. Additionally, Chen et al. [3] provided a lower bound for how much a node’s out-degree increases in the graph’s square. They concluded that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Bolter et al [2] where their work confirms that almost all orientations of $G(n,p)$ are satisfy the Seymour Second Neighborhood Conjecture.
The topic of $m$-free graphs is of particular interest to this paper because Daamouch’s [7] work is very similar to what we see here. In his paper, he introduces induced subgraphs and speaks of their roles in degree properties on vertices, which will also be a key thing in this paper.
The Seymour Second Neighborhood Conjecture is not strictly a graph theory problem. Although it has historically been studied within the field of mathematics, it has practical applications across several domains, including social media, social networking, epidemiology, disease control, and many other fields. In the critical context of disease control, the nodes in this conjecture can represent individuals, and the orientation of the arcs symbolizes one person infecting another. Here, the application of SSNC becomes relevant by asking: Is there a person whose influence in their second radius of interaction (i.e., people they indirectly connect with) exceeds the number of people in their first radius (i.e., direct contacts)?
In this paper, we will prove the Seymour Second Neighborhood Conjecture is true. We will do this by first assuming that it is false and attempting to construct a counterexample. We will explore the conditions under which no node’s out-degree doubles will double in the square of the oriented graph. This approach will focus initially on minimum out-degree nodes. We will analyze how their out-degree could fail to double. We will then extend out algorithm to other nodes based on their distance from the minimum out-degree node.
The paper is organized as follows: Section 2 begins with some common terminology used in graph theory. Section 3 introduces our approach to the problem. In section 4 we look at some small examples of this style of thinking. Section 5 introduces the terminology and identifying the conditions under which their out-degree might fail to double. Section 6 examines the back arcs and their place in oriented graphs. Section 7 looks at the conditions which would cause an interior degree to double, and its importance to the Seymour Second Neighborhood Conjecture. Section 8 further looks at containers under the context of density constraints. In Section 9 considers transitive triangles and Seymour diamonds and show that they are not a hindrance to our algorithm. Section 10 presents this process the full Decreasing Sequence Property algorithm as well as the theorem stating it’s solution to the Conjecture.
2 Graph Theory Terminology
A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $(x,x)$ where $x$ is a node in $G$) and no symmetric arcs, that is no arcs of the form $(x,y)$ and $(y,x)$ where $x$ and $y$ are nodes in $G$. The square of a graph $G$, denoted $G^{2}$, consists of the same node set as $G$ and an arcs exists from node $x$ to node $z$ if there is a node $y$ such that both arcs $(x,y)$ and $(y,z)$ exist in $G$ as well as the arcs already in $G$. Essentially, $G^{2}$ captures two-step connections from $x$ to $z$.
Definition 2.1.
Let $G^{2}=(V,E^{2})$ where $G=(V,E)$ is the original graph, and $E^{2}$ is the set of edges defined as:
$E^{2}={(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E}$
We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$. These are the nodes that $v$ has arcs pointing to. Similarly, $N^{++}(v)$ refers to the second neighbors of $v$ that is, nodes that are two steps away from $v$ in $G$.
Let $G=(V,E)$ be a graph and let $S\subseteq V$ be any subset of vertices. The induced subgraph $G[S]$ is the graph whose vertices are the set $S$, and whose edge set consists of all edges in $E$ that have both endpoints in $S$. For any node $v$, we refer to the neighbor-induced subgraph on $N^{+}(v)$, consisting of $v$’s out-neighbors. Let $u,v$ be nodes in $G$. The distance between $dist(u,v)$, is $k$ if the shortest path from $u$ to $v$ has length $k$.
Example 2.1.
[…]
Figure 1: A five-vertex graph illustrating node out-degrees and distances. Out-degrees are labeled near each node, and the dashed rectangle indicates an induced subgraph on vertices A, B, and C. Previously used in [9]
3 Approach and Motivation
We begin our analysis by considering general oriented graphs. We make no initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs. We know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
From this foundation, we can now partition the graph into $d_{v_{0}}+1$ subgraphs corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm. No nodes are being removed to form these partitions. This is a conceptual idea. Instead, Think of it as putting a box around the partition. Now the nodes inside the box are can work with the other nodes inside the box and solve their problem independent of the other partitions. These partitions are still connected to one another, but thinking of them in this way, allows us to do more.
This represents a stark contrast to most previous approaches to the Seymour Second Neighborhood Conjecure. Many of these approaxhes were conceptual or theoretical, saying only that a node whose degree doubled existed, but the notion of a path to that node was not part of the paper. Secondly, most of these paper focused on local search techniques that statrted at or near the desired degree doubling node. That presents a problem because oriented graphs can be varying in nature, This is the first exhaustive search technique on the SSNC problem.
We have laid the groundwork for our algorithm to tackle the oriented graphs. With a minimum degree node, we can partition this graph into independent subproblems. This is now ready for an exhaustive search algorithm.
4 Initial Lemmas
These first lemmas will seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. For this section, unless otherwise noted, assume that $v_{0}$ is a node with minimum out-degree in our oriented graph $G$. We will also introduce some standard notation from graph theory. The work in this section agrees with what was already done by [11] and is not meant to serve as original content. It does however, provide insight into our reasoning.
Example 4.1.
$v_{0}$$v_{2}$$v_{3}$$v_{1}$$v_{6}$$v_{5}$$v_{4}$
$\Rightarrow H$
Figure 2: Example illustrating how a minimum out-degree node in $G$ has its degree doubled in $G^{2}$ under certain conditions.
Lemma 4.1.
Minimum Out-Degree $<$ 3 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3, then $v_{0}$’s out-degree will at least double in $G^{2}$.
Proof.
We will consider the cases where $d^{+}(v_{0})=0$, $1$, and $2$.
Case 1: Assume that $d^{+}(v_{0})=0$. Then $v_{0}$ has no outgoing edges in $G$. In $G^{2}$, it will still have out-degree 0. Hence its out-degree trivially satisfies the condition of at least doubling.
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. We know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. This node $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. The node $v_{0}$ cannot be one of these nodes. Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4. Thus, its out-degree will double.
Therefore, in all cases, $v_{0}$’s out-degree at least doubles in $G^{2}$.
Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs. We are learning that, in order to construct our counter-example, we need out minimum degree to be at least 3. This will also place that same minimum value restriction on every other node in the graph. In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
Lemma 4.2.
Minimum Out-Degree 3 with Neighbor Out-Degree 0 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3, and at least one of its neighbors has out-degree 0 in the induced subgraph of its neighbors, then the out-degree of $v_{0}$ will at least double in $G^{2}$.
Proof.
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$. Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$. Now, assume that $v_{1}$, has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$, As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in $G$.
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$ shares no neighbors with $v_{0}$. This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors, and $v_{1}$ must have at least 3 first neighbors.
Example 4.2.
0
$\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$ H
Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$. Each of these nodes ($v_{1}$, $v_{2}$, and $v_{3}$) connects to the nodes $v_{4}$ and $v_{5}$. The neighbor of $v_{4}$ is $v_{5}$, and the neighbors of $v_{5}$ are $v_{6}$, $v_{7}$, and $v_{8}$. While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$. Thus, in $G^{2}$, the degree of $v_{0}$ will double.
Lemma 4.3.
Minimum Out-Degree 3 with Neighbors 1 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3 and all of its neighbors have an out-degree of 1 in the induced subgraph of its neighbors, then $v_{0}$ or one of its neighbors will have its out-degree at least double in $G^{2}.$
Proof.
Assume $d^{+}(v_{0})=3,$ so $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ We know that $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$. Let $G[v_{1},v_{2},v_{3}]$ represent the induced subgraph by $v_{1},v_{2},v_{3}$.
We can reason that if any of $v_{1},$ $v_{2},$ or $v_{3}$ has an out-degree greater than 3, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$.
Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}].$ If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume they all connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree $<$ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one edge between $v_{4}$ and $v_{5}$. One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$. As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$.
Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) answers a question that we will come back to throughout this paper - this is an exhaustive search algorithm. In order for that to work, it has to allow some node’s out-degree to not double. In this case, what we see is that the initial node in the example $v_{0}$ does not have its degree double. Instead, we are forced to search for the next node. That comes at a neighbor of $v_{0}$, $v_{1}$ whose degree does double in this example.
This is an important distinction because we did not need the ”right” starting node. We chose a minimum degree node because of its ability to partition the nodes in the graph, not for its potential closeness to the double degree node.
For the remainder of this paper, we will refer to graphs like those in Lemma 4.2 using double arrows to indicate that a node or set of nodes connects to every node in the subsequent subgraph. This can be done because the nodes within a common induced subgraph are all relating to the same set of nodes.
These lemmas provide insight for our motivation, and a framework for theoretical foundation. What is coming in the next sections will build on these results and will refer back to them.
5 Theoretical Foundation
In order to understand this next section, we will introduce some new terms. We will give a formal name to the partitions we spoke of about previously - containers. We will also speak about the decreasing sequence property and talk about its importance. Then we will talk about interior, exterior and back arcs. Then we will close with the GLOVER data structure.
Example 5.1.
[…]
Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from the minimum degree node, with no assumptions on container relationships or sizes.
Definition 5.1.
Given a minimum degree node $v_{0}$, a container $C_{i}$ is the set of nodes that are at distance $i$ from $v_{0}$, along with the arcs between those nodes. More formally, $C_{i}=\{u\in V(G)\text{ .s.t. }dist(v_{0},u)=i\}\text{ and A}(C_{i})=\{(u,v)|u,v\in C_{i}\}$ ,where $V(G)$ represents the vertex set of the graph $G$, and $dist(v_{0},u)$ is the shortest path distance between $v_{0}$ and $u$.
Daamouch used induced subgraphs similar in his study of $m$-free graphs [7]. The primary differece here is the neighboring containers (see image) which will be important in the proofs.
With the concept of containers we can explore the structure of oriented graphs from two perspectives. We have the option of exploring it from the perspective of the minimum degree node, or from the perspective of containers.
  • Containers are seen based on their distance from the
  • Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
minimum degree node.
The case where $d^{+}(v_{0})=3$, and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) in particular, is an important one because it provides a glimpse into what a counterexample to the conjecture may look like. We were able to construct a graph $G$ where $v_{0}$’s out-degree does not double in $G^{2}$. Simple graphs of out-degree 3 are the smallest possibilities of cycles in undirected graphs. So if we orient the edges in the same direction we can get a cycle of order 3 among the nodes adjacent to $v_{0}$. If we look at this situation further we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. In this section we will begin exploring the sufficient conditions for other nodes in the graph to not have their degrees double.
To continue along this line of reasoning though we will need to define new terms. First we will need a term for a node’s out-degree not doubling. A node in $G$ whose out-degree does not double in $G^{2}$ is said to have the decreasing sequence property. We have seen that examples exist where $v_{0}$ does not have its out-degree double in $G^{2}$. We will look into what other nodes can have the decreasing sequence property, and what properties must exist in the graph in order for this to take place.
Example 5.2.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
Figure 5: This is an illustration where node $v_{0}$ has the Decreasing Sequence Property (DSP). There are more first neighbors $d$ of $v_{0}$ than second neighbors.
Definition 5.2.
For a node $x\in V(G)$ in an oriented graph $G$, we say that x has the decreasing sequence property if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e.,$|N^{+}(x)|>|N^{++}(x)|$.
We have now introduced two concepts: the minimum degree node and the decreasing sequence property (DSP). A minimum degree node, defined as the node with the fewest connections, serves as a critical starting point both for our algorithm and degree values. By selecting it as the starting point it, it serves as the anchor. Now, the minimum degree node is forcing the remaining nodes into the properties that will emerge in the counterexample construction. It is acting as a catalyst.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
Lemma 5.1.
Total Order Induced by Containers Relative to the Minimum Degree Node
Let $G$ be a graph with a minimum degree node $v_{0}$. Further, let all nodes in $G$ satisfy the Decreasing Sequence Property (DSP). Construct a series of containers $C_{1},C_{2},\ldots,C_{k}$ ordered by distance relative to $v_{0}$. Each container $C_{i}$ will consist of nodes at that respective distance from $v_{0}$. Then this structure imposes a total order on the nodes in $G$ i.e. every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
Proof.
To prove a total order, we need to show that every node in $G$ can be compared to every other node in $G$. Let $u,v$ be two nodes in $G$ such that $u\neq v$. It is assumed that $dist(v_{0},u)$ and $dist(v_{0},v)$ are finite for vertices $u$ and $v$ in the containers $C_{i}$. This assumption implies that both vertices are reachable from $v_{0}$, although it does not necessitate that all vertices are connected. We know that distance is a total ordering, so we know that one of three things are true: either $dist(v_{0},u)<dist(v_{0},v)$, $dist(v_{0},u)=dist(v_{0},v)$, or $dist(v_{0},u)>dist(v_{0},v)$. The situation where $dist(v_{0},u)<dist(v_{0},v)$, implies that the node $u$ is in an closer container to $v_{0}$ than the node $v$. The situation where $dist(v_{0},u)=dist(v_{0},v)$ implies that the nodes $u\text{ and }v$ are in the same container. And the situation $dist(v_{0},u)>dist(v_{0},v)$, implies that the node $u$ is in a further container to $v_{0}$. This shows that any two nodes can be compared and shows that it is a total order. ∎
This is a good start as it imposes order on our graph, but the containers do more than just impose a total order. They essentially allow for grouped comparisons among nodes at the same distance from a reference node, $v_{0}$. This is something that a standard total order does not account for. This ”containerized ordering” enables us to distinguish nodes not only by their distance from $v_{0}$ but also by specific structural or adjacency properties within each distance group.
Unlike a simple total order, which could be any sequence with a single, consistent ordering (like lexicographic ordering), the container structure allows nodes at the same distance to be further organized. This organization is significant because it facilitates comparisons within each distance layer (container) by additional criteria, such as adjacency or neighborhood properties.
Definition 5.3.
A Level Total Order (LTO) on a set $S$ can be defined as follows:
  • Total Order: For any two elements $u$ and $v$ within the set $S$, exactly one of the following must hold:
    $u$ is less than $v$ (denoted as $u<v$), $u$ is greater than $v$ (denoted as $u>v$), or $u$ is equal to $v$ (denoted as $u=v$).
  • Leveled Structure: The elements of $S$ are partitioned into a leveling of ”containers” or ”levels,” $C_{1},C_{2},,,,,C_{n}$, with each container itself forming a subset of $S$ that maintains a total order. The containers are sequentially ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
  • Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
  • Comparability within Containers: For any two elements $u,v\in Ci$, their order can be determined based on a specific criterion, such as a designated metric or property relevant to the application.
  • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
This order means that we have order in two ways. First, the containers themselves are ordered. Second, the nodes within the containers are ordered. This ordering must also be consistent within the containers.
The notion of a non-triviality requirement is to distinguish this ordering from other traditional total orderings on graphs.
Lemma 5.2.
Decreasing Sequence Property Lemma If a node in an oriented graph $G$ does not have the decreasing sequence property, then its out-degree at least doubles in $G^{2}$.
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$, and those that are not neighbors of $x$. In the proofs of Lemma 4.2(Minimum Out-Degree 3 with Neighbor Out-Degree 0), and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor induced subgraph (of $v_{0}$) and in the graph with the neighbor induced subgraph removed. We will now give terms to these concepts.
Definition 5.4.
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
$int(x,y)=N^{+}(x)\cap N^{+}(y)$
,where $N^{+}(z)$ is the set of out-neighbors of $z$. The number of interior neighbors is the interior degree of $y$ with respect to $x$.
Now that we have introduced containers, we can talk arcs where both nodes are inside the containers, these are interior neighbors. Interior neighbors will be important to our study on Seymour’s Second Neighborhood Conjecture and our search for a counterexample.
In particular we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1 and we see that when all the neighbors in that subgraph had outdegree of 1, $v_{0}$’s degree did not double.
Definition 5.5.
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The exterior neighbors of $y$ with respect to $x$, denoted $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$:
$ext(x,y)=N^{+}(y)\cap N^{++}(x)$
, where $N^{+}(y)$ is the set of out-neighbors of $y$, and $N^{++}(x)$ is the set of second neighbors of $x$. The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers. This interaction is again critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container.
As an exercise to visualize interior and exterior neighbors, consider example 5.2. Exterior arcs are the double arrows between the boxes. Interior arcs are not drawn but can be mentally placed either between any two nodes in the red or blue box.
Definition 5.6.
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. In order to do so though, we need to consider the option of back-arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to the minimum degree node. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture where they may force us to double count nodes.
Having defined our arcs, we can talk about the decreasing sequence property. This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decreases as we get further from the minimum degree node. Exterior degrees, though, are a part of a node’s degree equation. So there’s another part of it. As we move further from the minimum degree node, the size of the interior degrees increases. This phenomenon reflects a yin and yang relationship between the interior and exterior degrees.
In all of this, the minimum degree node is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s outdegrees.
With this definition of interior and exterior nodes now introduced, we are now ready to now introduce the Graph Level Order (GLOVER) data structure.
Definition 5.7.
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Leveled Structure: The elements of $S$ are partitioned into levels of ”containers” or ”rankings,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
To prove that any oriented graph can be represented in a GLOVER, we’ll need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
Theorem 5.1.
Given an oriented graph $G$ and a minimum degree node $v_{0}$, where the nodes of $G$ are partitioned into containers $C_{1},C_{2},\dots,C_{k}$ based on their distance from $v_{0}$ and assuming no back arcs, and the interior arcs in each container and exterior arcs between each container forms a Graph Level Order (GLOVER) under the distance total order.
Proof.
To establish that the structure $\{C_{1},C_{2},\dots,C_{k}\}$ satisfies the GLOVER properties, we verify each condition in the GLOVER definition:
  • Total Order on Elements within $G$: - By partitioning nodes based on their distance from $v_{0}$, we establish a total order on the elements of $G$: for any two nodes $u$ and $v$, their order is determined by their respective containers. Nodes in closer containers (e.g., $C_{i}$) are considered ”less than” nodes in further containers (e.g., $C_{j}$ with $i<j$).
    - Within each container, nodes may be ordered based on additional criteria (such as their individual distances to $v_{0}$ if there are multiple paths). Thus, we achieve a total order among all nodes in $G$.
  • Leveled Structure with Containers: - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$. The elements within each container are in a total order relative to each other, and the containers themselves are ordered by distance from $v_{0}$, establishing a multi-level ranked structure.
  • Non-triviality Requirement: - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node. This fulfills the non-triviality requirement, as we avoid having all singleton containers.
  • Comparability within Containers: - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$ (for example, their shortest-path distance if there are multiple paths). Thus, comparability is ensured within containers.
  • Consistency of Comparisons Across Containers: - Since the graph is oriented and there are no back arcs (i.e., no edges from nodes in $C_{j}$ to nodes in $C_{i}$ for $j>i$), the ordering between containers is respected. This guarantees that if $C_{i}$ is ”less than” $C_{j}$, then every element in $C_{i}$ is ”less than” every element in $C_{j}$, ensuring consistency across containers.
  • Interior and Exterior Neighbors: Nodes $u$ and $v$ of the same distance from the chosen minimum degree node $v_{0}$ would be in the same container and thus be interior neighbors. If these nodes were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance different by 1.
    Each criterion for an GLOVER is satisfied by the structure defined by partitioning nodes based on distance from $v_{0}$. Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a Ranked Total Order under the distance total order, completing the proof.
What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers. This ordering is also ordering the nodes within the containers, so the GLOVER data structure is a multi-level total order.
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each container maintains its place in the total ordering.
Next, we will show the power of this data structure and why is introduced in this paper.
Lemma 5.3.
GLOVER Supports Induction: In any oriented graph, the minimum degree node $v_{0}$ within a Glover structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
Proof.
Let $v_{0}$ be the node with the minimum out-degree in the oriented graph $G$. By definition, for any node $v\in G$, $d^{+}(v_{0})\leq d^{+}(v)$, where $d^{+}(v)$ is the out-degree of node $v$. This ordering provides a baseline for comparison because all other nodes are either equal to or have greater out-degree than $v_{0}$.
The analysis begins at $v_{0}$, which serves as the foundation of the GLOVER structure. All outgoing arcs from $v_{0}$ lead to nodes at the next level (or ”container”) $C_{1}$. By the properties of the GLOVER structure (e.g., no back arcs), the relationships between nodes are strictly directed, ensuring no circular dependencies.
Assume that for any node $v$ in a container $C_{i}$, the following properties hold:
  • $d^{+}(v)\geq d^{+}(v_{0})$.
  • All properties derived from $v_{0}$ apply recursively to nodes in $C_{i}$.
Consider any node $u\in C_{i+1}$ that is connected to a node $v\in C_{i}$. By the inductive hypothesis, $v$ satisfies all conditions derived from $v_{0}$. Since $u$ receives its inbound arcs from $v$, and $d^{+}(v)\geq d^{+}(v_{0})$, the relationships in $C_{i+1}$ are consistent with the properties of the Glover structure. Furthermore, no node $u$ in $C_{i+1}$ can violate the inductive assumptions because $v_{0}$ ensures a well-ordering, propagating constraints downward.
Starting from $v_{0}$, the well-ordering propagates through all containers $C_{i}$, ensuring that:
  • The out-degree of any node is anchored to $v_{0}$.
  • There is no upward deviation in degree that contradicts the structure’s constraints.
Thus, the GLOVER structure supports induction by leveraging the minimum degree node $v_{0}$ as an anchor, ensuring consistency throughout the graph. ∎
Example 5.3.
[…]
Figure 6: This example illustrates back arcs in a graph. The two blue arcs from $x3$ to $y3$ and $y3$ to $z3$ represent exterior neighbors indicate forward directing arcs. Additionally, a red arc points backwards from $z5$ to $x1$. Although the containers are not outlined, they are implied by the difference in variables for each group.
Lemma 5.4.
Interior Degree Greater than 1 In any oriented graph $G$, represented by a GLOVER with $v_{0}$ as a minimum out-degree node satisfying the decreasing sequence property, In an attempt to prevent a degree doubling node, every node $x$ in the neighboring container $N^{+}(v_{0})$ has an interior degree of at least 1. That is, for $x\in N^{+}(v_{0})$,
$|\text{int}(v_{0},x)|\geq 1.$
Proof.
By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$
where $N^{+}(x)$ denotes the set of first neighbors (out-neighbors) of $x$, and $N^{++}(x)$ denotes the set of second neighbors (out-neighbors of $N^{+}(x)$).
For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $y$, and then from $y$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $y\in N^{+}(v_{0})$.
If $y$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $y$’s neighbors, which are $v_{0}$’s second neighbors. If $y$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $y$’s first neighbors.
To avoid this doubling, $y$ must connect to at least one interior arc. This means that the interior degree of $y$ with respect to $v_{0}$ must satisfy
$|int(v_{0},y)|\geq 1$
This argument holds for any neighbor $y\in N^{+}(v_{0})$, ensuring that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
Thus, for all $x\in N^{+}(v_{0})$,$|int(v_{0},x)|\geq 1$, completing the proof. ∎
Lemma 5.4 (Interior Degree Greater than 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling. In particular, we see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree to double.
Example 5.4.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
Lemma 5.5.
Exterior Less than Degree In any oriented graph $G$, represented by a GLOVER In an attempt to prevent a degree doubling node, if a node $x_{i}$ is in container $C_{i}$ has the decreasing sequence property, In an attempt to prevent a degree doubling node, then for all $y\in N^{+}(x)\subseteq C_{i+1}$, $|ext(x,y)|<d^{+}(x)$.
Proof.
Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|ext(x,y)|\geq d^{+}(x)$. By definition, $ext(x,y)$ consists of the set of out-neighbors of $y$, specifically:
$ext(x,y)=\{z\in N^{++}(x)\mid z\text{ is an out-neighbor of }y\}.$
Thus, $|ext(x,y)|\subseteq N^{++}(x)$, where $N^{++}(x)$ denotes the set of second out-neighbors of $x$. If $|ext(x,y)|\geq d^{+}(x)$, then:
$|N^{++}(x)|\geq|N^{+}(x)|\geq d^{+}(x),$
which suggests that the out-degree of $x$ would at least double in the graph $G^{2}$ (the second power of the graph, where edges represent paths of length 2). This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property, where the degree of nodes cannot increase as we move through the containers.
Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|ext(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
Lemma 5.5 is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree Greater than 1), with the notable exception that Lemma LABEL:extbd1 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
The results presented in 5.4(Interior Degree Greater then 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1. There, he demonstrated that a minimum degree vertex with no cycles in its induced graph has a degree that effectively doubles.
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we’re developing.
Lemma 5.6.
[…]
Then, for all $z\in ext(x,y)$, we have $|ext(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
Proof.
We begin by recalling the conditions for a node to have the decreasing sequence property (DSP). A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
Now, assume the conditions of the lemma hold: 1. The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increases significantly in $G[N^{+}(x)]$. 2. For all $z\in ext(x,y)$, $ext(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
Now, suppose for contradiction that there exists $z\in ext(x,y)$ such that $|ext(y,z)|\geq|ext(x,y)|$. This would mean that the exterior degree of $y$ is at least as large as the exterior degree of $x$. However, this contradicts the assumptions about the DSP of $y$, which requires $y$ to have fewer second neighbors than first neighbors.
From $|int(x,y)|$ doubling, $y$’s influence in $G[N^{+}(x)]$ grows, yet its exterior degree should decrease due to the DSP. The absence of back arcs prevents any increase in $y$’s exterior degree from its second neighbors.
Thus, the assumption $|ext(y,z)|\ \geq|\text{ext}(x,y)|$ leads to a contradiction,
confirming that $|\text{ext}(y,z)|<|\text{ext}(x,y)|$ for all $z\in\text{ext}(x,y)$.
Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$. ∎
The first thing to notice in this proof are the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions has an important purpose.
First we will talk about the lack of back arcs. We have already talked about how back arcs can be a hindrance to our algorithm overall. This is one of the main points where they can be a hindrance. In particular there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist. We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
Secondly, we assumed that a node’s interior degree doubled. This is similar to the notion of solving the containers separately, then joining them together again to solve the main problem. We are assuming that this not $y$ was in another container and its interior degree doubled in that container. In doing so, $y$ is now in a better position to have its overall degree double, needing only to have its exterior degree double.
Example 5.5.
0
$\Rightarrow$ 111122223333444455556666 $\Rightarrow$ 7777888899991010101011111111 $\Rightarrow$ H
Figure 7: Nodes in each container connect to all nodes in the subsequent container. Here, the minimum out-degree node has out-degree 6, with connectivity patterns shown to demonstrate degree doubling in $G^{2}$.
Table 1: Intersection Results for Triplets of Nodes in 5.5
Triplet(A, B, C) int(A, B) ext(B, C) Intersection Result
[…]
We have now defined the data structure that we will use throughout the rest of this paper. In doing so, we have a better understanding of interior and exterior arcs which will help guide us in our search for a counterexample.
6 Back Arcs
With a theoretical foundation underway, we are ready to launch an investigation into back arcs. We will look into whether these arcs can exist in a counterexample, and what other properties of oriented graphs must be true.
It is important to restate that back arcs and exterior neighbors are distinct concepts. Back arcs are not a subset of exterior neighbors. While back arcs connect to any previous container, exterior neighbors are limited to connecting with the next container. Additionally, exterior neighbors do not represent all possible external neighbors. Rather, an exterior neighbor ext(x, y) specifically refers to nodes at a distance of two from $x$ that are also neighbors of $y$.
If we look at Example 4.2 again we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$. Similarly we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
One of the things we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. This next lemma formalizes that concept.
Example 6.1.
[…]
Figure 8: Graph representation of back arcs (shown in red) and forward-directed arcs (in blue) between containers, illustrating how these arcs impact connectivity patterns in an oriented graph.
Lemma 6.1.
(Regular Interior Degrees) Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node, with minimum degree node $v_{0}$ partitioned by containers $C_{i}$. Suppose that two nodes $u$ and $v$ in a container $C_{i}$ have a common parent $v_{p}$ in the preceding container $C_{i-1}$, and that $v_{p}$’s interior degree doubles. If the interior degrees of $u$ and $v$ differ by at least 1, then the exterior degree of $v_{p}$ in container $C_{i-1}$ will double as well.
Proof.
Assume $u$ and $v$ are two nodes in container $C_{i}$, both with a common parent $v_{p}$ in container $C_{i-1}$, and let $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$, where $\text{int}(v_{p},x)$ denotes the interior degree of $x$ relative to $v_{p}$’s container $C_{i-1}$.
By the Decreasing Sequence Property (DSP), if the interior degree of $v$ is higher than $u$, then $v$ will require fewer exterior connections to maintain the required degree restrictions in the graph. Meanwhile, $u$, with fewer interior connections, will require more exterior connections to compensate.
Now, the imbalance in interior degrees means that $u$’s exterior connections are effectively increasing the degree of $v_{p}$. This is because $u$ must connect to more external nodes to maintain its degree, leading to an increase in the number of exterior neighbors of $v_{p}$.
Consequently, the exterior degree of $v_{p}$ must double to accommodate the increased number of connections needed by $u$ and $v$ in the graph, violating the degree constraints intended by the partitioning into containers.
Thus, the differing interior degrees of $u$ and $v$ in container $C_{i}$, combined with the effects of the DSP, will cause the exterior degree of $v_{p}$ in $C_{i-1}$ to double. This leads to a contradiction, so the claim holds. ∎
Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples. In particular, it shows that all the out-neighbors of interior arcs within a container must have the same degree.
In Example 5.5 and Table 1, we examined a minimum degree node $v_{0}$ with degree 6 in a regular graph. The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth.
A back arc is defined as an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$. If such a back arc were introduced, it would imply that the source node $y$ gains new neighbors at a lower level, potentially doubling its degree. This contradicts the hierarchical organization of containers, where each level is strictly defined by shortest-path distances from $v_{0}$.
Lemma 6.2.
No Back Arc to Early Container Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $ext(u_{i},v_{k})$ is at least the degree of $v_{k}$. Furthermore, if the interior degree of $u_{i}$ doubles, this will cause the total degree of $u_{i}$ to double as well.
Proof.
Assume that $u_{i}\in C_{i}$ has its first back arc to $v_{k}\in C_{k}$ where $k<i-1$. By definition of exterior neighbors:
$|ext(u_{i},v_{k})|=|N^{+}(v_{k})\cap N^{++}(u_{i})|$
where $N^{+}(v_{k})$ is the set of immediate neighbors of $v_{k}$, and $N^{++}(u_{i})$ is the set of nodes at distance two from $u_{i}$.
Since $N^{+}(v_{k})$ includes all direct neighbors of $v_{k}$, it follows that:
$|N^{+}(v_{k})|=deg(v_{k}).$
By the structure of GLOVER and the absence of back arcs from $u_{i}$ to $v_{0},\dots,v_{k-1}$, the size of $ext(u_{i},v_{k})$ satisfies:
$|ext(u_{i},v_{k})|\leq|N^{+}(v_{k})|=deg(v_{k}).$
Now, consider the relationship between interior and total degree. If the interior degree of $u_{i}$ doubles, the total degree of $u_{i}$ must also double. This is because the exterior neighbors are tightly constrained by the back arc and are not impacted by the increase in interior connections. Thus, any increase in the interior degree directly translates into a proportional increase in the total degree.
Finally, by the Weak Decreasing Exteriors Lemma, the size of $ext(u_{i},v_{k})$ decreases as we move further in the sequence of containers. This ensures that the total degree of $u_{i}$ remains bounded and behaves as expected under the given conditions, completing the proof. ∎
Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container structured graph. It has effectively reduced this case of back arcs to one of interior degree doubling or back arcs mapping to the previous container. This is important because the previous section introduced the GLOVER framework as a total order of containers. Back arcs can disrupt this order. What we see next is a special case of how back arcs could disrput this order.
Lemma 6.3.
Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
1. Suppose $x\in C_{i}$ and $y\in C_{i+1}$, both satisfying the Decreasing Sequence Property (DSP). 2. The interior degree of $y$ with respect to $x$ doubles in $G[N^{+}(x)]^{2}$. 3. There exists a $w\in\text{ext}(x,y)$ such that the $int(y,w)$ doubles.
Then, for all $z\in\text{ext}(x,y)$, we have:
$|\text{ext}(y,z)|<|\text{ext}(x,y)|$
First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
Consequently, the arc $w,z$ is an early back arc as we saw in 6.2(No Back Arc to Early Container). These arcs can lead to double counting. An out-neighbor of $w$ might be an out-neighbor of $x$, and a second out-neighbor of $y$. This implies that under the assumption that $w$’s interior degree double’s, $w$’s overall degree will double through that early back arc.
Lemma 6.4.
First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph represented by a GLOVER with minimum degree $d$ in an attempt to prevent a degree doubling node, and let $v_{0}$ be a node of minimum degree. Suppose $G$ is an oriented graph, where arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will violate the structure of the graph.
Proof.
Assume that the oriented graph $G$ is partitioned into containers $C_{1},C_{2},\dots,C_{i}$ such that nodes in each container $C_{j}$ (for $j<i$) only have either arcs to fellow nodes in $C_{j}$, or they map forward to nodes in $C_{j+1}$, with no back arcs present between containers up to $C_{i}$.
Now, let the first back arc appear from a node $u_{i}\in C_{i}$ to a node $v_{i-1}\in C_{i-1}$. In this oriented graph, by definition, all arcs are supposed to point in the forward direction—from container $C_{j}$ to container $C_{j+1}$. By Lemma 6.1 (Regular Interior Degrees), we know that all nodes inside container $C_{i-1}$ have the same interior degree. Because of their matching interior degrees, they will also share the same exterior degrees. Then since the node $v_{i-1}$ is in $C_{i-i}$, $v_{i-1}$ also maps to these exterior nodes. If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist{v_{0},u_{i}}-i$ so $u_{i}$ cannot belong to container $C_{i}$.
Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ leads to a fundamental contradiction in the oriented nature of the graph, and no back arcs can exist prior to $C_{i}$. ∎
Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees). From this, we can see that a back arc to a parent container violates the orientated nature of the graph.
Lemma 6.5.
No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$,$C_{2}$, $\dots$,$C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
Proof.
This was proved in the Lemmas 6.2 (No Back Arc to Early Container and 6.4 (First Back Arc Between Consecutive Containers) . ∎
[…]
Figure 9: Illustration of the proof of 7.2 (Interior Degrees Double). Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.
[…]
We stated that back arcs could be detrimental to our algorithm, so the removal of them in this section was important. Now we can better navigate the other properties of oriented graphs.
7 Interior Degrees Doubling
This section introduces the concept of interior degree doubling. This is similar to solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s the interior degree inside of a container. This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
Lemma 7.1.
Indistinguishable Regular Interior Nodes In a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER in an attempt to prevent a degree doubling node. If a node $v_{p}$’s interior degree doubles then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree double.
Proof.
Let $C_{i}$ be a regular container of size $n$, and let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$. The interior degree of a node $v_{i}$ in container is defined as the number of arcs shared $v_{i}$ between $v_{i}$ and a neighbor or parent node. In a regular container, all nodes are indistinguishable in terms of their degree. They also have the same parents because of the DSP. Any automorhism we construct including a random automorphism $\phi$ of the container $C_{i}$ will preserve the degree structures. This means any node will have its interior degree double. ∎
The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others. Previously we were searching for a node in a container whose interior degree had doubled. That would be the key. What indistinguishabable tells us is that if one node’s interior degree doubles inside a container, then all the node’s interior degrees double. That means that we can select any of those nodes.
Lemma 7.2.
Interior Degrees Double Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have its interior degree doubled.
Proof.
We prove this by strong induction on the size of the graph.
Base Case: In Lemmas 4.1 (Minimum Out-Degree ¡ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4.3 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3 there exists at least one node in each container with a doubled interior degree. These cases form our base, supporting the induction.
Inductive Hypothesis: Assume that for any graph of size $n^{\prime}<n$, each container $C_{i}$ contains at least one node whose interior degree doubles, thus satisfying the DSP by proportionally reducing exterior degrees.
Inductive Step: Consider a graph of size $n$ where we need to show that at least one node in each container $C_{i}$ has its interior degree doubled.
Suppose we construct $G$ such that no node at distance $i$ from $v_{0}$ initially has a doubled interior degree in $C_{i}$. Define a sequence of subgraphs $H_{0}$,$H_{1}$,…,$H_{k}$ for $C_{i}$ as follows:
$H_{0}=\{u_{i}\}$
$H_{1}=N^{+}(u_{i})$
with nodes in $H_{1}$ having doubled interior degrees by the I.H., as applied to each subgraph. By the Induction Hypothesis (I.H.),there exists a node $y_{1}\in H_{1}$ such that the interior degree of $y_{1}$ with respect to $u_{i}$ doubles in $H_{1}$. We can now define $H_{2}=ext(u_{i},y_{1})$. Likewise, by the I.H., there will be a node $y_{2}inH_{2}$ whose interior degree doubles in $H_{2}$. We can define $H_{3}=ext(y_{1},y_{2})$. Continuing this process, we can define $H_{j}=ext(y_{j-2},y_{j-1})$, where $y_{j-1}$ is a node whose interior degree doubles in $H_{j-1}$.
By the Weak Exterior Size Decreasing lemma, we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers. By the DSP, an increase in $int(x,y)$ requires a proportional reduction in $ext(y,z)$, preventing exterior degrees from exceeding interior degrees. This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$, as required by DSP.
The strictly decreasing nature of this sequence implies that it must terminate when $H_{i}\leq 2$. At this point, the remaining nodes in $H_{i}$ either have no further exterior neighbors or their degrees lie entirely within the interior set.
Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$. Thus, the parent of $y^{*}$, denoted $y^{*}-1$, in $G$ achieves a doubled degree within $C_{i}$.
This doubling process terminates as DSP constraints prevent exterior degrees from surpassing interior degrees without contradicting the DSP.
Therefore, for a graph of size $n$, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step. ∎
Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. By finding a positive answer to this, and not just a positive answer, but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. Once again, we see why the Decreasing Sequence Property (DSP) received its name.
This doubling of interior degrees will help us with the construction of our algorithm in search for a counterexample. It does, however, give us a bit of light into the side of the Conjecture being true.
8 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The Weak Exteriors Decreasing Lemma implies that containers will decrease in size. We need to see how that decrease in size will impact the degrees of the nodes.
Lemma 8.1.
Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in $G$ has the Decreasing Sequence Property (DSP). Then the regular interior degrees will result in regular exterior degrees in $G$.
Proof.
Let $x_{i}$ and $x_{j}$ be two nodes at the same distance $i$ from the minimum degree node $v_{0}$ in the oriented graph $G$, both contained in the same container $C_{i}$.
Assume, for contradiction, that the sizes of their exterior sets differ by at least one. Without loss of generality, let there exist a node $y$ such that $|ext(x_{i},y)|=|ext(x_{j},y)|+1$. This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
Since the size of the next container $C_{i+1}$ is determined by the exterior degrees of nodes in $C_{i}$, this irregularity in the exterior degrees would force the size of $C_{i+1}$ to differ, contradicting the assumption that the containers are of regular size.
By Lemmas 6.1 (Regular Interior Degrees) and 7.2 (Interior Degree Double), we know that if the interior degrees within a container are regular, the exterior degrees must also follow this regularity. Specifically, if there is an irregularity, any parent node of $x_{i}$ and $x_{j}$ will have its exterior degree double, which would in turn double their total degree.
This doubling of the parents’ total degree would violate the Decreasing Sequence Property (DSP).
Therefore, the assumption that $|ext(x_{i},y)|$ and $|ext(x_{j},y)|$ differ is false. Hence, the exterior degrees in $G$ must be regular if the interior degrees are regular. ∎
We proved Lemma 6.1 (Regular Interior Degrees) on our way to proving that back arcs did not exist. LABEL:refintdeg is an important property of the graph though that we are going to use to get other properties of the graph, like 8.1.
Lemma 8.2.
Regular Graph All Nodes Have Equal Our-Degree Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node, with $v_{0}$ as a minimum degree node. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $d_{v_{0}}=d^{+}(v_{0})$.
Proof.
We will prove this lemma by induction on the number of nodes in the graph $G$.
For the base case, consider the node $v_{0}$, which, by definition, has the minimum degree $d$, so $d^{+}(v_{0})=d_{v_{0}}$.
Assume that for some node $v_{k}$ in $G$, the degree of $v_{k}$ satisfies $d^{+}(v_{k})=d_{v_{0}}$.
Now consider an arbitrary node $v_{k+1}$ in $G$. Suppose, for contradiction, that $d^{+}(v_{k+1})>d_{v_{0}}$.
Consider a neighbor $v_{n}$ of $v_{k+1}$. By the inductive hypothesis, the degree of $v_{n}$ is exactly $d_{v_{0}}$. Since $v_{k+1}$ has more than $d_{v_{0}}$ neighbors, the neighbors of $v_{k+1}$ are second neighbors of $v_{n}$ in the graph $G^{2}$. Therefore, $v_{n}$ would have more than $d_{v_{0}}$ second neighbors in $G^{2}$.
This implies that $v_{n}$’s degree has effectively doubled, which contradicts the DSP, which requires that $v_{n}$ must have degree $d_{v_{0}}$. Hence, the assumption that $d^{+}(v_{k+1})>d_{v_{0}}$ leads to a contradiction.
Thus, we must conclude that:
$d^{+}(v_{k+1})=d_{v_{0}}.$
By the principle of mathematical induction, it follows that for every node $v$ in $G$, the degree of $v$ is equal to $d_{v_{0}}$. ∎
We are incrementally refining the properties that must hold for our counterexamples. The two most crucial aspects are the Decreasing Sequence Property and the identification of the minimum degree node. However, additional constraints, such as the overall graph regularity established by Lemma 8.2 (Regular Graph), help simplify future arguments. This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
Lemma 8.3.
Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER In an attempt to prevent a degree doubling node, with no back arcs, and let $v_{0}$ be the chosen minimum degree node in $G$. For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1)$
for some $z\in N^{+}(x_{i})$ in order to keep any node’s degree from doubling in the oriented graph.
Proof.
We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$.
Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
$\{y\in\text{ext}(x_{1},z)\text{ s.t. }dist(v_{0},y)=2\}$
is bounded by $d^{+}(v_{0})-1$, since each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2, as there are no back arcs to $v_{0}$.
Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1).$
Inductive Step: We need to show that this bound holds for nodes at distance $i+1$ from $v_{0}$. Let $x_{i+1}$ be a node in container $C_{i+1}$ relative to $v_{0}$. Suppose, for contradiction, that
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|>d^{+}(v_{0})-i.$
This would imply that the exterior neighbors of $x_{i+1}$ exceed the allowed bound, violating the decreasing sequence property (DSP) conditions of $x_{i}$. By Lemmas 7.2 (Interior Degrees Double), 6.1(Regular Interior Degree), and 7.1 (Indistinguishable Regular Interior Nodes), this would cause the interior degree of $x_{i}$ to double, thereby forcing an overall doubling of the degree for $x_{i}$.
This contradiction implies that our assumption must be false. Therefore, we conclude:
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|\leq d^{+}(v_{0})-i.$
Lemma (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma), to give us a way of calculating the maximum size of a container based only in its distance from the minimum degree node and the minimum degree node’s degree.
The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” is named.
Lemma 8.4.
Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER, In an attempt to prevent a degree doubling node, and let $v_{0}$ be a minimum degree node in $G$, where every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided $(v_{0},x)\in G$.
Proof.
Let $x$ be a node at distance $i$ from $v_{0}$. By Lemma 8.3 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$, $|ext(v_{i-1},x)|$, is bounded by $d^{+}(v_{0})-i$. This places a limit on the number of nodes $x$ can map to at distance $i+1$ from $v_{0}$.
Now, since $x$ has at least the same out-degree as $v_{0}$ (by the assumption of DSP), $x$ must map to at least $i$ other nodes. Given that back arcs are not allowed (as shown previously), $x$ cannot map to nodes at earlier distances without violating the structure of the graph.
Therefore, the only remaining possibility is that $x$ maps to interior nodes. Since $x$’s degree must account for at least $i$ mappings, and these must be to interior nodes, we conclude that $x$ must map to at least $i$ interior nodes.
Thus, $|int(v_{i-1},x)|\geq i$, as required. ∎
Having already placed an upper bound on the exterior sizes and containers, it is only natural that the there is an opposite bound for the interiors. Once again the minimum degree node is the driving force behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
9 Transitive Triangles and Seymour Diamonds
Before we continue with the development of our DSP algorithm and show how it solves the Seymour Second Neighborhood Conjecture, we need to address some potential problems and show that they will not hinder our algorithm.
Definition 9.1.
Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
Here, we have $x$ pointing to $y$ and $u$; and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles we might improperly count out second neighbors and thus not have a correct algorithm.
Definition 9.2.
Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
The Seymour Diamond is a graph theory concept in graph there where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$. This is basically a four vertex oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and more importantly the Seymour’s Second Neighborhood Conjecture.
The notion of interior neighbors is another way of describing what [1] referred to as transitive triangles. They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e. one where every node has the decreasing sequence property. We provided a similar result with Lemma 5.4 (Interior Degree Greater than 1). When a minimum degree node $v_{0}$ has the decreasing sequence property, all of its neighbors have interior degree of at least 1. That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$ will form a transitive triangle. We can extend that further for any two constructive containers.
xyu
Lemma 9.1.
Transitive Triangles from Containers In a graph $G$ represented by a GLOVER, for any two consecutive containers $C_{i}$ and $C_{i+1}$ in a graph satisfying the Decreasing Sequence Property (DSP), there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle. That is, there exist directed edges from $u$ to $w$, $v$ to $w$, and from $u$ to $v$, or from $v$ to $u$.
Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm.
Lemma 9.2.
Seymour Diamonds from Containers
In a graph $G$ represented by a GLOVER, for any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
$u\in C_{i},\quad v,w\in C_{i+1},\quad y\in C_{i+2}$
The following arcs hold:
1. There are forward arcs $x\to y$ and $y\to u$, 2. There are forward arcs $x\to w$ and $w\to u$, 3. No back arcs exist between containers.
Similar to Lemma 9.1, Lemma 9.2 (Seymour Diamonds from Containers) lets us know that we have been interacting with Seymour Diamonds all along without knowing it. We didn’t know it because they they did not disrupt our algorithm and they are not a hindrance to our algorithm.
xywu
Lemma 9.3.
Seymour Diamonds and Decreasing Path Algorithm Let $G$ be an oriented graph represented by the GLOVER, in an attempt to prevent a degree during node, where we are applying the DSP path algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP path algorithm.
Proof.
  • A Seymour diamond consists of two vertices, $x$ and $u$, connected by two parallel paths via $y$ and $w$.
  • The DSP path algorithm can utilize one of these paths while disregarding the other, as both connect $x$ and $u$, thus maintaining degree constraints.
  • The presence of redundant paths does not diminish the outgoing neighbors of any node, ensuring that the minimum degree condition holds throughout the process.
  • Consequently, these additional connections offer alternative routes without introducing cycles or otherwise disrupting the algorithm’s operation.
None of these case are a deterrent to the execution of the DSP Path algorithm. ∎
Lemma 9.4.
Interaction of Minimum Degree Nodes and Transitive Triangles Let $G$ be an oriented graph represented by a GLOVER, containing a minimum degree node $v_{0}$ with degree $d$. Consider a transitive triangle formed by nodes $z$, $y$, and another vertex $x$, where $z$ and $y$ are part of this triangle. The following cases are possible:
  • If $d^{+}(y)=d$ and $d^{+}g(z)=d$
  • If $d^{+}(y)=d$ and $d^{+}(z)>d$
  • If $d^{+}(y)>d$
  • If $d^{+}(y)<d$
Proof.
  • Case 1: If $d^{+}(y)=d$ and $d^{+}(z)=d$: Both $y$ and $z$ have exactly $d$ outgoing connections. For the triangle to be transitive, they must not share all neighbors, meaning there exists at least one neighbor of $y$ that is not a neighbor of $z$. Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
  • Case 2: If $d^{+}(y)=d$ and $d^{+}(z)>d$: $z$ covers some neighbors of $y$, but since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
  • Case 3: If $d^{+}(y)>d$: $y$ has more outgoing connections than required. It can utilize these additional connections to map to unique neighbors without impacting the minimum degree condition, ensuring $y$ retains neighbors not covered by $z$.
  • Case 4: If $d^{+}(y)<d$: This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections. Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
Now that we see that transitive triangles and Seymour diamonds are not a problem for our algorithm, we can continue our investigation of oriented graphs. Next we will blend everything together into our main results.
10 Main Theorem: Seymour’s Conjecture is True
10.1 Introduction
Here is where we sum up all the lemmas we have derived from our analysis on oriented graphs. We will state the final algorithm and some final results.
So far, this battle has been a battle between a catalyst - the chosen minimum degree node, and the bottleneck - the Decreasing Sequence Property (DSP). We have proved some important things - Back arcs cannot exist; interior degrees must double inside each container, the exterior degrees decrease in size; interior degrees proportionally increase in size.
There is a limit though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
We first restate the Lemma 6.5 (Non-Existence of Back-Arcs), from section 6, but slightly rephrased.
Lemma 10.1.
Let $G$ be an oriented graph represented by a GLOVER, in an attempt to prevent a degree doubling node, with all nodes satisfying the Decreasing Sequence Property (DSP), where $v_{0}$ is a minimum degree node in $G$, and all nodes are placed into containers. If there are an back arcs in $G$, then a node’s degree will double.
10.2 Theorem Statement
Degree Doubling by Exterior Decreasing
Theorem 10.1.
Let $G$ be an oriented graph represented by a GLOVER with all nodes the Decreasing Sequence Property (DSP). If all nodes in $G$ are placed into containers, where interior degrees are maximized, then Seymour’s conjecture holds true. Specifically, a node in $G$ will have its exterior size decrease to less than 2.
Algorithm 1 Algorithm for Analyzing Degree Doubling in Containers
1:Initialization:
2:Choose the minimum degree node $v_{0}$.
3:Divide nodes into containers $C_{1},C_{2},\ldots$ based on their degree and proximity to $v_{0}$.
4:for each container $C_{i}$ in $G$ do:
5:     Locate a node $x_{i}$ in $C_{i}$ whose interior degree doubles.
6:     Move to a neighboring container $C_{i+1}$.
7:     if there is a back arc from $C_{i+1}$ to a prior container then
8:         Locate child of $x_{i}$ whose exterior degree doubles, causing $x_{i}$’s overall degree to double.
9:         Halt.
10:     else
11:         Since there is no back arc, the container $C_{i+1}$ will be bound in size to be 1 less than the size of container $C_{i}$.
12:         Repeat from Step 4.
13:     end if
14:end for
Start with graph size $n$Select minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots$Identify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to prior containerLocate child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
Figure 10: Flowchart depicting the degree-doubling algorithm applied within containers, demonstrating the process of identifying nodes with the Decreasing Sequence Property.
Corollary 10.1.
Degree Doubling by Container Density In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},C_{2},...,C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
There are two possibilities:
  1. 1.
    As the containers shrink in size and the interior degrees grow in size, we reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
  2. 2.
    For small minimum degree nodes $v_{0}$, The container’s shrinkage is not fast enough to cause this collision.
Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and increased number of nodes. This was represented by the yin and yang effects early on. We see it full circle here. Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbor neighbors represented the size of the containers.
In small graphs like 4.3 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur. To understand the reasoning behind this, understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1. The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$. So for small graphs, we are just not going out far enough to catch up. For larger graphs though, we do start to see this.
Corollary 10.2.
[…]
$i=\frac{k-1}{2}$
where $k$ is the minimum degree of $G$. For even values of $k$, the complete graph $K_{n}$ will not occur on any actual even number of nodes but may appear in the interval between two odd numbers.
Proof.
We begin by restating some previously shown examples:
  • For $k=3$: $K_{3}$ (triangle) occurs in $C_{1}$.
  • For $k=5$: $K_{5}$ occurs in $C_{2}$.
  • For $k=7$: $K_{7}$ occurs in $C_{3}$.
These observations hold because the GLOVER structure ensures that each container $C_{i}$ increases in regularity and density as the degrees of nodes double, creating the conditions necessary for complete graphs to emerge.
For even values of $k$, such as $k=4$ or $k=6$, the symmetry and density required for a complete graph $K_{n}$ cannot form directly at any container level. This is because the degree-doubling effect and the odd-number dependency of $K_{2i+1}$ mean that the requirements for even $n$ are never satisfied in full. Instead, it transitions through intermediate states between odd complete graphs, and complete graphs may appear to ”exist” in a fractional or approximate form between $K_{2i-1}$ and $K_{2i+1}$.
The relationship $i=\frac{k-1}{2}$ reflects the interplay between the minimum degree $k$ and the growth of density in containers. This relationship confirms that the GLOVER structure consistently generates complete graphs $K_{2i+1}$ in odd-numbered increments, with no direct occurrence for even $n$.
This emergence of orientations of complete graphs as containers decrease in size adds a layer of beauty to the Seymour Second Neighborhood Conjecture. There was never a question of these arising. Some questioned multiple nodes. Some questioned zero nodes. A complete graph occurrence though is a thing of beauty and it helps give rise to another result.
Corollary 10.3.
Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $k$, and let the complete graph occur in container $C_{i}$. Then every node in container $C_{i}$ has its degree doubled in the representation.
The discovery of multiple degree-doubling nodes lifts the Seymour Second Neighborhood Conjecture from a mere question of existence to one in abundance. There were many who doubted that one even existed. Now there are many.
Second is the way we find it. It’s the presence of an oriented complete graph as these containers decrease in size. There’s an element of mathematical beauty there.
Then third we can look at the applications. We may be looking for the person who has the most influence, but this is saying that there may not be just one person there may be a whole set of people, a tightly connected group.
10.3 Decreasing Sequence Property Path Algorithm
We will now give a pseudocode algorithm to identify the node whose out-degree doubles in the square of an oriented graph. This is the culmination of the lemmas, theorems, and corollaries discussed throughout the paper, providing a cohesive framework for finding degree-doubling nodes.
  • The containers reach a size two causing a node in the previous container to have its degree double, as stated in Theorem 10.2 (Degree Doubling by Exterior Decreasing).
  • The interior degree required for nodes inside containers will be greater than the possible for a container of that size causing a node in the previous container to have its degree double, as stated in Corollary 10.1
In both these cases we are able to find a node whose degree doubles.
Here is the algorithm in pseudocode. Input: An oriented graph $G$ with minimum out-degree $d$.
Output: A node $v$ in $G^{2}$ whose out-degree at least doubles.
Algorithm:
1. Initialize the Graph Structure: - Identify the minimum degree node $v_{0}$ in $G$. - Partition $G$ into containers $C_{1},C_{2},\ldots,C_{d}$, where each container $C_{i}$ groups nodes by their distance $i$ from $v_{0}$.
2. Check for Back Arcs: - For each container $C_{i}$, check for arcs leading to any container $C_{j}$ where $j<i$. - If a back arc exists from a node $u_{i}\in C_{i}$ to a node $v_{j}\in C_{j}$, record $u_{i}$ as a degree-doubling node by Lemma 7.2 (Interior Degree Doubling) and Lemma 6.5 (Non-Existence of Back Arcs). - If no back arcs exist, proceed with Step 3.
3. Establish Connectivity Properties: - For each node $v$ in container $C_{i}$, define: - Interior neighbors $\text{int}(v)$: Nodes within the same container. - Exterior neighbors $\text{ext}(v)$: Nodes in the next container.
4. Iterative Degree-Doubling Check: - For each container $C_{i}$: - Calculate the out-degree of each node in $C_{i}$. - If a node in $C_{i}$ has fewer second neighbors than first neighbors (DSP violation), record it as a potential degree-doubling node.
5. Decreasing Sequence Property Verification: - Verify that the size of each exterior set $\text{ext}(v)$ is strictly less than the size of the corresponding interior set $\text{int}(v)$ in $G^{2}$. - If not, apply the conditions set in Lemma 7.2 to ensure any doubling in interior degrees results in a matching exterior degree.
6. Termination and Output: - Return the identified node(s) with doubled degrees as proof of degree-doubling in $G^{2}$.
11 Conclusion and Future Work
In this work, we explored the Seymour Second Neighborhood Conjecture in oriented graphs. We did so through the lenses of graph theory, combinatorics, algorithms and data structures. This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide and conquer paradigm.
One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph. This effectively solves the Seymour Second Neighborhood Conjecture with an answer in the affirmative. The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree doubling node was in the graph. Instead the algorithm selects a minimum degree node and is able to partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
This approach introduced a new perspective on the Seymour Second Neighborhood Conjecture. It was no longer framed as merely a graph theory problem. By utilizing the divide and conquer paradigm and solving different parts independently, this algorithm shows that the Seymour Second Neighborhood Conjecture could effective be treated with computer science techniques. This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists. Alternatively, it may be that there is not enough interdisciplinary communication about these problems and the potential techniques to solve them.
The Seymour Second Neighborhood Conjecture started as a question in 1990. Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers. It has been talked about in books. on message boards, in classrooms, and at conferences. This problem has a long history.
With a solution now in hand, its time to begin looking forward. The Seymour Second Neighborhood Conjecture can be represented as a social media problem instead of an oriented graph. Here, the nodes are people, where no one follows back. Now instead of wondering if there is a person whose sphere of influence doubles, we would simply be searching for that person. We just need to implement the algorithm to find the person. We could have the graph represent epidemiology. Here the nodes be infected people, and the arcs would be one way infections. The question of interest would be can we implement the algorithm to find the person who is a super-spreader, and do so quickly? There are many more real life examples where we can apply this problem and this algorithm. There are fields like security, telecommunications, urban planning, neural networks. There are too many to name in this short paper.
Secondly, we introduced a data structure, Graph Level Order-GLOVER. This data structure allowed us to get a lot done on the conjecture. Many of the things inherent to SSNC made it easier for GLOVER - the well ordering and the dual metrics of distance and degrees.
Future research and development should explore questions of complexity, scalability. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
12 Acknowledgments
I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:6 has been a guiding light throughout this journey. I am profoundly grateful to my family—my mother, father, and wife—for their unwavering support and encouragement.
Special thanks to the late Dr. Nate Dean for his contributions to this work and to Dr. Michael Ball for his support. I want to sincerely thank everyone who has taken the time to pray for me, share encouraging words, and offer constructive feedback. Your support has meant so much to me, and I’m truly grateful for the kindness and thoughtfulness you’ve shown.
Finally, I extend heartfelt gratitude to my family and friends for their patience and belief in me throughout this journey.
The GLOVER data structure is protected by a provisional patent application filed with the U.S. Patent and Trademark Office.
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Ranked Structure: The elements of $S$ are partitioned into a ranking of ”containers” or ”levels,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Definition 13.4.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the interior neighbors of $y$ with respect to $x$, denoted as $int(x,y)$, are the out-neighbors of $y$ that are also out-neighbors of $x$. Formally,
$int(x,y)=N^{+}(x)\cap N^{+}(y),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$ in the graph. The number of interior neighbors of $y$ with respect to $x$ is called the interior degree of $y$.
Definition 13.5.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the exterior neighbors of $y$ with respect to $x$, denoted as $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$. Formally,
$ext(x,y)=N^{+}(y)\cap N^{++}(x),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$, and $N^{++}(x)$ denotes the set of out-neighbors of the out-neighbors of $x$ (i.e., the second neighbors of $x$). The number of exterior neighbors of $y$ with respect to $x$ is called the exterior degree of $y$.
Definition 13.6.
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
Definition: Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
Definition 13.7.
Definition: Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
14 Appendix 2: Applicability of the Graph Level Order
The deciding factor that held the algorithm together and helped solve the Seymour Second Neighborhood Conjecture was the Graph Level Order (GLOVER) data structure. This research did not begin as an investigation into data structures. Instead, a series of operations were conducted on the oriented graphs - partitioning, ordering those partitions, adding interior and exterior arcs, until what was left was an ordering of both the partition and nodes within the partition. Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric. Unlike the Seymour Second Neighborhood Conjecture, exterior arcs do not always have to only be defined to the next container. That was a problem specific definition.
Moving forward, the GLOVER data structure should be well suited to tackle problems in graph theory and beyond. By introducing containers, another way of doing induction on graphs has been introduced. The concept of divide and conquer, which has dominated much of computer science, shows up in this paper as well. We are able to treat these containers independent of one another and solve their problems locally before bringing them back to the global problem. This shows the how versatile the GLOVER data structure is moving forward.
What makes GLOVER special are things like total order on containers. Now there is a two way ranking of nodes, as opposed to traditional one way rankings like lexicographical sorting. This agrees with many real world systems where we have dual metrics, that are often competing, like price and performance. This is hard to measure on a single scale.
Next these nodes are not just treated like clusters inside these containers, but like nodes in graph theory. The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth first search, or centrality to be called. What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
This is still a new data structure, so we are learning a lot about it. There are many applications that have not been tried yet. However, since it brings the strengths of a total order and a graph into one unified structure, it should be able to model complex relationships. This should include domains of machine learning like clustering, dynamic scheduling, resource management, project management, and natural language processing.
15 Appendix 3: Visualization and Accessibility of Oriented Graphs
When we take a course on graph theory, oriented graphs are probably not on the course load. Second neighbors or squares of graphs are probably not either. To fully grasp this problem though we need to be able to delve into these definitions through visualizations.
For example, research on the Seymour Second Neighborhood Conjecture led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs. The paper addressed how these terms were intertwined with one another. One example is where the paper says that there is a monotonically decreasing sequence.
James Robert Brown’s book Philosophy of Mathematics speaks of the importance of visualizations in mathematics. Proofs are still the foundation, but sometimes visualizations can serve as an inspiration or a spark for those proofs.
This realization motivated the creation of a complimentary website. The site serves as an interactive tool to understand research. It also has things like mathjax installed to read lemmas, definitions and proofs from this paper. Where possible, things like HTML canvas and JavaScript D3.js told are used to visualize these concepts and help improve understanding.
Readers are encouraged to explore the website for additional insights and examples at:
https://glovermethod.com
This is an evolving project. Additional examples and applications will continuously be added
Readers are invited to explore the website. Please provide feedback on any aspects of the site or the content for improvements.
Our goal here was to help users connect with the Seymour Second Neighborhood Conjecture. Hopefully, this tool not only serves as support for this paper, but as an educational resource in the fields of graph theory and computer science. Thank you for your engagement