### v6

# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. This paper first presents a data structure, GLOVER (Graph Level Order), which is a partitioning of the nodes into a total order of neighborhoods. This data structure allows us to introduce a total order on oriented graphs and construct a decreasing sequence of subsets. This decreasing sequence of subsets is what ensures that no counterexample can exist. Further, this algorithmic approach lets us know where the node is. Later findings extend the results to multiple nodes. This is true for all oriented graphs. These findings also have practical applications in network optimization and algorithm design.
1 Introduction
Seymour’s Second Neighborhood Conjecture (SSNC) was originally proposed by Paul Seymour in 1990. It asks the question of the existence of a node in an oriented graph whose out-degree at least doubles in the graph’s square. The conjecture was first published by Nathaniel Dean and Brenda J. Latka in 1995 [4], where they posed a similar conjecture specific to tournaments. This version of the conjecture was first proven by Fischer [8], with a second proof provided by Havet and Thomasse [10].
Conjecture 1.1.
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}$ represent the first and second neighborhoods of $v$ respectively.
This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [9]). Here, we hope to tackle the entire class of oriented graphs based on the structure laid out in that paper.
The Seymour Second Neighborhood Challenge is a difficult graph theory problem. The problem itself asks us to focus on the question of oriented graphs and second neighbors. There have been several papers that have looked into special cases. Kaneko and Locke [11] showed that the conjecture was true for graphs with a minimum degree of at most six. Brantner et al. [1] showed that, if the conjecture is false, infinitely many counterexamples must exist. Additionally, Chen et al. [3] provided a lower bound for how much a node’s out-degree increases in the graph’s square. They concluded that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Bolter et al. [2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the Seymour Second Neighborhood Conjecture.
The topic of $m$-free graphs is of particular interest to this paper because Daamouch’s [7] work is very similar to what we see here. In his paper, he introduces induced subgraphs and speaks of their roles in degree properties on vertices, which will also be a key thing in this paper.
The Seymour Second Neighborhood Conjecture is not strictly a graph theory problem. Although it has historically been studied within the field of mathematics, it has practical applications across several domains, including social media, social networking, epidemiology, disease control, and many other fields. In the critical context of disease control, the nodes in this conjecture can represent individuals, and the orientation of the arcs symbolizes one person infecting another. Here, the application of SSNC becomes relevant by asking: Is there a person whose influence in their second radius of interaction (i.e., people they indirectly connect with) exceeds the number of people in their first radius (i.e., direct contacts)?
In this paper, we will prove the Seymour Second Neighborhood Conjecture is true. We will do this by first assuming that it is false and attempting to construct a counterexample. We will explore the conditions under which no node’s out-degree doubles will double in the square of the oriented graph. This approach will focus initially on minimum out-degree nodes. We will analyze how their out-degree could fail to double. We will then extend our algorithm to other nodes based on their distance from the minimum out-degree node.
The paper is organized as follows: Section 2 begins with some common terminology used in graph theory. Section 3 introduces our approach to the problem. In section 4 we look at some small examples of this style of thinking. Section 5 introduces the terminology and identifies the conditions under which their out-degree might fail to double. Section 6 examines the back arcs and their place in oriented graphs. Section 7 looks at the conditions that would cause an interior degree to double and its importance to the Seymour Second Neighborhood Conjecture. Section 8 further looks at containers under the context of density constraints. In Section 9, we consider transitive triangles and Seymour diamonds and show that they are not a hindrance to our algorithm. Section 10 presents this process, the full Decreasing Sequence Property algorithm, as well as the theorem stating its solution to the Conjecture.
2 Graph Theory Terminology
A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $(x,x)$ where $x$ is a node in $G$) and no symmetric arcs, that is, no arcs of the form $(x,y)$ and $(y,x)$ where $x$ and $y$ are nodes in $G$. The square of a graph $G$, denoted $G^{2}$, consists of the same node set as $G$, and an arc exists from node $x$ to node $z$ if there is a node $y$ such that both arcs $(x,y)$ and $(y,z)$ exist in $G$ as well as the arcs already in $G$. Essentially, $G^{2}$ captures two-step connections from $x$ to $z$.
Definition 2.1.
Let $G^{2}=(V,E^{2})$ where $G=(V,E)$ is the original graph, and $E^{2}$ is the set of edges defined as:
$E^{2}={(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E}$
We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$. These are the nodes that $v$ has arcs pointing to. Similarly, $N^{++}(v)$ refers to the second neighbors of $v$, that is, nodes that are two steps away from $v$ in $G$.
Let $G=(V,E)$ be a graph and let $S\subseteq V$ be any subset of vertices. The induced subgraph $G[S]$ is the graph whose vertices are the set $S$ and whose edge set consists of all edges in $E$ that have both endpoints in $S$. For any node $v$, we refer to the neighbor-induced subgraph on $N^{+}(v)$, consisting of $v$’s out-neighbors. Let $u,v$ be nodes in $G$. The distance between $dist(u,v)$ is $k$ if the shortest path from $u$ to $v$ has length $k$.
Example 2.1.
[…]
Figure 1: A five-vertex graph illustrating node out-degrees and distances. Out-degrees are labeled near each node, and the dashed rectangle indicates an induced subgraph on vertices A, B, and C. Previously used in [9]
Conjecture 2.1.
(Square version of Seymour’s Second Neighborhood Conjecture). Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when you square the oriented graph.
Now that we have two different conjectures, we can investigate how they relate to one another. Conjecture 1.1 relates the notion of a node’s degree doubling to its neighborhoods, while conjecture 2.1 relates this same concept to the square graph. The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling, and its second neighborhood being greater than or equal to the first. By the definition of an arc in the square graph, we can see that an arc is defined in this square graph only for paths of length two in the original graph. This is why these two conjectures are seen as identical.
We will focus our attention on Conjecture 1.1 for the reason that the conjecture 2.1 requires two different graphs - the original graph and the square graph. While the construction of the square graph is possible, we have found that it is not as helpful to our research at this time. Instead, focusing on the neighborhood analysis brought many insights and characteristics that helped to develop our algorithm.
Next, we turn our attention to the term degree-doubling nodes. By this term, we aim to get to the heart of conjecture 1.1. That conjecture is looking for a node whose degree at least doubles, and we are doing so by looking at the first and second neighborhoods of nodes and finding one whose second neighborhood is at least as large as the first, as the conjecture asks for. A similar definition, a Seymour vertex, fulfills conjecture 2.1 - it’s out-degree in the squared graph is at least double its original out-degree. We could equivalently search for a Seymour vertex or a degree-doubling node.
3 Approach and Motivation
We begin our analysis by considering general oriented graphs. We make no initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs. We know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
From this foundation, we can now partition the graph into $d_{v_{0}}+1$ subgraphs corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm. No nodes are being removed to form these partitions. This is a conceptual idea. Instead, think of it as putting a box around the partition. Now the nodes inside the box can work with the other nodes inside the box and solve their problem independent of the other partitions. These partitions are still connected to one another, but thinking of them in this way allows us to do more.
This represents a stark contrast to most previous approaches to the Seymour Second Neighborhood Conjecture. Many of these previous approaches were conceptual or theoretical. They would say only that a node whose degree doubled existed, but the notion of a path to that node was not part of the paper. Secondly, most of these papers focused on local search techniques that started at or near the desired degree-doubling node. That presents a problem because oriented graphs can be varying in nature. This is the first exhaustive search technique on the SSNC problem.
When speaking of an exhaustive search algorithm, complexity becomes a factor and questions of how the algorithms will handle large graphs compared to smaller graphs. This should not be a problem as we will see in the following sections; the slowest part of this algorithm is building the data structure. That will require partitioning the nodes and edges of the graph, in a worst case complexity of $O(V+E)$.
By partitioning the graph, even larger graphs, often with higher minimum degree nodes and thus more partitions, are able to be broken into sub-problems and solved independent of one another. These subproblems are each their own instances of the Seymour Second Neighborhood Conjecture, just a localized version. What the partition is doing is breaking a node’s individual degree into two parts-the degree within the container and the degree to the next container. The degree within the container is the subproblem that will be solved separately, and the degree to the next container is how they connect back to the global problem.
We still need to show that this is a formidable foundation. We have laid the groundwork for our algorithm to tackle the oriented graphs. With a minimum degree node, we can partition this graph into independent subproblems. This is now ready for an exhaustive search algorithm.
4 Initial Lemmas
These first lemmas will seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. For this section, unless otherwise noted, assume that $v_{0}$ is a node with minimum out-degree in our oriented graph $G$. We will also introduce some standard notation from graph theory. The work in this section agrees with what was already done by [11] and is not meant to serve as original content. It does, however, provide insight into our reasoning.
Example 4.1.
$v_{0}$$v_{2}$$v_{3}$$v_{1}$$v_{6}$$v_{5}$$v_{4}$
$\Rightarrow H$
Figure 2: Example illustrating how a minimum out-degree node in $G$ has its degree doubled in $G^{2}$ under certain conditions.
Lemma 4.1.
Minimum Out-Degree $<$ 3 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double in $G^{2}$.
Proof.
We will consider the cases where $d^{+}(v_{0})=0$, $1$, and $2$.
Case 1: Assume that $d^{+}(v_{0})=0$. Then $v_{0}$ has no outgoing edges in $G$. In $G^{2}$, it will still have out-degree 0. Hence its out-degree trivially satisfies the condition of at least doubling.
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. We know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. This node $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. The node $v_{0}$ cannot be one of these nodes. Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4. Thus, it’s out-degree will double.
Therefore, in all cases, $v_{0}$’s out-degree at least doubles in $G^{2}$.
Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs. We are learning that, in order to construct our counterexample, we need our minimum degree to be at least 3. This will also place that same minimum value restriction on every other node in the graph. In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
Lemma 4.2.
Minimum Out-Degree 3 with Neighbor Out-Degree 0 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3. Suppose also that at least one of $v_{0}$’s neighbors has out-degree 0; then the out-degree of $v_{0}$ will at least double in $G^{2}$.
Proof.
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, and $v_{3}\in V(G)$. Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$. Now, assume that $v_{1}$ has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$. As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in $G$.
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$, shares no neighbors with $v_{0}$. This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors, and $v_{1}$ must have at least 3 first neighbors.
Example 4.2.
0
$\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$ H
Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$. Each of these nodes ($v_{1}$, $v_{2}$, and $v_{3}$) connects to the nodes $v_{4}$ and $v_{5}$. The neighbor of $v_{4}$ is $v_{5}$, and the neighbors of $v_{5}$ are $v_{6}$, $v_{7}$, and $v_{8}$. While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$. Thus, in $G^{2}$, the degree of $v_{0}$ will double.
Minimum Out-Degree 3 with Neighbors 1
Lemma 4.3.
Assume that in an oriented graph, the minimal out-degree node $v_{0}$ has an out-degree of 3. Additionally, suppose that in the induced subgraph of its neighbors, every neighbor has an out-degree of 1. The out-degree of $v_{0}$ or one of its neighbors will then at least double in $G^{2}.$.
Proof.
Assume $d^{+}(v_{0})=3$. So $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ We know that $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$. Let $G[v_{1},v_{2},v_{3}]$ represent the induced subgraph by $v_{1},v_{2},v_{3}$.
We can reason that if any of $v_{1},$ $v_{2},$ or $v_{3}$ has an out-degree greater than 3, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$.
Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}].$ If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume they all connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree $<$ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one edge between $v_{4}$ and $v_{5}$. One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$. As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$. ∎
Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) provides an exhaustive search strategy, which is a question we shall revisit throughout this paper. It must prevent a node’s out-degree from doubling in order for that to function. In this instance, we observe that the degree of the starting node in the example $v_{0}$ is not doubled. We are compelled to look for the next node instead. In this case, the degree of the neighbor of $v_{0}$ $v_{1}$ doubles.
We did not require the ”right” starting node; therefore, this is a crucial distinction. Not because of its possible proximity to the double degree node, but rather because of its capacity to divide the nodes in the graph, we selected a minimal degree node.
We shall use double arrows to show that a node or collection of nodes connects to every node in the next subgraph for graphs similar to those in Lemma 4.2 throughout the rest of this work. Because all of the nodes in a shared induced subgraph relate to the same set of nodes, this is possible.
These lemmas offer a theoretical framework and insight into our rationale. These findings will be expanded upon and referenced in the upcoming sections.
5 Theoretical Foundation
We will introduce some new terms to help comprehend this section. The partitions we discussed earlier will now be formally referred to as containers. Additionally, we shall discuss the decreasing sequence property and its significance. After that, we will discuss interior, exterior, and back arcs. We will show the importance of these arcs with some foundational lemmas. The GLOVER data structure will then be used to wrap up.
Example 5.1.
[…]
Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from the minimum degree node, with no assumptions on container relationships or sizes.
Definition 5.1.
A container $C_{i}$ is the set of nodes at distance $i$ from $v_{0}$, together with the arcs connecting those nodes, given a minimum degree node $v_{0}$. Formally speaking,
The formula $C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where $V(G)$ denotes the graph $G$’s vertex set and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
Similar to this, Daamouch employed induced subgraphs in his research on $m$-free graphs [7]. The neighboring containers (see image), which will be crucial in the proofs, are the main distinction here.
We will investigate the structure of oriented graphs from two angles by using the idea of containers. Alternatively, we will investigate it from the viewpoint of containers or the minimum degree node.
  • Containers are seen based on their distance from the minimum degree node.
  • Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
Specifically, the case where $d^{+}(v_{0})=3$ and Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve. A graph $G$ was created in which the out-degree of $v_{0}$ does not double in $G^{2}$. A simple graph of degree three corresponds to the lowest cycle that can exist in an undirected graph. Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the nodes adjacent to $v_{0}$. A closer examination of this case reveals that the out-degree of a node that is an out-neighbor of $v_{0}$ is (at least) twice in this situation. This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
However, we will need to define new concepts in order to proceed with this line of thinking. To prevent a node’s out-degree from doubling, we must first find a term. The decreasing sequence property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$. There are instances where $v_{0}$ lacks its out-degree double in $G^{2}$, as we have seen. We’ll investigate which other nodes can possess the decreasing sequence property and which network features are necessary for this to occur.
Example 5.2.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
Figure 5: This is an illustration where node $v_{0}$ has the Decreasing Sequence Property (DSP). There are more first neighbors $d$ of $v_{0}$ than second neighbors.
Definition 5.2.
If the size of a node $x\in V(G)$ in an oriented graph $G$ is strictly bigger than the size of its second out-neighborhood, that is, $|N^{+}(x)|>|N^{++}(x)|$, then x is said to have the decreasing sequence property.
Two new ideas have been presented thus far: the decreasing sequence property (DSP) and the minimal degree node. A crucial starting point for our approach and degree values is a minimal degree node, which is the node with the fewest connections. It acts as the anchor by being chosen as the beginning point. The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node. It is serving as a stimulant.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
Lemma 5.1.
Total Order Induced by Containers Relative to the Minimum Degree Node
Let $G$ be a graph with a minimum-degree-node $v_{0}$. Construct a series of containers $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{k}$ ordered by distance relative to $v_{0}$. Each container $C_{i}$ is a set of nodes at distance $i$ from $v_{0}$. This imposes a total order on the nodes in $G$. This means that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
Proof.
To prove a total order, we need to show that any two nodes in $G$ can be compared. Let $u,v$ be two nodes in $G$ such that $u\neq v$. We assume that both $u$ and $v$ are reachable from $v_{0}$. Distance is a total ordering, so one of three things is true: either $dist(v_{0},u)<dist(v_{0},v)$, $dist(v_{0},u)=dist(v_{0},v)$, or $dist(v_{0},u)>dist(v_{0},v)$. The situation where $dist(v_{0},u)<dist(v_{0},v)$ implies that $u$ is in a closer container than $v$. The situation where $dist(v_{0},u)=dist(v_{0},v)$ implies that the nodes $u\text{ and }v$ are in the same container. And the situation $dist(v_{0},u)>dist(v_{0},v)$ implies that $u$ is in a further container. This shows that it is a total order. ∎
While imposing order on our graph is a solid first step, the containers do more than just enforce overall order. In essence, they enable group comparisons between nodes that are equally spaced from a reference node, $v_{0}$. A typical total order, such as lexicographical order, does not take this into consideration. This ”containerized ordering” allows us to differentiate nodes based on particular adjacency qualities inside each distance group in addition to their distance from $v_{0}$.
Definition 5.3.
A Level Total Order (LTO) on a set $S$ can be defined as follows:
  • Total Order: For any two elements $u$ and $v$ within the set $S$, exactly one of the following must hold:
    • $u$ is less than $v$ (denoted as $u<v$),
    • $u$ is greater than $v$ (denoted as $u>v$), or
    • $u$ is equal to $v$ (denoted as $u=v$).
  • Leveled Structure: The elements of $S$ are partitioned into a leveling of containers. $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$ Each container itself forms a subset of $S$. This maintains a total order. The containers are sequentially ordered relative to each other. This establishes a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
  • Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$. This avoids trivial total orderings where all elements are isolated in singleton containers.
  • Comparability within Containers: For any two elements $u,v\in C_{i}$, their order can be determined based on a specific criterion.
  • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves.
This order means that we have order in two ways. First, the containers themselves are ordered. Second, the nodes within the containers are ordered. This ordering must also be consistent within the containers.
The notion of a non-triviality requirement is to distinguish this ordering from other traditional total orderings on graphs.
Lemma 5.2.
Decreasing Sequence Property Lemma Suppose a node in an oriented graph $G$ does not have the decreasing sequence property. Then this node will have its out-degree at least double in $G^{2}$.
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$. In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed. We will now give terms to these concepts.
Definition 5.4.
Let $x$ be a node with the decreasing sequence property and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
$int(x,y)=N^{+}(x)\cap N^{+}(y)$
The number of interior neighbors is the interior degree of $y$ with respect to $x$.
Now that we have introduced containers, we can talk about arcs where both nodes are inside the containers; these are interior neighbors. Interior neighbors will be important to our study on SSNC and our search for a counterexample.
In particular, we saw things about interior neighbors in 4 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
Definition 5.5.
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The exterior neighbors of $y$ with respect to $x$, denoted $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$:
$ext(x,y)=N^{+}(y)\cap N^{++}(x)$
The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers. This interaction is again critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container.
As an exercise to visualize interior and exterior neighbors, consider example 5.2. Exterior arcs are the double arrows between the boxes. Interior arcs are not drawn but can be mentally placed either between any two nodes in the container boxes in example 5.2 or example 5.1.
Definition 5.6.
Let $v_{0}$ be a minimum degree node. Suppose that $x$ is a node in the component $C_{i}$. A back arc is defined as an arc $(x,y)$ such that $y\in N^{+}(x)$ and $y\in C_{j}$, where $j<i$.
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. In order to do so, though, we need to consider the option of back-arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to the minimum degree node. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes.
Having defined our arcs, we can talk about the decreasing sequence property. This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from the minimum degree node. Exterior degrees, though, are a part of a node’s degree equation. There’s another part of it though. As we move further from the minimum degree node, the size of the interior degrees increases.
This phenomenon reflects a yin and yang relationship between the interior and exterior degrees. In all of this, the minimum degree node is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees. With this definition of interior and exterior nodes now introduced, we are now ready to introduce the Graph Level Order (GLOVER) data structure.
Definition 5.7.
A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
Leveled Structure: The elements of $S$ are partitioned into levels of containers $C_{1},C_{2},\ldots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
To prove that any oriented graph can be represented in a GLOVER, we’ll need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
Theorem 5.1.
Assume we have an oriented graph $G$. Assuming that there are no back arcs in $G$, this oriented graph can be represented by a GLOVER.
Proof.
To establish a Graph Level Order, we first need a minimum degree node. We can obtain this by taking the minimum value on the degree vector of nodes. We will call that in the node $v_{0}$.
Now that we know the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$.
Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes whose distance from the node $v_{0}$ is equal. These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
  • Total Order on Elements within $G$ Because nodes are partitioned based on their distance from $v_{0}$, we have a total order on the elements of $G$. For any two nodes $u$ and $v$, this total order is determined by their containers.
  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$.
  • Non-triviality Requirement - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node.
  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$.
  • Consistency of Comparisons Across Containers - Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
  • Interior and Exterior Neighbors:
    • Nodes $u$ and $v$ of the same distance from $v_{0}$ would be in the same container and thus be interior neighbors.
    • If the nodes $u$ and $v$ were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance differs by 1.
    Each criterion for a GLOVER is satisfied by partitioning nodes based on distance from $v_{0}$. Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a GLOVER under the distance total order, completing the proof.
What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers. This ordering is also ordering the nodes within the containers, so the GLOVER data structure is a multi-level total order.
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each container maintains its place in the total ordering.
Next, we will show one of the other benefits of this data structure in mathematical induction via well ordering.
Lemma 5.3.
GLOVER Supports Induction
In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
Proof.
We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering.
A set is considered well-ordered if every non-empty subset of that set has a least element.
A discrete total order is a total order where each element has a distinct ”next” element in the order.
Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order. This fulfills the condition of being well-ordered.
Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set.
Example 5.3.
[…]
Figure 6: This example illustrates back arcs in a graph. The two blue arcs from $x_{3}$ to $y_{3}$ and $y_{3}$ to $z_{3}$ represent exterior neighbors indicating forward-directing arcs. Additionally, a red arc points backwards from $z_{5}$ to $x_{1}$. Although the containers are not outlined, they are implied by the difference in variables for each group.
Lemma 5.4.
Interior Degree at least 1
Suppose we have an oriented graph $G$, represented by a GLOVER. Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property. In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $N^{+}(v_{0})$ must have an interior degree of at least 1. That is, for $x\in N^{+}(v_{0})$,
$|\text{int}(v_{0},x)|\geq 1.$
Proof.
By definition, a node satisfies the decreasing sequence property if
$|N^{+}(x)|>N^{++}(x)|$
For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $x$, and then from $x$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $x\in N^{+}(v_{0})$.
If $x$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $x$’s neighbors, which are $v_{0}$’s second neighbors. If $x$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $x$’s first neighbors.
To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
$|int(v_{0},x)|\geq 1$
This argument holds for any neighbor $x\in N^{+}(v_{0})$. This ensures that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
Thus, for all $x\in N^{+}(v_{0})$, $|int(v_{0},x)|\geq 1$.
Lemma 5.4 (Interior Degree at least 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling. In particular, we see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree to double.
Example 5.4.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
Lemma 5.5.
Exterior Less than Degree
Suppose we have an oriented graph $G$, represented by a GLOVER. In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i}),|\text{ext}(x,y)|<d^{+}(x)$.
Proof.
Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$.
By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$
Thus, $|\text{ext}(x,y)|\subseteq N^{++}(x)$, then:
$|N^{++}(x)|\geq|N^{+}(x)|\geq d^{+}(x),$
suggesting that the degree of outwardness of $x$ would at least double on graph $G^{2}$. This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property.
Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$.
Lemma 5.5 (Exterior Less than Degree) is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree at least 1), with the notable exception that Lemma 5.5 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
The results presented in 5.4 (Interior Degree At Least 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1. There, he demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we are developing.
Lemma 5.6.
Weak Decreasing Exteriors Lemma
Suppose that $G$ is an oriented graph represented by a GLOVER, and that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
Suppose that:
  1. 1.
    The interior degree of $y$ with respect to $x$, $|int(x,y)|$, doubles in the subgraph $G[N^{+}(x)]^{2}$.
  2. 2.
    For all $z\in ext(x,y)$, we have $ext(y,z)\cap N^{+}(x)=\emptyset$, i.e., no back arcs from the second neighbors of $x$ to the first neighbors of $x$ exist.
Then, for all $z\in\text{ext}(x,y)$, we have $|\text{ext}(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
Proof.
We begin by recalling the conditions for a node to have the decreasing sequence property (DSP). A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
Now, assume the conditions of the lemma hold:
  1. 1.
    The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increase significantly in $G[N^{+}(x)]$.
  2. 2.
    For all $z\in\text{ext}(x,y)$, $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
Now, suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$.
The node $y$ has a degree that can be broken down into two components, its interior degrees and its exterior degrees. By the assumption that $y$’s interior degree doubles, if $y$’s exterior degree were to double, it would lead to a contradiction.
If there exists a node $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$, then $y$’s out-degree will be impacted by $\text{ext}(x,y)$ and $\text{ext}(y,z)$.
$\begin{split}|\text{ext}(y,z)|&\geq|\text{ext}(x,y)|\\ |\text{ext}(y,z)|+|ext(x,y)|&\geq|\text{ext}(x,y)|+|\text{ext}(x,y)|\\ |\text{ext}(y,z)|+|\text{ext}(x,y)|&\geq 2*|\text{ext}(x,y)|\end{split}$ (1)
This would mean that the exterior degree of $y$ is at least as large as the exterior degree of $x$. However, this contradicts the assumptions about the DSP of $y$, which requires $y$ to have fewer second neighbors than first neighbors.
From $|int(x,y)|$ doubling, $y$’s influence in $G[N^{+}(x)]$ grows, yet its exterior degree should decrease due to the DSP. The absence of back arcs prevents any increase in $y$’s exterior degree from its second neighbors.
Thus, the assumption $|\text{ext}(y,z)|\allowbreak\geq\allowbreak|\text{ext}(x,y)|$ leads to a contradiction, confirming that $|\text{ext}(y,z)|\allowbreak<\allowbreak|\text{ext}(x,y)|$ for all $z\allowbreak\in\allowbreak\text{ext}(x,y)$.
Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$.
The first thing to notice in this proof is the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions are important.
First, we will talk about the lack of back arcs. We have already talked about how back arcs can be a hindrance to our algorithm overall. This is one of the main points where they can be a hindrance. In particular, there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist. We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
Secondly, we assumed that the interior degree of a node doubled. This is similar to the notion of solving the containers separately, then joining them together again to solve the main problem. We are assuming that this node $y$ was in another container and its interior degree doubled in that container. In doing so, $y$ is now in a better position to have its overall degree double, only needing to have its exterior degree double.
The more popular representations of graphs are adjacency matrices, arraylists and edge lists. These suffer from the same limitations of graphs as mentioned above, They scatter the nodes along the two-dimensional plane without attempts to group into an ordering. The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. In order to represent graphs in a GLOVER we would use any markup language that can encode graphs. There needs to be three types of data- containers, nodes, and arcs. The containers will be given ids, the nodes will then be assigned to their unique containers, and arcs will be determined to be interior, exterior, or back, depending on the containers of the endpoints of the arcs.
This determination of the number of containers and the assignment of nodes to containers does not need to be declared beforehand. Instead, just as we have been talking about the distance of a node from a minimum degree node $v_{0}$ as the determination of container assignment, we would have the same or a similar function to assign nodes to containers.
With this container representation of the SSNC, by way of the GLOVER data structure, we can begin to visualize elements of these lemmas. For example, the lemma 5.6 (Weak Decreasing Exteriors Lemma) can be visualized by example 5.5. This shows a series of ordered containers. Each container has fewer nodes. The fewer nodes implies that the container circle can be drawn with a smaller radius, giving light to the decrease in size.
Example 5.5.
[…]
To further illustrate the Graph Level Order (GLOVER) data structure, imagine a social media example where users can follow one another but no-one follows back. This network is then an oriented graph. Running the Seymour Second Neighborhood Conjecture in this context means looking for an account whose indirect sphere of influence is greater than their direct sphere of influence. Here, we are defining the indirect sphere of influence as those accounts they can reach through others, and we are defining the direct sphere of influence as those accounts they can directly connect with. Essentially, these are the same definitions as the second neighbor and first neighbor respectively that we have been working with. With this in mind, after finding the person who connects to the fewest accounts in this social media graph, we could set up containers just as before.
Now, lemma 5.4 (Interior Degree at least 1) says that each person in container 1 must follow another person in container 1. What this is saying is that, inside this first container, we will have one cycle of users where the messages are traversed about, or a number of disjoint cycles where messages are spread among themselves. Considering example 5.6, which takes the former of these approaches. These cycles in the first container are what stops the user Andi from being an influencer.
This brings us to the next container. Lemma 5.6 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 5 people - (Hugo, Isla, Jade, Kai, and Lei). If any person in this container followed less than two other people in this same container, then they would need to follow at least five people in the next container (container 3). From this we would conclude that a person in container 1, whose indirect sphere of influence is this persons sphere of influence, would have its first degree double and its decond degree double, implying that its overall degree doubles.
Example 5.6.
Andi
$\Rightarrow$ BellChadDawnElseFayeGus $\Rightarrow$ HugoIslaJadeKaiLei $\Rightarrow$ H
Figure 7: This is an example of a social media setup of the Seymour Second Neighborhood Conjecture. The Second Container has ten arcs and 5 nodes, which represents a perfect graph. This implies that in the next container $C_{3}$ (not shown), with 4 people can only have 6 arcs. Due to this fact, at least one person in $C_{3}$ will relate to at least five people in $C_{4}$. We can conclude then that at least one person in $C_{1}$ will be an influencer and have their degree double.
Table 1: Intersection Results for Triplets of Nodes in 5.6
Triplet(A, B, C) int(A, B) ext(B, C) Intersection Result
[…]
We have now defined the data structure that we will use throughout the rest of this paper. In doing so, we have a better understanding of interior and exterior arcs, which will help guide us in our search for a counterexample.
6 Back Arcs
With a theoretical foundation underway, we are ready to launch an investigation into back arcs. We will look into whether these arcs can exist in a counterexample and what other properties of oriented graphs must be true.
It is important to restate that back arcs and exterior neighbors are distinct concepts. Back arcs are not a subset of exterior neighbors. While back arcs connect to any previous container, exterior neighbors are limited to connecting with the next container. Additionally, exterior neighbors do not represent all possible external neighbors. Rather, an exterior neighbor ext(x, y) specifically refers to nodes at a distance of two from $x$ that are also neighbors of $y$.
If we look at Example 4.2 again, we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$. Similarly, we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
One of the things we notice about Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. This next lemma formalizes that concept.
Example 6.1.
[…]
Figure 8: Graph representation of back arcs (shown in red) and forward-directed arcs (in blue) between containers, illustrating how these arcs impact connectivity patterns in an oriented graph.
Lemma 6.1.
(Regular Interior Degrees) Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node of $G$. Suppose that the containers of $G$ are $C_{i}$. In an attempt to prevent a degree-doubling node, suppose that two nodes $u$ and $v$ in a container $C_{i}$ have a common parent $v_{p}$ in the preceding container $C_{i-1}$, and that $v_{p}$’s interior degree doubles. If the interior degrees of $u$ and $v$ differ by at least 1, then the exterior degree of $v_{p}$ in container $C_{i-1}$ will double as well.
Proof.
Assume that $u$ and $v$ are two nodes in container $C_{i}$, and that they share a parent $v_{p}$ in container $C_{i-1}$. Suppose that $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$
In order to preserve the necessary degree limits in the graph, $v$ will need fewer external connections if its inner degree is greater than $u$. $u$, on the other hand, will need more external connections to make up for its lack of inner connections.
As a result of the imbalance in internal degrees, the external connections of $u$ are essentially raising the second neighbors of degree of $v_{p}$.
As a result, the exterior degree of $v_{p}$ will double to make room for the additional connections that $u$ in the graph requires.
Thus, these different interior degrees of $u$ and $v$ in container $C_{i}$, will cause the exterior degree of $v_{p}$ in $C_{i-1}$ to double. This leads to a contradiction, so the claim holds. ∎
Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples. In particular, it shows that all the out-neighbors of interior arcs within a container must have the same degree.
In Example 5.6 and Table 1, we examined a minimum degree node $v_{0}$ with degree 6 in a regular graph. The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth.
The definition of a back arc is an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$ respectively. It is implied that the source node $y$ acquires new neighbors at a lower level, maybe doubling its degree, if such a back arc were included. The shortest-path lengths from $v_{0}$ strictly define each level in the hierarchical container organization, which is in contrast to this.
Lemma 6.2.
No Back Arc to Early Container Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node. If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $\text{ext}(u_{i},v_{k})$ is at least the degree of $v_{k}$. Furthermore, if the interior degree of $u_{i}$ doubles, this will cause the total degree of $u_{i}$ to double as well.
Proof.
Assume that $u_{i}\in C_{i}$ has its first back arc to $v_{k}\in C_{k}$ where $k<i-1$. By definition of exterior neighbors:
$|ext(u_{i},v_{k})|=|N^{+}(v_{k})\cap N^{++}(u_{i})|$
Since $N^{+}(v_{k})$ includes all direct neighbors of $v_{k}$, it follows that:
$|N^{+}(v_{k})|=deg(v_{k}).$
By the structure of GLOVER and the absence of back arcs from $u_{i}$ to $v_{0},\dots,v_{k-1}$, the size of $ext(u_{i},v_{k})$ satisfies:
$|ext(u_{i},v_{k})|\leq|N^{+}(v_{k})|=deg(v_{k}).$
Each node in $u_{i}$’s exterior is matched up with a node in $v_{k}$’s neighborhood. This means that the back arc will double $u_{i}$’s exterior degree. Then, because we’ve already assumed that $u_{i}$’s interior degree will double, we can say that $u_{i}$’s overall degree will double. This is true because its interior degree and its exterior degree will double.
Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container-structured graph. It has effectively reduced this case of back arcs to one of interior degree doubling or back arcs mapping to the previous container. This is important because the previous section introduced the GLOVER framework as a total order of containers. Back arcs can disrupt this order. What we see next is a special case of how back arcs could disrupt this order.
Lemma 6.3.
Back Arcs Special Case
Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
  1. 1.
    Suppose $x\in C_{i}$ and $y\in C_{i+1}$, both satisfying the Decreasing Sequence Property (DSP).
  2. 2.
    The interior degree of $y$ with respect to $x$ doubles in $G[N^{+}(x)]^{2}$.
  3. 3.
    There exists a $w\in\text{ext}(x,y)$ such that the $int(y,w)$ doubles.
Then, for all $z\in\text{ext}(x,y)$, we have:
$|\text{ext}(y,z)|<|\text{ext}(x,y)|$
First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e. from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e. not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
Consequently, the arc $w,z$ is an early back arc as we saw in 6.2 (No Back Arc to Early Container). These arcs can lead to double counting. An out-neighbor of $w$ might be an out-neighbor of $x$ and a second out-neighbor of $y$. This implies that under the assumption that $w$’s interior degree doubles, $w$’s overall degree will double through that early back arc.
Lemma 6.4.
First Back Arc Between Consecutive Containers in an Oriented Graph:
Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$. Suppose that we represent this graph by a GLOVER. Suppose we are attempting to prevent a degree-doubling node and that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
Proof.
Assume that the oriented graph $G$ is partitioned into containers $C_{1},C_{2},\ldots,C_{i}$ such that nodes in each container $C_{j}$ (for $j<i$) only have interior arcs in $C_{j}$ or exterior arcs to nodes in $C_{j+1}$, with no back arcs present between containers up to $C_{i}$.
Now, let the first back arc appear from a node $u_{i}\in C_{i}$ to a node $v_{i-1}\in C_{i-1}$. In this oriented graph, by definition, all arcs are supposed to point in the forward direction—from container $C_{j}$ to container $C_{j+1}$. By Lemma 6.1 (Regular Interior Degrees), we know that all nodes inside container $C_{i-1}$ have the same interior degree. Because of their matching interior degrees, they will also share the same exterior degrees. Then since the node $v_{i-1}$ is in $C_{i-i}$, $v_{i-1}$ also maps to these exterior nodes. If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist{v_{0},u_{i}}-i$ so $u_{i}$ cannot belong to container $C_{i}$.
Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ contradicts the oriented nature of $G$, and no back arcs can exist prior to $C_{i}$.
Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees). From this, we can see that a back arc to a parent container violates the orientated nature of the graph.
Lemma 6.5.
No Back Arcs Non-Existence of Back Arc
In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree-doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
Proof.
This was proved in the lemmas 6.2 (No Back Arc to Early Container) and 6.4 (First Back Arc Between Consecutive Containers).
[…]
Figure 9: Illustration of the proof of 7.2 (Interior Degrees Double). This is another illustration of the Interior Degree Doubling. Each node inside a container represents a node whose degree is doubled by the inductive hypothesis. This process cannot continue forever so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
[…]
Figure 10: Illustration of the proof of 7.2 (Interior Degrees Double). Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.
[…]
We stated that back arcs could be detrimental to our algorithm, so the removal of them in this section was important. Now we can better navigate the other properties of oriented graphs.
7 Interior Degrees Doubling
This section introduces the concept of interior degree-doubling. This is similar to solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s interior degree inside of a container. This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
Lemma 7.1.
Indistinguishable Regular Interior Nodes
Suppose we have a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER. We are attempting to prevent a degree-doubling node. If a node $v_{p}$ ’s interior degree doubles, then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree doubled.
Proof.
Let $C_{i}$ be a regular container of size $n$. Let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$. The interior degree of a node $v_{i}$ in a container is defined as the number of arcs shared by $v_{i}$ between $v_{i}$ and a neighbor or parent node. In a regular container, all nodes are indistinguishable in terms of their degree. They also have the same parents because of the DSP. Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve the degree structures. This means any node will have its interior degree doubled.
The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others. Previously we were searching for a node in a container whose interior degree had doubled. That would be the key. What indistinguishable tells us is that if one node’s interior degree doubles inside a container, then all the node’s interior degrees double. That means that we can select any of those nodes.
Lemma 7.2.
Interior Degrees Double
Let $G$ be an oriented graph represented by a GLOVER. We are attempting to prevent a degree-doubling node. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have its interior degree doubled.
Proof.
We prove this by strong induction on the size of the graph.
Base Case: In Lemmas 4.1 (Minimum Out-Degree ¡ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3, there exists at least one node in each container with a doubled interior degree. These cases form our base, supporting the induction.
Inductive Hypothesis: Assume that for any graph of size $n^{\prime}<n$, each container $C_{i}$ contains at least one node whose interior degree doubles.
Inductive Step: Consider a graph of size $n$. We need to show that at least one node in each container $C_{i}$ has its interior degree doubled.
Suppose we construct $G$ such that no node at distance $i$ from $v_{0}$ initially has a doubled interior degree in $C_{i}$. Define a sequence of subgraphs $H_{0}$, $H_{1}$,…,$H_{k}$ for $C_{i}$ as follows:
$H_{0}=\{u_{i}\}$
, where $u_{i}$ is the minimum degree node in $C_{i}$.
$H_{1}=N^{+}(u_{i})$
with nodes in $H_{1}$ having doubled interior degrees by the I.H., as applied to each subgraph. By the Induction Hypothesis (I.H.), there exists a node $y_{1}\in H_{1}$ such that the interior degree of $y_{1}$ with respect to $u_{i}$ doubles in $H_{1}$. We can now define $H_{2}=ext(u_{i},y_{1})$. Likewise, by the I.H., there will be a node $y_{2}\in H_{2}$ whose interior degree doubles in $H_{2}$. We can define $H_{3}=ext(y_{1},y_{2})$. Continuing this process, we can define $H_{j}=ext(y_{j-2},y_{j-1})$, where $y_{j-1}$ is a node whose interior degree doubles in $H_{j-1}$.
By Lemma 5.6 (Weak Decreasing Exteriors Lemma), we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers. An increase in $int(x,y)$ requires a proportional reduction in $ext(y,z)$. This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$.
The strictly decreasing nature of this sequence implies that it must terminate when $H_{i}\leq 2$. At this point, there are not enough arcs (1) in $H_{i}$ for both nodes to have an interior degree of 2. So, one node must have all its degree be exterior.
Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i-1}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$. Thus, the parent of $y^{*}$, denoted $y^{*}-1$, in $G$ achieves a doubled degree within $C_{i}$.
Therefore, for a graph of size $n$, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step.
Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. By finding a positive answer to this, and not just a positive answer, but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. Once again, we see why the Decreasing Sequence Property (DSP) received its name.
This double of interior degrees will help us to build our algorithm in the search for a counterexample. It does, however, give us a bit of light into the side of the Conjecture being true.
8 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The Weak Decreasing Exteriors Lemma implies that containers will decrease in size. We need to see how that decrease in size will impact the degrees of the nodes.
Lemma 8.1.
Regular Exterior Degrees
Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in $G$ has the Decreasing Sequence Property (DSP). Then the nodes in a container in $G$ will result in regular exterior degrees.
Proof.
Let $x_{i}$ and $x_{j}$ be two nodes in the same container $C_{i}$.
Assume, for contradiction, that the sizes of their exterior sets differ by at least one. Without loss of generality, let there exist a node $y$ such that $|ext(x_{i},y)|=|ext(x_{j},y)|+1$. This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
By Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs), we can imply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here. This says that with each distance from the minimum degree node, our exterior size is decreased by 1.
This stopped at the container $C_{i}$. That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$. So, a node in $C-{i-2}$ can have its exterior degree double. By 7.2 and 7.1 (Indistinguishably of Regular Nodes), we know that this node has its interior degree double as well.
So, this assumption led to a contradiction.
This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the Weak Decreasing Exteriors Lemma in this proof because the two go hand in hand.
Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double. The DSP is a bottleneck in this way because it is really limiting what can take place in this graph.
Lemma 8.2.
Strong Decreasing Exteriors Lemma
Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node, so for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1)$
for some $z\in N^{+}(x_{i})$.
Proof.
We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$.
Base Case ($i=1$):
Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
$\{y\in\text{ext}(x_{1},z)\text{ s.t. }dist(v_{0},y)=2\}$
is bounded by $d^{+}(v_{0})-1$. Each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2. We have already proved that there are no back arcs to $v_{0}$.
Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1).$
Inductive Step: We need to show that this bound holds for nodes at distance $i+1$ from $v_{0}$. Let $x_{i+1}$ be a node in container $C_{i+1}$ relative to $v_{0}$. Suppose, for contradiction, that
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|>d^{+}(v_{0})-i.$
This would imply that the exterior neighbors of $x_{i+1}$ exceed the allowed bound, violating the decreasing sequence property (DSP) conditions of $x_{i}$. By Lemmas 7.2 (Interior Degrees Double), 6.1 (Regular Interior Degree), and 7.1 (Indistinguishable Regular Interior Nodes), this would cause the interior degree of $x_{i}$ to double, thereby forcing an overall doubling of the degree for $x_{i}$.
This contradiction implies that our assumption must be false. Therefore, we conclude:
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|\leq d^{+}(v_{0})-i.$
Lemma 8.2 (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma) to give us a way of calculating the maximum size of a container based only on its distance from the minimum degree node and the minimum degree node’s degree.
The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” was originally named.
Lemma 8.3.
Interior Degrees Lower Bound
Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. Suppose every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided $(v_{0},x)\in G$.
Proof.
Let $x$ be a node at distance $i$ from $v_{0}$. By Lemma 8.2 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$, $|ext(v_{i-1},x)|$, is bounded by $d^{+}(v_{0})-i$. This places a limit on the number of nodes $x$ can map to at distance $i+1$ from $v_{0}$.
Now, since $x$ has at least the same out-degree as $v_{0}$, $x$ must map to at least $i$ other nodes. Given that back arcs are not asked, $x$ cannot map to nodes at earlier distances.
Therefore, the only remaining possibility is that $x$ maps to interior nodes. Since $x$’s degree must account for at least $i$ mappings. These must be to interior nodes. So, we conclude that $x$ must map to at least $i$ interior nodes.
Thus, $|int(v_{i-1},x)|\geq i$, as required.
Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is an opposite bound for the interiors. Once again, the minimum degree node is the driving force behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
Lemma 8.4.
Regular Graph
All Nodes Have Equal Out-Degree
Let $G$ be an oriented graph represented by a GLOVER. Suppose $v_{0}$ is a minimum degree node of $G$. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node. Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $\delta=d^{+}(v_{0})$.
Proof.
With the proofs on the lack of back arcs and the regularity of interior arcs and the regularity of exterior arcs, we arrive at the following equation for each container:
$\delta=|int(v_{p},u_{i})|+|ext(v_{p},u_{i})|$
where $u_{i}$ is in container $C_{i}$, and $v_{p}$ is a parent of $u_{i}$ in container $C_{i-1}$.
For a node $u_{i}$ in container $C_{i}$, with degree $\delta+1$, there would be the possible places for this degree to be located—back arcs, interior arcs, or exterior arcs. We have already shown that back arcs cannot exist. If this extra degree went to the exterior arcs, then it would cause the degree of a parent node to double.
The only remaining possibility is for this extra degree to go to the interior arcs of $u_{i}$. While this is not a problem, it will cause the interior degrees to increase at a faster rate than we just showed in Lemma 8.3 (Interior Degree Lower Bound), as this would increase the number of interior arcs.
This more rapid reduction in the interior degree leads to container density faster. To avoid this, regularity is key.
Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph.
We are incrementally refining the properties that must hold for our counterexamples. The two most crucial aspects are the Decreasing Sequence Property and the identification of the minimum degree node. However, additional constraints, such as the overall graph regularity established by Lemma 8.4 (Regular Graph), help simplify future arguments. This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
9 Transitive Triangles and Seymour Diamonds
Before we continue with the development of our DSP algorithm and show how it solves the Seymour Second Neighborhood Conjecture, we need to address some potential problems and show that they will not hinder our algorithm.
Definition 9.1.
Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, and $u$ form a transitive triangle.
Here, we have $x$ pointing to $y$ and $u$ and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
Definition 9.2.
Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, and $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$. This is a four-vertex-oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture.
The notion of interior neighbors is another way of describing what [1] referred to as transitive triangles. They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e., one where every node has the decreasing sequence property. We provided a similar result with Lemma 5.4 (Interior Degree at least 1). When a minimum degree node $v_{0}$ has the Decreasing Sequence Property, all of its neighbors have an interior degree of at least 1. That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$, will form a transitive triangle. We can extend that further for any two constructive containers.
xyu
Lemma 9.1.
Transitive Triangles from Containers
Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). For any two consecutive containers $C_{i}$ and $C_{i+1}$, there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle.
Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm.
Lemma 9.2.
Seymour Diamonds from Containers
Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). for any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
$u\in C_{i},\quad v,w\in C_{i+1},\quad y\in C_{i+2}$
The following arcs exist:
1. There are forward arcs $x\to y$ and $y\to u$,
2. There are forward arcs $x\to w$ and $w\to u$,
3. No back arcs exist between containers.
Similar to Lemma 9.1, Lemma 9.2 (Seymour Diamonds from Containers) lets us know that we have been interacting with Seymour Diamonds all along without knowing it. We didn’t know it because they did not disrupt our algorithm.
xywu
Lemma 9.3.
Seymour Diamonds and Decreasing Path Algorithm
Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node, where we are applying the DSP algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP algorithm.
Proof.
A Seymour diamond consists of two vertices, $x$ and $u$, connected by two parallel paths via $y$ and $w$.
The DSP algorithm can utilize one of these paths while disregarding the other. Both paths connect $x$ and $u$, thus maintaining degree constraints.
The presence of redundant paths does not diminish the outgoing neighbors of any node. This ensures that the minimum degree condition holds throughout the process.
Consequently, these additional connections offer alternative routes without disrupting the algorithm’s operation.
None of these cases are a deterrent to the execution of the DSP algorithm.
Lemma 9.4.
Interaction of Minimum Degree Nodes and Transitive Triangles
Suppose $G$ is an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node with degree $d$. Consider a transitive triangle formed by nodes $z$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(z)=0$.
Then the following cases are possible:
  • $d^{+}(y)=d$ and $d^{+}g(z)=d$
  • $d^{+}(y)=d$ and $d^{+}(z)>d$
  • $d^{+}(y)>d$
  • $d^{+}(y)<d$
Proof.
  • Case 1: $d^{+}(y)=d$ and $d^{+}(z)=d$:
    Both $y$ and $z$ have exactly $d$ outgoing connections. For the triangle to be transitive, they must not share all neighbors. This means that there exists at least one neighbor of $y$ that is not a neighbor of $z$. Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
  • Case 2: $d^{+}(y)=d$ and $d^{+}(z)>d$:
    Then $z$ may cover some neighbors of $N^{+}(y)$. Since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
  • Case 3: If $d^{+}(y)>d$:
    $y$ has more outgoing connections than required. It can utilize these additional connections to map to unique neighbors without impacting the minimum degree condition, ensuring $y$ retains neighbors not covered by $z$.
  • Case 4: If $d^{+}(y)<d$:
    This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections. Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
Now that we see that transitive triangles and Seymour diamonds are not a problem for our algorithm, we can continue our investigation of oriented graphs. Next, we will blend everything together into our main results.
10 Main Theorem: Seymour’s Conjecture is True
10.1 Introduction
Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some final results.
So far, this battle has been a battle between a catalyst, the chosen minimum degree node and the bottleneck, the Decreasing Sequence Property (DSP). We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
There is a limit, though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
We first restate the Lemma 6.5 (Non-Existence of Back-Arcs), from section 6, but slightly rephrased.
Lemma 10.1.
Let $G$ be an oriented graph represented by a GLOVER, with all nodes satisfying the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node. If there are back arcs in $G$, then a node’s degree will double.
10.2 Theorem Statement
Degree Doubling by Exterior Decreasing
Theorem 10.1.
Let $G$ be an oriented graph represented by a GLOVER with all nodes having the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node where we set up our partition. Then Seymour’s conjecture holds true.
Proof.
We have set up the containers by a GLOVER data structure. There are no back arcs, else we have a degree-doubling node. Next is lemma 7.2 (Interior Degree Doubling), which states that in every container we will have a node whose interior degree doubles.
These lemmas are what allowed us to set up and use lemma 5.6 (Weak Decreasing Exteriors Lemma). That lemma gives us a decreasing set of positive numbers. This set is bound to reach the size of two. When it does, one of the nodes in that set will have all of its arcs to the next container (thereby being exterior arcs). This will cause a parent of that node’s degree to double.
Algorithm 1 Algorithm for Analyzing Degree Doubling in Containers
1:Initialization:
2:Choose the minimum degree node $v_{0}$.
3:Divide nodes into containers $C_{1},C_{2},\ldots$ based on their degree and proximity to $v_{0}$.
4:for each container $C_{i}$ in $G$ do:
5:     Locate a node $x_{i}$ in $C_{i}$ whose interior degree doubles.
6:     Move to a neighboring container $C_{i+1}$.
7:     if there is a back arc from $C_{i+1}$ to a prior container then
8:         Locate child of $x_{i}$ whose exterior degree doubles, causing $x_{i}$’s overall degree to double.
9:         Halt.
10:     else
11:         Since there is no back arc, the container $C_{i+1}$ will be bound in size to be 1 less than the size of container $C_{i}$.
12:         Repeat from Step 4.
13:     end if
14:end for
Start with graph size $n$Select minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots$Identify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to prior containerLocate child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
Figure 11: Flowchart depicting the degree-doubling algorithm applied within containers, demonstrating the process of identifying nodes with the Decreasing Sequence Property.
Corollary 10.1.
Degree Doubling by Container Density
In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},\allowbreak C_{2},\allowbreak\ldots\allowbreak C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
There are two possibilities:
  • Case 1: The containers continue to shrink in size. Simultaneously, the interior degrees continue to grow in size. We reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
  • Case 2: For small minimum degree nodes, the container’s shrinkage is not fast enough to cause this collision.
Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and increased number of nodes. This was represented by the yin and yang effects early on. We see it come full circle here. Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbors represented the size of the containers.
In small graphs like 4 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur. To understand the reasoning behind this, understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1. The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$. So, for small graphs, we are just not going out far enough to catch up. For larger graphs, though, we do start to see this.
Corollary 10.2.
Occurrence of Last Dense Container
Let $G$ be an oriented graph represented by a GLOVER with minimum degree $k$. Then there will be a container that reaches maximal density. Further, if that container is $C_{i}$, it will be $i$-out-regular.
Proof.
We begin by stating lemma 6.1 (Regular Interior Degrees). While these are called interior degrees, these are the out-degrees of the nodes in a container. This lemma says that every container must have regular out-degrees.
The second part of this proof is lemma 8.2 (Strong Exterior Decreasing), which says that in container $i$, the size of the exterior is at most $d^{+}(v_{0})-i$. Lemma 8.3 (Interior Degree Lower Bound) tells us that each node must have degree $i$ in this container.
Thus, we have a series of regular containers that are getting smaller as we get further from the minimum degree none. At the same time, these containers are expected to hold fewer nodes of higher degree, i.e., a graph of more arcs. This is progressively getting to a denser container. As Corollary 10.1 (Degree Doubling by Container Density), this cannot continue forever. There must be a last, most dense container.
The Seymour Second Neighborhood Conjecture is made more beautiful by the appearance of these orientations of out-regular graphs as containers get smaller. Not only has the conjecture been proven true, but with an inherent symmetry and order. There is a visual pattern emerging inside these containers that is not just beautiful to graph theorists, but there is a structure that is easily appealing to non-mathematicians.
Corollary 10.3.
Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $k$, and let the most dense graph in container $C_{i}$. Then every node in container $C_{i}$ has its degree doubled in the representation.
The discovery of multiple degree-doubling nodes lifts the Seymour Second Neighborhood Conjecture from a mere question of existence to one in abundance. There were many who doubted that one even existed. Now there are many.
Second is the way we find it. It’s the presence of an oriented regular graph as these containers decrease in size. There’s an element of mathematical beauty there.
Then third we can look at the applications. We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group.
Theorem 10.2.
Suppose we have a graph $G$. The Decreasing Sequence Property Algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ represents the number of vertices in $G$ and $|E|$ represents the number of edges in $G$.
Proof.
We will first show that the complexity of building the partitions has a worst-case complexity of $O(|V|+|E|)$.
Given a graph $G$, the search for a minimum degree node that’s $O(|V|)$ nodes. If a priority queue is used, this can be optimized further. Then, to evaluate the degrees of each of those nodes is a constant factor. This makes the search for the minimum degree node $O(|V|)$.
Once we have this value, we will place every node and every edge of $G$ into a preprocessing array ready to be partitioned. Each edge is examined, contributing $O(|E|)$ complexity, while nodes contribute $O(|V|)$. Each node will individually be processed based on its distance from the minimum degree node, along with the connected outgoing arcs. This will require $O(|V|)$ comparisons and $O(|V|+|E|)$ stores into containers. This means that the partitioning has a complexity of $O(|V|+|E|)$.
Next, we will look into the complexity of the run-time analysis of the Decreasing Sequence Algorithm.
First, remember that we proved in lemma 7.2 (Interior Degree Doubles) that every container has a node whose interior degree doubles. Secondly, we also note that the lemma 6.1 (Regular Interior Degrees) implies that all the nodes are regular inside a container. Third, lemma 7.1 (Indistinguishable Regular Interior Nodes) lets us know that nodes within a container are indistinguishable from one another. The combination of these three lemmas let us know that every node inside a container has their interior degree double. This gives the complexity of finding a node whose interior degree has doubled $O(1)$.
What remains is to consider the complexity necessary to combine these interior degree nodes to solve the overall problem. Solving the full problem requires determining the number of containers necessary for this partition, which is bounded above by $O(|V|)$. The total number of containers cannot exceed the number of nodes, $|V|$, since containers are based on partitions. Processing each container’s exterior arcs and remaining possible nodes is bounded by $O(|V|+|E|)$. This gives the run time a complexity of $O(|V|+|E|)$.
This complexity analysis shows that the Decreasing Sequence Property Algorithm can effectively find a degree-doubling node in polynomial time. Moreover, this will work for any oriented graph. Complexity analysis shows that it is not hindered by larger graphs.
A further point is to consider the concept in theorem 10.2 of all nodes inside a container having their degrees double. It was not spoken about in the paper because the focus was entirely set on proving the SSNC, but this is another example of multiple degree-doubling nodes, just a localized instance.
11 Conclusion and Future Work
In this work, we explored the Seymour Second Neighborhood Conjecture in oriented graphs. We did so through the lenses of graph theory, combinatorics, algorithms, and data structures. This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide-and-conquer paradigm.
One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph. This effectively solves the Seymour Second Neighborhood Conjecture with an answer in the affirmative. The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree-doubling node was in the graph. Instead, the algorithm selects a minimum degree node and is able to partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
This approach introduced a new perspective on the Seymour Second Neighborhood Conjecture. It was no longer framed as merely a graph theory problem. By utilizing the divide-and-conquer paradigm and solving different parts independently, this algorithm shows that the Seymour Second Neighborhood Conjecture could effectively be treated with computer science techniques. This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists. Alternatively, it may be that there is not enough interdisciplinary communication about these problems and the potential techniques to solve them.
The Seymour Second Neighborhood Conjecture started as a question in 1990. Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers. It has been talked about in books. on message boards, in classrooms, and at conferences. This problem has a long history.
With a solution now in hand, it’s time to begin looking forward. The Seymour Second Neighborhood Conjecture can be represented as a social media problem instead of an oriented graph. Here, the nodes are people, where no one follows back. Now instead of wondering if there is a person whose sphere of influence doubles, we would simply be searching for that person. We just need to implement the algorithm to find the person. We could have the graph represent epidemiology. Here the nodes will be infected people, and the arcs would be one-way infections. The question of interest would be, can we implement the algorithm to find the person who is a super-spreader and do so quickly? There are many more real-life examples where we can apply this problem and this algorithm. There are fields like security, telecommunications, urban planning, and neural networks. There are too many to name in this short paper.
Secondly, we introduced a data structure, Graph Level Order-GLOVER. This data structure allowed us to get a lot done on the conjecture. Many of the things inherent to SSNC made it easier for GLOVER—the well ordering and the dual metrics of distance and degrees.
Future research and development should explore questions of complexity and scalability. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
12 Acknowledgments
I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:6 has been a guiding light throughout this journey. I am profoundly grateful to my family—my mother, father, and wife—for their unwavering support and encouragement.
Special thanks to the late Dr. Nate Dean for his contributions to this work and to Dr. Michael Ball for his support. I want to sincerely thank everyone who has taken the time to pray for me, share encouraging words, and offer constructive feedback. Your support has meant so much to me, and I’m truly grateful for the kindness and thoughtfulness you’ve shown.
Finally, I extend heartfelt gratitude to my family and friends for their patience and belief in me throughout this journey.
The GLOVER data structure is protected by a provisional patent application filed with the U.S. Patent and Trademark Office.
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Ranked Structure: The elements of $S$ are partitioned into a ranking of ”containers” or ”levels,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Definition 13.4.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the interior neighbors of $y$ with respect to $x$, denoted as $int(x,y)$, are the out-neighbors of $y$ that are also out-neighbors of $x$. Formally,
$int(x,y)=N^{+}(x)\cap N^{+}(y),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$ in the graph. The number of interior neighbors of $y$ with respect to $x$ is called the interior degree of $y$.
Definition 13.5.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the exterior neighbors of $y$ with respect to $x$, denoted as $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$. Formally,
$ext(x,y)=N^{+}(y)\cap N^{++}(x),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$, and $N^{++}(x)$ denotes the set of out-neighbors of the out-neighbors of $x$ (i.e., the second neighbors of $x$). The number of exterior neighbors of $y$ with respect to $x$ is called the exterior degree of $y$.
Definition 13.6.
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
Definition: Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
Definition 13.7.
Definition: Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
14 Appendix 2: Applicability of the Graph Level Order
The deciding factor that held the algorithm together and helped solve the Seymour Second Neighborhood Conjecture was the Graph Level Order (GLOVER) data structure. This research did not begin as an investigation into data structures. Instead, a series of operations were conducted on the oriented graphs - partitioning, ordering those partitions, adding interior and exterior arcs, until what was left was an ordering of both the partition and nodes within the partition. Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric. Unlike the Seymour Second Neighborhood Conjecture, exterior arcs do not always have to only be defined to the next container. That was a problem specific definition.
Moving forward, the GLOVER data structure should be well suited to tackle problems in graph theory and beyond. By introducing containers, another way of doing induction on graphs has been introduced. The concept of divide-and-conquer, which has dominated much of computer science, shows up in this paper as well. We are able to treat these containers independent of one another and solve their problems locally before bringing them back to the global problem. This shows how versatile the GLOVER data structure is moving forward.
What makes GLOVER special are things like total order on containers. Now there is a two-way ranking of nodes, as opposed to traditional one-way rankings like lexicographical sorting. This agrees with many real-world systems where we have dual metrics that are often competing, like price and performance. This is hard to measure on a single scale.
Next these nodes are not just treated like clusters inside these containers, but like nodes in graph theory. The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth first search, or centrality to be called. What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
This is still a new data structure, so we are learning a lot about it. There are many applications that have not been tried yet. However, since it brings the strengths of a total order and a graph into one unified structure, it should be able to model complex relationships. This should include domains of machine learning like clustering, dynamic scheduling, resource management, project management, and natural language processing.
15 Appendix 3: Visualization and Accessibility of Oriented Graphs
When we take a course on graph theory, oriented graphs are probably not on the course load. Second neighbors or squares of graphs are probably not, either. To fully grasp this problem, though, we need to be able to delve into these definitions through visualizations.
For example, research on the Seymour Second Neighborhood Conjecture led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs. The paper addressed how these terms were intertwined with each other. An example is where the paper says that there is a monotonically decreasing sequence.
James Robert Brown’s book Philosophy of Mathematics speaks of the importance of visualizations in mathematics. Proofs are still the foundation, but sometimes visualizations can serve as inspiration or a spark for those proofs.
This realization motivated the creation of a complimentary website. The site serves as an interactive tool to understand the research. It also has things like mathjax installed to read lemmas, definitions, and proofs from this paper. Where possible, things like HTML canvas and JavaScript D3.js are used to visualize these concepts and help improve understanding.
Readers are encouraged to explore the website for additional information and examples at: https://glovermethod.com This is an evolving project. Additional examples and applications will continue to be added. Readers are invited to explore the website. Please provide feedback on any aspects of the site or content for improvement.
Our goal here was to help users connect with the Seymour Second Neighborhood Conjecture. Hopefully, this tool not only serves as support for this paper, but as an educational resource in the fields of graph theory and computer science. Thank you for your engagement