### v7

# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

Abstract
Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph. In this paper, we first present a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers. This data structure establishes a well-ordering on oriented graphs and allows for the construction of a decreasing sequence of subsets of nodes. This sequence proves the non-existence of counterexamples to the SSNC and precisely identifies the required node. Further, our approach finds the occurrence of dense regular graphs inside containers. This finding extends the SSNC to the discovery of multiple nodes satisfying the degree doubling property. Beyond theoretical implications, the algorithm and data structure have practical applications in network optimization and algorithm design.
1 Introduction
The Seymour Second Neighborhood Conjecture (SSNC), proposed by Paul Seymour in 1990, posits that in every oriented graph, there exists a node whose out-degree in the graph’s square is at least twice its out-degree in the original graph. This conjecture was first formally published by Dean and Latka in 1995 [4], where they proposed a related version specific to tournaments. The tournament case was proved by Fisher [8], with an alternative proof later provided by Havet and Thomassé [10].
Conjecture 1.1.
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}(v)$ denote the first and second neighborhoods of $v$, respectively.
Despite decades of research, the SSNC remains one of graph theory’s challenging open problems. Special cases of the conjecture have been explored extensively. Kaneko and Locke [11] showed that the SSNC holds for graphs with a minimum degree of at most six. Brantner et al. [1] demonstrated that if the conjecture were false, infinitely many counterexamples would exist. Chen et al. [3] established a lower bound, proving that every oriented graph contains a vertex whose second neighborhood is at least $lambda$ times larger than its first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Botler et al. [2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
The concept of $m$-free graphs is particularly relevant to this paper, as it closely aligns with the work of Daamouch [7]. A directed graph $D$ is $m$-free if it contains no directed cycles of length at most $m$. In his paper, Daamouch introduces induced subgraphs and explores their role in the degree properties of vertices—an idea that will also play a significant role in the discussions and results presented here.
This paper builds on these foundations, generalizing an approach initially developed for the minimum degree 7 case [9]. Our key contribution is the introduction of a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers. This structure enables the construction of a decreasing sequence of subsets of nodes, forming the basis of our algorithmic proof of the SSNC. Our approach not only proves the conjecture but also identifies the specific node(s) whose out-degree doubles in the graph’s square.
Beyond its theoretical significance, the SSNC has practical applications in domains like social networks, epidemiology, and algorithm design. For instance, in the context of disease control, the conjecture helps identify individuals whose influence in their second-degree connections surpasses their direct contacts, offering insights into the spread of infections.
Unlike traditional proof-theoretic approaches, this work emphasizes the development of an efficient algorithm to locate the target node. We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems. The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.
The remainder of this paper is structured as follows: Section 2 introduces common terminology used in graph theory. Section 3 outlines our approach to the problem, while Section 4 provides illustrative examples. Section 5 focuses on conditions under which a node’s out-degree may fail to double, and Section 6 examines the role of back arcs in oriented graphs. Section 7 explores the conditions that lead to doubling of interior degrees, examining their significance in relation to the SSNC. It introduces a divide-and-conquer approach, where the graph is partitioned, and each subproblem is solved independently before being integrated into the broader structure. Section 8 explores density constraints in containers. Section 9 investigates the impact of transitive triangles and Seymour diamonds on our algorithm. Finally, Section 10 presents the full Decreasing Sequence Property algorithm and its proof of the conjecture.
2 Graph Theory Terminology
A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $(x,x)$ where $x$ is a node in $G$) and no symmetric arcs, that is, no arcs of the form $(x,y)$ and $(y,x)$ where $x$ and $y$ are nodes in $G$. The square of a graph $G$, denoted $G^{2}$, consists of the same node set as $G$, and an arc exists from node $x$ to node $z$ if there is a node $y$ such that both arcs $(x,y)$ and $(y,z)$ exist in $G$ as well as the arcs already in $G$. Essentially, $G^{2}$ captures two-step connections from $x$ to $z$.
Definition 2.1.
Let $G^{2}=(V,E^{2})$ where $G=(V,E)$ is the original graph, and $E^{2}$ is the set of edges defined as:
$E^{2}={(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E}$
We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$. These are the nodes that $v$ has arcs pointing to. Similarly, $N^{++}(v)$ refers to the second neighbors of $v$, that is, nodes that are two steps away from $v$ in $G$.
Let $G=(V,E)$ be a graph and let $S\subseteq V$ be any subset of vertices. The induced subgraph $G[S]$ is the graph whose vertices are the set $S$ and whose edge set consists of all edges in $E$ that have both endpoints in $S$. For any node $v$, we refer to the neighbor-induced subgraph on $N^{+}(v)$, consisting of $v$’s out-neighbors. Let $u,v$ be nodes in $G$. The distance between $dist(u,v)$ is $k$ if the shortest path from $u$ to $v$ has length $k$.
Example 2.1.
[…]
Figure 1: A five-vertex graph illustrating node out-degrees and distances. Out-degrees are labeled near each node, and the dashed rectangle indicates an induced subgraph on vertices A, B, and C. Previously used in [9]
Conjecture 2.1.
(Square version of Seymour’s Second Neighborhood Conjecture). Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when you square the oriented graph.
Now that we have two different conjectures, we can investigate how they relate to one another. Conjecture 1.1 relates the notion of a node’s degree doubling to its neighborhoods. Similarly, conjecture 2.1 relates this same concept to the square graph. The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling, and its second neighborhood being greater than or equal to the first. By the definition of an arc in the square graph, we can see that an arc is defined in this square graph only for paths of length two in the original graph. This is why these two conjectures are seen as identical.
We will focus our attention on Conjecture 1.1 and Conjecture 2.1. Conjecture 1.1 is easier to work with initially and understand why a node’s degree will double in the square graph. Concentrating on neighborhood analysis yielded numerous insights and features that aided in the development of our algorithm.
Next, we turn our attention to the term degree-doubling nodes. By this term, we aim to get to the heart of conjecture 1.1. That conjecture is looking for a node whose degree at least doubles. In order to do this, we examine the node’s first and second neighborhoods. Then identify a node whose second neighborhood is at least as large as the first, as required by the conjecture. A similar definition, a Seymour vertex, fulfills conjecture 2.1 - it’s out-degree in the squared graph is at least double its original out-degree. We could equivalently search for a Seymour vertex or a degree-doubling node.
3 Approach and Motivation
We begin our analysis by considering general oriented graphs. We make no initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs. We know that there is an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
From this foundation, we can now partition the graph into $d_{v_{0}}+1$ subgraphs corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm. No nodes or arcs are being removed to form these partitions. This is a conceptual idea. Instead, think of it as putting a box or circle around the nodes and arcs. Now the nodes inside the box can work with the other nodes inside the box and solve their problem independent of the other partitions. These partitions are still connected to one another and will have to reconnect to solve the global problem. However, thinking of them in this way allows us to use their independent nature.
This represents a stark contrast to most previous approaches to the SSNC. Many of these previous approaches were conceptual or theoretical. Papers would mention the existence of a node whose degree doubled. The notion, though, of a path to that node was not part of the paper. Secondly, most of these papers focused on local search techniques that started at or near the desired degree-doubling node. That presents a problem because oriented graphs can be varying in nature. This is the first global search technique on the SSNC problem.
Complexity enters the picture when discussing global search algorithms. This raises concerns about how well they will handle larger graphs in comparison to smaller ones. This should not be a problem, as we will see in the following sections. The slowest part of this algorithm is building the data structure. That will require partitioning the nodes and edges of the graph, in a worst-case complexity of $O(V+E)$.
When the graph is partitioned, even larger graphs, often with higher value minimum degree nodes and thus more partitions, are able to be broken into subproblems. These subproblems are then able to be solved independent of one another. These subproblems are each their own instances of the SSNC, just a localized version. What the partition is doing is breaking a node’s individual degree into two parts—the degree within the container and the degree to the next container. The degree within the container is the subproblem that will be solved separately, and the degree to the next container is how they connect back to the global problem.
We still need to show that this is a formidable foundation. We have laid the groundwork for our algorithm to tackle the oriented graphs. With a minimum degree node, we can partition this graph into independent subproblems. This is now ready for a global search algorithm.
4 Initial Lemmas
These first lemmas will seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. For this section, unless otherwise noted, assume that $v_{0}$ is a node with minimum out-degree in our oriented graph $G$. The work in this section agrees with what was already done by [11] and is not intended as original content. However, it does provide insight into our reasoning.
Example 4.1.
$v_{0}$$v_{2}$$v_{3}$$v_{1}$$v_{6}$$v_{5}$$v_{4}$
$\Rightarrow H$
Figure 2: Example illustrating how a minimum out-degree node in $G$ has its degree doubled in $G^{2}$ under certain conditions.
Lemma 4.1.
Minimum Out-Degree $<$ 3 Suppose that a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
Proof.
We will consider the cases where $d^{+}(v_{0})=0$, $1$, and $2$.
Case 1: Assume that $d^{+}(v_{0})=0$. Then $v_{0}$ has no outgoing edges in $G$. Then $v_{0}$’s first and second neighborhoods are of size 0. Hence, its out-degree trivially satisfies the condition of at least doubling.
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. We know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. This node $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$. This means that $(v_{1},v_{2})\in G$. Hence $v_{2}$ is a second neighbor of $v_{0}$, thus making it at least as large as $v_{0}$’s first neighborhood.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. Consider the induced subgraph $G[v_{1},v_{2}]$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph. This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. The node $v_{0}$ cannot be one of these nodes. Hence $v_{0}$ will have at least two second neighbors - $v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first. Thus, its out-degree will double in $G^{2}$.
Therefore, in all cases, $v_{0}$’s out-degree at least doubles.
Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs. We are learning that, to construct our counterexample, we need the minimum degree of our graph to be at least 3. In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
Lemma 4.2.
Minimum Out-Degree 3 with Neighbor Out-Degree 0 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3. Suppose also that at least one of $v_{0}$’s neighbors has out-degree 0. Then the out-degree of $v_{0}$ will at least double in $G^{2}$.
Proof.
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, and $v_{3}\in V(G)$. Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$.
Now, assume that $v_{1}$ has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
We know that $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has at least 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$. This means that $v_{0}$ has at least three second neighbors. As a result, $v_{0}$ has at least as many second neighbors as first neighbors, and equivalently, $v_{0}$’s out-degree doubles in $G^{2}$. ∎
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$, shares no neighbors with $v_{0}$. This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors, and $v_{1}$ must have at least 3 first neighbors.
Example 4.2.
0
$\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$ H
Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$. Each of these nodes ($v_{1}$, $v_{2}$, and $v_{3}$) connects to the nodes $v_{4}$ and $v_{5}$. The neighbor of $v_{4}$ is $v_{5}$, and the neighbors of $v_{5}$ are $v_{6}$, $v_{7}$, and $v_{8}$. While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$. Thus, in $G^{2}$, the degree of $v_{1}$ will double. See lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1)
Lemma 4.3.
Minimum Out-Degree 3 with Neighbors 1 Assume that in an oriented graph, the minimal out-degree node $v_{0}$ has an out-degree of 3. Additionally, suppose that in the induced subgraph of its neighbors, every neighbor has an out-degree of 1. Then the out-degree of $v_{0}$ or one of its neighbors will then at least double in $G^{2}$.
Proof.
Assume $d^{+}(v_{0})=3$. So $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ We know that $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$. Let $G[v_{1},v_{2},v_{3}]$ represent the induced subgraph by $v_{1},v_{2},v_{3}$.
We can reason that if any of $v_{1},$ $v_{2},$ or $v_{3}$ has an out-degree greater than 3, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$.
Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}]$. If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume that they all connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
This reduces the situation to Case 3 in Lemma 4.1 (Minimum Out-Degree $<$ 3), where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one edge between $v_{4}$ and $v_{5}$. One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$. As a result, $v_{4}$ will have three second neighbors, and that is at least as many as its first neighbors. This means that degree doubles in $G^{2}$. ∎
Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) provides a global search strategy, which is a question we shall revisit throughout this paper. It must prevent a node’s out-degree from doubling for that to function. In this instance, we observe that the degree of the starting node in the example $v_{0}$ is not doubled. We are compelled to look for the next node instead. In this case, the degree of the neighbor of $v_{0}$, $v_{1}$, doubles.
We did not require the ”right” starting node. Therefore, this is a crucial distinction. Not because of its possible proximity to the degree-doubling node. Instead, we chose a minimal degree node due to its ability to split the graph’s nodes.
We shall use double arrows to show that a collection of nodes connects to every node in the next induced subgraph as shown in example 4.2. Because all nodes in a shared induced subgraph relate to the same set of nodes, this is possible.
These lemmas offer a theoretical framework and insight into our rationale. These findings will be expanded upon and referenced in the upcoming sections.
5 Theoretical Foundation
We will introduce some new terms to help with the rest of this paper. The partitions we discussed earlier will now be formally referred to as containers. Additionally, we shall discuss the decreasing sequence property and its significance. After that, we will discuss interior, exterior, and back arcs. We will show the importance of these arcs with some foundational lemmas. The GLOVER data structure will then be used to wrap up.
Example 5.1.
[…]
Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from the minimum degree node, with no assumptions on container relationships or sizes.
Definition 5.1.
A container $C_{i}$ is the set of nodes at distance $i$ from $v_{0}$, together with the arcs connecting those nodes, given a minimum degree node $v_{0}$. Formally speaking,
The formula
$C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where
$V(G)$ denotes the graph $G$’s vertex set and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
Similar to this, Daamouch employed induced subgraphs in his research on $m$-free graphs [7]. The neighboring containers (see example 5.1), which will be crucial in the proofs, are the main distinction here.
We will investigate the structure of oriented graphs from two angles by using the idea of containers. Alternatively, we will investigate it from the viewpoint of containers or the minimum degree node. Containers are seen based on their distance from the minimum degree node. Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
Specifically, the case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve. A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double in $G^{2}$. A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph. Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the nodes adjacent to $v_{0}$. If we look at this situation further, we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
However, we will need to define new concepts to proceed with this line of reasoning. To prevent a node’s out-degree from doubling, we must first define a term. The decreasing sequence property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$. There are instances where $v_{0}$ lacks its out-degree double in $G^{2}$, as we have seen. We’ll investigate which other nodes can possess the decreasing sequence property and which network features are necessary for this to occur.
Example 5.2.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
Figure 5: This is an illustration where node $v_{0}$ has the Decreasing Sequence Property (DSP). There are more first neighbors $d$ of $v_{0}$ than second neighbors.
Definition 5.2.
If the size of a node $x\in V(G)$ in an oriented graph $G$ is strictly bigger than the size of its second out-neighborhood, that is, $|N^{+}(x)|>|N^{++}(x)|$, then $x$ is said to have the decreasing sequence property.
Two competing new ideas have been presented, the decreasing sequence property (DSP) and the minimal degree node. A crucial starting point for our approach and degree values is a minimal degree node. It acts as the anchor by being chosen as the beginning point. The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node. It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DSP introduces constraints on connectivity. This concept is akin to a bottleneck. A bottleneck in graph theory represents a point that limits flow or connectivity within the network, a role precisely filled by the DSP. Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
Lemma 5.1.
Total Order Induced by Containers Relative to the Minimum Degree Node
Let $G$ be a graph with a minimum-degree node $v_{0}$. Construct a series of containers $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{k}$ ordered by distance relative to $v_{0}$. Each container $C_{i}$ is a set of nodes at distance $i$ from $v_{0}$. This imposes a total order on the nodes in $G$. This means that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
Proof.
To prove a total order, we need to show that any two nodes in $G$ can be compared. Let $u,v$ be two nodes in $G$ such that $u\neq v$. We assume that both $u$ and $v$ are reachable from $v_{0}$. Distance is a total ordering, so one of three things is true: either $dist(v_{0},u)<dist(v_{0},v)$, $dist(v_{0},u)=dist(v_{0},v)$, or $dist(v_{0},u)>dist(v_{0},v)$. The situation where $dist(v_{0},u)<dist(v_{0},v)$ implies that $u$ is in a closer container than $v$. The situation where $dist(v_{0},u)=dist(v_{0},v)$ implies that the nodes $u\text{ and }v$ are in the same container. And the situation $dist(v_{0},u)>dist(v_{0},v)$ implies that $u$ is in a further container. This shows that it is a total order. ∎
While imposing order on our graph is a solid first step, the containers do more than just enforce overall order. In essence, they allow for the grouping of comparisons between nodes that are equally spaced from a reference node, $v_{0}$. A typical total order, like lexicographical order, does not take this into consideration. This ”containerized ordering” allows us to differentiate nodes based on particular adjacency qualities inside each distance group in addition to their distance from $v_{0}$.
Definition 5.3.
A Level Total Order (LTO) on a set $S$ can be defined as follows:
  • Total Order: For any two elements $u$ and $v$ within the set $S$, exactly one of the following must hold:
    • $u$ is less than $v$ (denoted as $u<v$),
    • $u$ is greater than $v$ (denoted as $u>v$), or
    • $u$ is equal to $v$ (denoted as $u=v$).
  • Leveled Structure:
    • The elements of $S$ are partitioned into a leveling of containers. $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$
    • Each container itself forms a subset of $S$.
    • The containers are sequentially ordered relative to each other.
  • Comparability within Containers: For any two elements $u,v\in C_{i}$, their order can be determined based on a specific criterion.
  • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$.
This order means that we have order in two ways. First, the containers themselves are ordered. Second, the nodes within the containers are ordered. This ordering must also be consistent within the containers.
Lemma 5.2.
Decreasing Sequence Property Lemma Suppose a node in an oriented graph $G$ does not have the decreasing sequence property. Then this node will have its out-degree at least double in $G^{2}$.
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$. In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed. We will now give terms to these concepts.
Definition 5.4.
Let $x$ be a node with the decreasing sequence property and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
$int(x,y)=N^{+}(x)\cap N^{+}(y)$
The number of interior neighbors is the interior degree of $y$ with respect to $x$.
Now that we have introduced containers, we can talk about arcs where both nodes are inside the containers; these are interior neighbors. Interior neighbors will be important to our study on SSNC and our search for a counterexample.
In particular, we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1). The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
Definition 5.5.
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The exterior neighbors of $y$ with respect to $x$, denoted $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$:
$ext(x,y)=N^{+}(y)\cap N^{++}(x)$
The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
We contrast interior neighbors with exterior neighbors, which reveals how a container interacts with adjacent containers. This interaction is again critical for the SSNC, as the nodes within the current container depend on the connections to the next container.
As an exercise to visualize interior and exterior neighbors, consider example 5.2. Exterior arcs are the double arrows between the boxes. Interior arcs are not drawn but can be mentally placed either between any two nodes in the container boxes in example 5.2 or example 5.1.
Definition 5.6.
Let $v_{0}$ be a minimum degree node. Suppose that $x$ is a node in the component $C_{i}$. A back arc is defined as an arc $(x,y)$ such that $y\in N^{+}(x)$ and $y\in C_{j}$, where $j<i$.
We have spoken of so much of our algorithm depending on the distance property of this oriented graph. We would like to continue this. To do so, though, we need to consider the option of back arcs and see how they might hurt our algorithm. These have the possibility of altering a container’s distance to the minimum degree node. Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes.
Having defined our arcs, we can talk about the decreasing sequence property. This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from the minimum degree node. Exterior degrees, though, are a part of a node’s degree equation. There’s another part of it though. As we move further from the minimum degree node, the size of the interior degrees increases.
This phenomenon reflects a yin and yang relationship between the interior and exterior degrees. In all of this, the minimum degree node is serving as a catalyst and driving force for the process. At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees. With this definition of interior and exterior nodes now introduced, we are now ready to introduce the Graph Level Order (GLOVER) data structure.
Definition 5.7.
A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
Leveled Structure: The elements of $S$ are partitioned into levels of containers $C_{1},C_{2},\ldots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Next we will prove that any oriented graph can be represented in a GLOVER. We will need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties. This includes maintaining a total order of the containers, allowing for the comparison of any two nodes, establishing a total order of those nodes within the containers, and establishing interior and exterior neighbors.
Theorem 5.1.
Given an oriented graph $G$. Assuming that there are no back arcs in $G$, this oriented graph can be represented by a GLOVER data structure.
Proof.
To establish a graph level order, we first need a minimum degree node. So we take the minimum value on the degree vector of nodes. We will call that node $v_{0}$. Knowing the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$. Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes at distance $i$ from $v_{0}$. These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
  • Total Order on Elements within $G$ Nodes are partitioned based on their distance from $v_{0}$. This is a total order on the nodes of $G$. For any two nodes $u$ and $v$, this total order is determined by their containers.
  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ each correspond to nodes at a specific distance from $v_{0}$.
  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric to distinguish $u$ from $v$.
  • Consistency of Comparisons Across Containers - Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
  • Interior and Exterior Neighbors:
    • Nodes $u$ and $v$ of the same distance from $v_{0}$ are in the same container and thus are interior neighbors.
    • If the nodes $u$ and $v$ were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance differs by 1.
    Each criterion for a GLOVER is satisfied by partitioning nodes based on distance from $v_{0}$. Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a GLOVER under the distance total order, completing the proof.
What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers. This ordering is more than just representative, though. The containers help to organize the nodes. What this suggests is that each partition, or set of nodes, is being totally ordered as we move away from the chosen minimum degree node. This is a clear distinction from standard graph theory. Graph nodes are generally thought of as able to be placed anywhere in 2-dimensional space. However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from the chosen minimum degree node.
The one thing that could disrupt this total ordering is a back arc. This could potentially lead to a feedback loop. By assuming no back arcs, we ensure that each container maintains its place in the total ordering. However, it is a bold assumption to claim that oriented graphs have no back arcs. Oriented graphs come in various types; surely some exist with back arcs.
This is an important claim to investigate, and the next section will look into it in depth. We will prove that these back arcs do not hurt the algorithm. Instead, they only lead to the faster path to a degree-doubling node. In addition, there will no longer be a need to represent the oriented graph in a GLOVER, as we will have already found the degree-doubling node.
Next, we will show one of the other benefits of this data structure in mathematical induction via well ordering. Mathematical induction is a proof technique that uses inference to prove statements. Remember that a well ordering is a concept in discrete mathematics where a total ordering on a set $S$ has the property that every non-empty subset of $S$ has a least element.
Lemma 5.3.
GLOVER Supports Induction In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
Proof.
We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering. A set is considered well-ordered if every non-empty subset of that set has a least element. A discrete total order is a total order where each element has a distinct ”next” element in the order. Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order. This fulfills the condition of being well-ordered. Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set. ∎
This becomes an important concept because now, in addition to the nodes, we also have the containers as subsets of $S$. These containers are totally ordered and thus can also be used as a means of mathematical induction. We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from the chosen minimum degree node $v_{0}$.
Example 5.3.
[…]
Figure 6: This example illustrates back arcs in a graph. The two blue arcs from $x_{3}$ to $y_{3}$ and $y_{3}$ to $z_{3}$ represent exterior neighbors indicating forward-directing arcs. Additionally, a red arc points backwards from $z_{5}$ to $x_{1}$. Although the containers are not outlined, they are implied by the difference in variables for each group.
Lemma 5.4.
Interior Degree at least 1 Suppose we have an oriented graph $G$, represented by a GLOVER. Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property. In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $C_{1}$ must have an interior degree of at least 1. That is, for $x\in C_{1}$,
$|\text{int}(v_{0},x)|\geq 1.$
Proof.
By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$ If $x$ has $d^{+}(v_{0})$ exterior arcs, this would make $v_{0}$’s degree double because $x$’s neighbors are $v_{0}$’s second neighbors. To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
$|int(v_{0},x)|\geq 1$
This is true for any neighbor $x\in C_{1}$. This ensures that every node in $C_{1}$ has an interior degree of at least 1. Thus, for all $x\in C_{1}$, $|int(v_{0},x)|\geq 1$. ∎
Lemma 5.4 (Interior Degree at least 1) is important because we can see some structural properties in generally oriented graphs that may coincide with degree doubling. We see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree doubled.
Lemma 5.5.
Exterior Less than Degree Suppose we have an oriented graph $G$, represented by a GLOVER. In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
Proof.
Assume, for the sake of contradiction, that there is a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$. By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$. Thus, $|\text{ext}(x,y)|\subseteq N^{++}(x)$. Then:
$|N^{++}(x)|\geq|N^{+}(x)|\geq d^{+}(x),$
This means that the out-degree of $x$ at least doubles on the graph $G^{2}$. This contradicts the assumption that $x$ has the decreasing sequence property.
This must be true for all $x$ in $G$, so we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
Lemma 5.5 (Exterior Less than Degree) is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree at least 1), with the notable exception that Lemma 5.5 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
The results presented in 5.4 (Interior Degree At Least 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1. There, he demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an global search solution with the algorithm we are developing.
Lemma 5.6.
Weak Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph represented by a GLOVER. Suppose further that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
Suppose that:
  1. 1.
    The interior degree of $y$ with respect to $x$, $|int(x,y)|$, doubles in the induced subgraph $G[N^{+}(x)]^{2}$.
  2. 2.
    For all $z\in ext(x,y)$, we have $ext(y,z)\cap N^{+}(x)=\emptyset$, i.e., no back arcs from the second neighbors of $x$ to the first neighbors of $x$ exist.
Then, for all $z\in\text{ext}(x,y)$, we have $|\text{ext}(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
Proof.
We begin by recalling the conditions for a node to have the decreasing sequence property (DSP) because it is very important for this proof. A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$. Now, assume the conditions of the lemma hold:
  1. 1.
    The interior degree $int(x,y)$ doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increase significantly in $G[N^{+}(x)]$.
  2. 2.
    For all $z\in\text{ext}(x,y)$, $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$; no back arcs exist.
Suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$. The node $y$ has a degree that can be broken down into two components, its interior degrees and its exterior degrees. By the assumption that $y$’s interior degree doubles, if $y$’s exterior degree were to double, it would lead to a contradiction. If there exists a node $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$, then $y$’s out-degree will be impacted by $\text{ext}(x,y)$ and $\text{ext}(y,z)$.
$\begin{split}|\text{ext}(y,z)|&\geq|\text{ext}(x,y)|\\ |\text{ext}(y,z)|+|ext(x,y)|&\geq|\text{ext}(x,y)|+|\text{ext}(x,y)|\\ |\text{ext}(y,z)|+|\text{ext}(x,y)|&\geq 2*|\text{ext}(x,y)|\end{split}$ (1)
This would mean that the exterior degree of $y$ is at least as large as the exterior degree of $x$. However, this contradicts the assumptions about the DSP of $y$, which requires $y$ to have fewer second neighbors than first neighbors.
From $|int(x,y)|$ doubling, $y$’s influence in $G[N^{+}(x)]$ grows, yet its exterior degree should decrease due to the DSP. The absence of back arcs prevents any increase in $y$’s exterior degree from its second neighbors.
Thus, the assumption $|\text{ext}(y,z)|\allowbreak\geq\allowbreak|\text{ext}(x,y)|$ leads to a contradiction, confirming that $|\text{ext}(y,z)|\allowbreak<\allowbreak|\text{ext}(x,y)|$ for all $z\allowbreak\in\allowbreak\text{ext}(x,y)$.
Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$. ∎
The first thing to notice in this proof is the assumptions. We assume that there are no back arcs from $x$’s second neighbors to $x$’s first neighbors and that a $y$’s interior degree doubles. Both of these assumptions are important.
First, we will talk about the lack of back arcs. We have already talked about how back arcs can be a hindrance to our algorithm overall. This is one of the main points where they can be a hindrance. There exists a way to double count nodes as both first neighbors and second neighbors if back arcs exist. We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
Secondly, we assumed that the interior degree of a node doubled. This is like the notion of solving the containers separately, then joining them together again to solve the main problem. We are assuming that this node $y$ was in another container and its interior degree doubled in that container. In doing so, $y$ is now in a better position to have its overall degree double, only needing to have its exterior degree double.
Notice the wording of interior degree doubling. That is a separate version of the SSNC, but localized to the containers. The conjecture in this case is can we find a node in each container whose degree doubles. If so, then it gives light to this divide-and-conquer approach, where we partition the containers, solve them individually, and then combine them again for this lemma 5.6 (Weak Decreasing Exteriors Lemma).
The more popular representations of graphs are adjacency matrices, arraylists, and edge lists. These suffer from the same limitations of graphs as mentioned above; they scatter the nodes along the two-dimensional plane without attempts to group them into an ordering. The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way. For the SSNC, that reasonable way was the distance metric. To represent graphs in a GLOVER, we can use any data serialization that can encode graphs, including JSON, YAML, and XML. There needs to be three types of data: containers, nodes, and arcs. The containers will be given IDs. The nodes will then be assigned to their unique containers. Arcs will also be determined to be interior, exterior, or back, depending on the endpoints of the nodes of the arcs.
The determination of the number of containers and the assignment of nodes to containers also do not need to be declared beforehand. Instead, just as the data structure uses the minimum function to find a node representing the minimum degree node $v_{0}$. This same function will return a value $\delta=d^{+}(v_{0})$, and $\delta$ will represent the maximum number of containers. We would have a similar function to assign nodes to containers.
With this container representation of the SSNC, by way of the GLOVER data structure, we can begin to visualize elements of these lemmas. For example, the lemma 5.6 (Weak Decreasing Exteriors Lemma) can be visualized by example 5.4. This shows a series of ordered containers. Each container has fewer nodes. The fewer nodes implies that the container circle can be drawn with a smaller radius, giving light to the decrease in size.
Example 5.4.
[…]
Figure 7: Illustration of the Weak Decreasing Exteriors Lemma. We have a node $x$ in container $C_{i}$ connecting to three nodes in container $C_{i+1}$. In order to prevent $x$’s degree from doubling, we need to have $|C_{i+2}|<|C_{i+1}|$.
To further illustrate the Graph Level Order (GLOVER) data structure, imagine a social media example where users can follow one another, but no one follows back. This network is then an oriented graph. Running the SSNC in this context means looking for an account whose indirect sphere of influence is at least as large as their direct sphere of influence. Here, we are defining the indirect sphere of influence as those accounts they can reach through others. Similarly, we are defining the direct sphere of influence as those accounts they can directly connect with. Essentially, these are the same definitions as the second neighbor and first neighbor, respectively, that we have been working with. After finding the person who connects to the fewest accounts in this social media graph, we could set up containers just as before.
Now, lemma 5.4 (Interior Degree at least 1) says that each person in container 1 must follow another person in container 1. What this is saying is that, inside this first container, we will have one cycle of users where the messages are traversed about or a number of disjoint cycles where messages are spread among themselves. Considering example 5.5, which takes the former of these approaches. These cycles in the first container are what stops the user Andi from being an influencer.
This brings us to the next container. Lemma 5.6 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 5 people—Hugo, Isla, Jade, Kai, and Lei. If any person in this container followed less than two other people in this same container, then they would need to follow at least five people in the next container (container 3). From this we would conclude that a person in container 1, whose indirect sphere of influence is this person’s sphere of influence, would have its first degree double and its second degree double, implying that its overall degree doubles.
Example 5.5.
Andi
$\Rightarrow$ BellChadDawnElseFayeGusHugoIslaJade $\Rightarrow$ KaiLeiMiaNovaOwenPatQinRyan $\Rightarrow$ SageTateUgoVeraWadeXiaYoel $\Rightarrow$ H
Figure 8: This is an example of a social media setup of the SSNC. The second container has 21 arcs and seven nodes, which represents an orientation of the complete graph.
We have now defined the data structure that we will use throughout the rest of this paper. In doing so, we have a better understanding of interior and exterior arcs, which will help guide us in our search for a counterexample.
6 Back Arcs
With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs. We will investigate whether these arcs can bring about a counterexample.
It is important to note that the back arcs and the exterior neighbors are distinct concepts. Back arcs are not a subset of exterior neighbors. While back arcs may connect to any previous container, exterior neighbors are limited to connecting with the next container. In addition, exterior neighbors do not represent all possible future containers. Rather, an exterior neighbor $ext(x,y)$ specifically refers to nodes at a distance of two from $x$ that are also distance one from $y$.
If we look at Example 4.2 again, we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$. Similarly, we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
One of the things we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. The next lemma formalizes that concept.
Example 6.1.
[…]
Figure 9: Graph representation of back arcs (shown in red) and forward-directed arcs (in blue) between containers, illustrating how these arcs impact connectivity patterns in an oriented graph.
Lemma 6.1.
(Regular Interior Degrees) Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node of $G$. Suppose that the containers of $G$ are $C_{i}$. In an attempt to prevent a degree-doubling node, suppose that two nodes $u$ and $v$ in a container $C_{i}$ have a common parent $v_{p}$ in the preceding container $C_{i-1}$, and that $v_{p}$’s interior degree doubles. If the interior degrees of $u$ and $v$ differ by at least 1, then the exterior degree of $v_{p}$ in container $C_{i-1}$ will double as well.
Proof.
Assume that $u$ and $v$ are two nodes in container $C_{i}$. Suppose that they share a parent $v_{p}$ in container $C_{i-1}$. Suppose also that $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$.
If $|int(v_{p},v)|>|int(v_{p},u)|+1$, then $v$ will have fewer exterior neighbors than $u$. Similarly, $u$ will have more external neighbors because of its fewer interior neighbors.
These exterior neighbors of $u$ are raising the second neighbors of $v_{p}$.
As a result, the exterior degree of $v_{p}$ will double. Because $v_{p}$ has already had its interior degree doubled, it will have its overall degree double
Thus, the different interior degrees of $u$ and $v$ in container $C_{i}$ leads to a contradiction, so the claim holds. ∎
Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within the containers of these counterexamples. It shows that all the out-neighbors of interior arcs within a container must have the same degree. This may not be surprising, as the examples we have been showing have all had this property. For example, both containers 1 and 2 in example 5.5 are orientations of regular graphs. The second container is actually an orientation of $K_{5}$, the complete graph on 5 nodes. This result begins to lend itself to some of the mathematical symmetry and structure that must be present in these containers and the SSNC.
The definition of a back arc is an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$ respectively. It is implied that the source node $y$ acquires new neighbors at a lower level, maybe doubling its degree, if such a back arc were included. The shortest-path lengths from $v_{0}$ strictly define each level in the hierarchical container organization, which contrasts with this.
Next, we begin to look at how back arcs interact with the data structure and the development of the algorithm. Up to this point, we have gotten around back arcs by assuming that they do not exist. However, that will not last forever. These next few lemmas address the possibility of what happens when we try to construct a GLOVER and encounter a back arc. We will see that although we are not able to construct a GLOVER, we are still able to execute our algorithm and find a degree-doubling node.
Consider the following table based on example 5.5.
Table 1: Consequences of Back Arcs in 5.5
Back Arc From(A) Back Arc To (B) Back Arc New Neighbors via $N^{+}(B)$
Sage Faye Gus, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Tate Gus Hugo, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Ugo Hugo Isla, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Vera Bell Chad, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Wade Chad Dawn, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Xia Dawn Else, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
Yoel Else Faye, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
This table looks at the possible back arcs from container 3 in example 5.5 and how they would impact the degree doubling of the node in question. The back arcs constructed in this table are attempts to go from container 3 $C_{3}$ to container $C_{1}$. In each case, we can see that $N^{+}(B)$ causes that degree of the source node to double. What follows is a lemma that generalizes this table.
Lemma 6.2.
No Back Arc to Early Container Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node. If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $\text{ext}(u_{i},v_{k})$ is at least the degree of $v_{k}$. Furthermore, if the interior degree of $u_{i}$ doubles, this will cause the total degree of $u_{i}$ to double as well.
Proof.
Assume that $u_{i}\in C_{i}$ has its first back arc to $v_{k}\in C_{k}$ where $k<i-1$. By definition of exterior neighbors:
$|ext(u_{i},v_{k})|=|N^{+}(v_{k})\cap N^{++}(u_{i})|$
We know there are no back arcs from $u_{i}$ to $v_{0},\dots,v_{k-1}$, so the size of $ext(u_{i},v_{k})$ satisfies:
$|ext(u_{i},v_{k})|\leq|N^{+}(v_{k})|=deg(v_{k}).$
This means that the back arc will double $u_{i}$’s exterior degree. Then, because we’ve already assumed that $u_{i}$’s interior degree will double, we can say that $u_{i}$’s overall degree will double. This is true because its interior degree and its exterior degree will double. ∎
Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container-structured graph. It has effectively reduced this case of back arcs to one of interior degree doubling or back arcs mapping to the previous container. This is important because the previous section introduced the GLOVER framework as a total order of containers. Back arcs can disrupt this order. What we see next is a special case of how back arcs could disrupt this order.
Lemma 6.3.
Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
  1. 1.
    Suppose $x\in C_{i}$ and $y\in C_{i+1}$, both satisfying the Decreasing Sequence Property (DSP).
  2. 2.
    The interior degree of $y$ with respect to $x$ doubles in $G[N^{+}(x)]^{2}$.
  3. 3.
    There exists a $w\in\text{ext}(x,y)$ such that the $int(y,w)$ doubles.
Then, for all $z\in\text{ext}(x,y)$, we have:
$|\text{ext}(y,z)|<|\text{ext}(x,y)|$
First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e., from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e., not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
Consequently, the arc $w,z$ is an early back arc as we saw in 6.2 (No Back Arc to Early Container). These arcs can lead to double counting. An out-neighbor of $w$ might be an out-neighbor of $x$ and a second out-neighbor of $y$. This implies that under the assumption that $w$’s interior degree doubles, $w$’s overall degree will double through that early back arc.
Lemma 6.4.
First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$. Suppose that we represent this graph by a GLOVER. Suppose we are attempting to prevent a degree-doubling node and that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
Proof.
Assume that the oriented graph $G$ is partitioned into containers $C_{1},C_{2},\ldots,C_{i}$ such that nodes in each container $C_{j}$ (for $j<i$) only have interior arcs in $C_{j}$ or exterior arcs to nodes in $C_{j+1}$, with no back arcs present between containers up to $C_{i}$.
Now, let the first back arc appear from a node $u_{i}\in C_{i}$ to a node $v_{i-1}\in C_{i-1}$. In this oriented graph, by definition, all arcs are supposed to point in the forward direction—from container $C_{j}$ to container $C_{j+1}$. By Lemma 6.1 (Regular Interior Degrees), we know that all nodes inside container $C_{i-1}$ have the same interior degree. Because of their matching interior degrees, they will also share the same exterior degrees. Then since the node $v_{i-1}$ is in $C_{i-i}$, $v_{i-1}$ also maps to these exterior nodes. If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist(v_{0},u_{i})-i$ so $u_{i}$ cannot belong to container $C_{i}$.
Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ contradicts the oriented nature of $G$, and no back arcs can exist prior to $C_{i}$. ∎
Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees). From this, we can see that a back arc to a parent container violates the oriented nature of the graph.
Lemma 6.5.
No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree-doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
Proof.
This was proved in the lemmas 6.2 (No Back Arc to Early Container) and 6.4 (First Back Arc Between Consecutive Containers). ∎
[…]
Figure 10: Illustration of the proof of 7.2 (Interior Degrees Double). This is another illustration of the Interior Degree Doubling. Each node inside a container represents a node whose degree is doubled by the inductive hypothesis. This process cannot continue forever, so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
[…]
Figure 11: Illustration of the proof of 7.2 (Interior Degrees Double). Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.
We stated that back arcs could be detrimental to our data structure. Now that we have shown that even if they are present, they will not hinder the development of our pathfinding algorithm, as a back arc will simply lead to a degree-doubling node faster. Now we can better navigate the other properties of oriented graphs.
7 Interior Degrees Doubling
This section introduces the concept of interior degree-doubling. This is like solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s interior degree inside of a container. This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
Lemma 7.1.
Indistinguishable Regular Interior Nodes Suppose we have a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER. We are attempting to prevent a degree-doubling node. If a node $v_{p}$ ’s interior degree doubles, then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree double.
Proof.
Let $C_{i}$ be a regular container. Let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$. In a regular container, all nodes are indistinguishable in terms of their degree. They also have the same parents because of containers. Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve the degree structures. This means any node will have its interior degree doubled under the automorphism $\phi$. ∎
The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others. Previously we were searching for a node in a container whose interior degree had doubled. That would be the key. What indistinguishable tells us is that if one node’s interior degree doubles inside a container, then all the nodes within that container will have their interior degrees double. That means that we can select any of those nodes.
We still need to prove that for each container there exists a node whose interior degree doubles. We have seen this in the example 5.5, but those could be too simple to observe any real patterns. In similar structure, what we saw in 4.1 (Minimum Out-Degree $<$ 3) was that it was easy to find degree-doubling nodes in graphs with a minimum degree of 0, 1, or 2. However, when that minimum degree was increased to 3, there were more cases that needed to be considered.
[…]
Lemma 7.2.
Interior Degrees Double Let $G$ be an oriented graph represented by a GLOVER. We are attempting to prevent a degree-doubling node. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have its interior degree doubled.
Proof.
We prove this by strong induction on the size of the graph. Base Case: In Lemmas 4.1 (Minimum Out-Degree $<$ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4.3 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3, there exists at least one node in each container with a doubled interior degree. These cases form our base, supporting the induction.
Inductive Hypothesis: Assume that for any graph of size $n^{\prime}<n$, each container $C_{i}$ contains at least one node whose interior degree doubles.
Inductive Step: Consider a graph of size $n$. We need to show that at least one node in each container $C_{i}$ has its interior degree doubled.
Suppose we construct $G$ such that no node at distance $i$ from $v_{0}$ initially has a doubled interior degree in $C_{i}$. Define a sequence of subgraphs $H_{0}$, $H_{1}$,…,$H_{k}$ for $C_{i}$ as follows:
$H_{0}=\{u_{i}\}$
where $u_{i}$ is the minimum degree node in $C_{i}$.
$H_{1}=N^{+}(u_{i})$
with nodes in $H_{1}$ having doubled interior degrees by the I.H., as applied to each subgraph. By the Induction Hypothesis (I.H.), there exists a node $y_{1}\in H_{1}$ such that the interior degree of $y_{1}$ with respect to $u_{i}$ doubles in $H_{1}$. We can now define $H_{2}=ext(u_{i},y_{1})$. Likewise, by the I.H., there will be a node $y_{2}\in H_{2}$ whose interior degree doubles in $H_{2}$. We can define $H_{3}=ext(y_{1},y_{2})$. Continuing this process, we can define $H_{j}=ext(y_{j-2},y_{j-1})$, where $y_{j-1}$ is a node whose interior degree doubles in $H_{j-1}$.
By Lemma 5.6 (Weak Decreasing Exteriors Lemma), we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers.
This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$.
The strictly decreasing nature of this sequence implies that it must terminate when $H_{i}\leq 2$. At this point, there are not enough arcs in $H_{i}$ for both nodes to have an interior degree of 2. So one node must have all of its degree be exterior.
Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i-1}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$. Thus, the parent of $y^{*}$, denoted $y^{*-1}$, in $G$ achieves a doubled degree in $C_{i}$.
Therefore, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step. ∎
Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture. We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles. By finding a positive answer to this, and not just a positive answer but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets. This cannot continue indefinitely. This gives rise to why the Decreasing Sequence Property (DSP) received its name. Prior to proofs, our research showed the Weak Decreasing Exteriors Lemma as an early observation through many examples. The questions of whether this was always true or whether back arcs like lemma 6.3 could force this into being a false positive.
This proof also helps with explaining the concept of the SSNC. Any oriented graph in which we want to seek a degree-doubling node can be represented in a GLOVER, as long as there are no back arcs. Once we have done this, there are containers spread out from a chosen minimum degree node. Now for each container we can simply run the conjecture on that container. The difference is that with lemma 7.2 (Interior Degrees Double), we already know that there is a node whose degree doubles inside that container. Then with lemma 7.1 (Indistinguishable Regular Interior Nodes), we know that any node we select in that container will also have their interior degree double.
This notion of interior degrees will help us to build our algorithm in the search for a counterexample. It also gives us a bit of light into the possibility of the conjecture being true.
8 Container Density
Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean. The Weak Decreasing Exteriors Lemma implies that containers will decrease in size. We need to see how that decrease will impact the degrees of the nodes, the other properties of containers, and the development of our algorithm.
Lemma 8.1.
Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in a container $C_{i}$ has the Decreasing Sequence Property (DSP). Then the nodes in $C_{i}$ will result in regular exterior degrees.
Proof.
Let $x_{i}$ and $x_{j}$ be two nodes in the same container $C_{i}$.
Assume, for contradiction, that the sizes of their exterior sets differ by at least one. Without the loss of generality, let there exist a node $y$ such that $|ext(x_{i},y)|\geq|ext(x_{j},y)|+1$. This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
Because Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs) both hold, we can apply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here. This says that with each increase in distance from the minimum degree node, our exterior size is decreased by 1.
This decreasing of the exteriors (and thus the containers) is stopped at the container $C_{i}$. This is where our assumption of the different exterior sizes comes into play. That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$. So, a node in $C-{i-2}$ can have its exterior degree double. By 7.2 and 7.1 (Indistinguishably of Regular Nodes), we know that this node has its interior degree double as well.
So, this assumption led to a contradiction. ∎
This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 5.6 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double. The DSP is a bottleneck in this way because it is really limiting what can happen in this graph in the instance of not having a node’s degree double.
Lemma 8.2.
Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node. For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors, and thus $C_{i}$ must be bounded by
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1)$
for some $z\in N^{+}(x_{i})$.
Proof.
We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$. Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
$\{y\in\text{ext}(x_{1},z)\text{ s.t. }dist(v_{0},y)=2\}$
is bounded by $d^{+}(v_{0})-1$. Each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2. We have already proved that there are no back arcs to $v_{0}$.
Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1).$
Inductive Step: We need to show that this bound holds for nodes at distance $i+1$ from $v_{0}$. Let $x_{i+1}$ be a node in container $C_{i+1}$ relative to $v_{0}$. Suppose, for contradiction, that
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|>d^{+}(v_{0})-i.$
This would imply that the exterior neighbors of $x_{i+1}$ exceed the allowed bound, violating the decreasing sequence property (DSP) conditions of $x_{i}$. By Lemmas 7.2 (Interior Degrees Double), 6.1 (Regular Interior Degree), and 7.1 (Indistinguishable Regular Interior Nodes), this would cause the interior degree of $x_{i}$ to double, thereby forcing an overall doubling of the degree for $x_{i}$.
This contradiction implies that our assumption must be false. Therefore, we conclude:
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|\leq d^{+}(v_{0})-i.$
Lemma 8.2 (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma) to give us a way of calculating the maximum size of a container based only on its distance from the minimum degree node and the minimum degree node’s degree.
In Section 5 (Theoretical Foundation), we demonstrated that mathematical induction could be applied to our framework. This proof highlights the utility of the lemma. It allowed us to construct a proof based on the container distance from the chosen minimum-degree node. It is possible that this result could have been proven without relying on induction. We could have used a different approach or simply avoided induction based on container distance from the minimum degree node entirely. Still, the use of induction provides an important method, particularly for graphs where the usage of induction is more limited. Induction may not always be necessary, but having a variety of alternative approaches to proofs strengthens our theoretical foundation and enhances our flexibility in tackling these problems.
Lemma 8.3.
Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node in $G$. Suppose every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a parent of $x$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$.
Proof.
Let $x$ be a node at distance $i$ from $v_{0}$.
By Lemma 8.2 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$, $|ext(v_{i-1},x)|$, is bounded by $d^{+}(v_{0})-i$. This places a limit on the number of nodes $x$ can map to at distance $i+1$ from $v_{0}$.
Now, since $x$ has at least the same out-degree as $v_{0}$, $x$ must map to at least $i$ other nodes. Given that back arcs are not allowed, $x$ cannot map to nodes at earlier distances.
Therefore, the only remaining possibility is that $x$ maps to interior nodes. These must be to interior nodes. So, we conclude that $x$ must map to at least $i$ interior nodes.
Thus, $|int(v_{i-1},x)|\geq i$, as required. ∎
Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is a contrasting bound for the interiors. Once again, the minimum degree node is the driving force, the catalyst, behind this. The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
Lemma 8.4.
Regular Graph All Nodes Have Equal Out-Degree Let $G$ be an oriented graph represented by a GLOVER. Suppose $v_{0}$ is a minimum degree node of $G$. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node. Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $\delta=d^{+}(v_{0})$.
Proof.
For a node $u_{i}$ in container $C_{i}$, with degree $\delta+1$, there would be the possible places for this degree to be located—back arcs, interior arcs, or exterior arcs. We have already shown that back arcs cannot exist. If this extra degree went to the exterior arcs, then it would cause the degree of a parent node to double.
The only remaining possibility is for this extra degree to go to the interior arcs of $u_{i}$. While this is not a problem, it will cause the interior degrees to increase at a faster rate than we just showed in Lemma 8.3 (Interior Degree Lower Bound), as this would increase the number of interior arcs.
This more rapid reduction in the interior degree leads to container density faster, which would simply lead to a degree doubling node faster. To avoid this, overall regularity is key.
Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph. ∎
We are incrementally refining the properties that must hold for our counterexamples. The two most crucial aspects are the decreasing sequence property and the identification of the minimum degree node. However, additional constraints, such as the overall graph regularity established by Lemma 8.4 (Regular Graph), help simplify future arguments. This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
Example 8.1.
0
$\Rightarrow$ 111122223333444455556666777788889999101010101111111112121212131313131414141415151515161616161717171718181818191919192020202021212121222222222323232324242424H
[…]
Example 8.1 shows some of the patterns that emerge with this regularity in the SSNC. Starting from the outside in, going inward represents the order of containers as they increase in distance from $v_{0}$. This example has visualized only the interior arcs, as the exterior arcs would complicate this visualization. However, just as in previous examples, all nodes in a previous container are mapping to all nodes in the next container.
Some of those patterns that emerge are first the equal out-degree of nodes. For even-sized containers, we are able to see symmetry about balance, uniformity, and a sense of order within each container. Also, while it is not a point of the conjecture, we do see the presence of cycles in each container.
9 Transitive Triangles and Seymour Diamonds
Before we continue with the development of our algorithm and show how it solves the SSNC, we need to address some potential problems and show that they will not hinder our algorithm.
Definition 9.1.
Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, and $u$ form a transitive triangle.
Here, we have $x$ pointing to $y$ and $u$ and $y$ pointing to $u$. The concept of a transitive triangle is an important one because it involves the risk of double counting. The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$. If we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
Definition 9.2.
Let $x\in G$ and $y,w\in N^{+}(x)$. Then $x$, $y$, $w$, and $u$ form a Seymour diamond if $(y,u)$ and $(w,u)\in G$.
The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$. This is a four-vertex-oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow y\rightarrow u$ in this diamond, and there is a path $x\rightarrow w\rightarrow u$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture.
The notion of interior neighbors is another way of describing what [1] referred to as transitive triangles. They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e., one where every node has the decreasing sequence property. We provided a similar result with Lemma 5.4 (Interior Degree at least 1). When a minimum degree node $v_{0}$ has the Decreasing Sequence Property, all of its neighbors have an interior degree of at least 1. That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$, will form a transitive triangle. We can extend that further for any two constructive containers.
xyu
Figure 12: Illustration of a transitive triangle with parent node $x$, child nodes $y$ and $u$, where $u$ is also a child of $y$
Lemma 9.1.
Transitive Triangles from Containers Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). For any two consecutive containers $C_{i}$ and $C_{i+1}$, there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle.
Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm. Transitive triangles represent a common situation with our containers. A parent node $x$ has two children $y$ and $u$ that are also connected by an arc. The purpose of the interior arcs is to ensure that the parent node $x$’s degree does not double. We have already shown through 8.3 (Interior Degrees Lower Bound) that every node in Container $i$ has an interior degree at least $i$.
The two arcs from a parent node and the interior nodes are what create these transitive triangles. What is not acknowledged in lemma 9.1 (Transitive Triangles from Containers) is that every node in the container takes both non-parent roles in different transitive triangles. Every node is both a child of a parent node in the previous container and has a degree of at least one. This is what creates the transitive triangles in every container.
Lemma 9.2.
Seymour Diamonds from Containers
Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). For any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
$u\in C_{i},\quad v,w\in C_{i+1},\quad y\in C_{i+2}$
The following arcs exist: 1. There are arcs $x\to y$ and $y\to u$, 2. There are arcs $x\to w$ and $w\to u$, 3. No back arcs exist between containers.
Similar to Lemma 9.1, Lemma 9.2 (Seymour Diamonds from Containers) lets us know that we have been interacting with Seymour Diamonds all along without knowing it. We didn’t know it because they did not disrupt our algorithm.
This is exactly what we are seeing happen every time there is a Seymour diamond present between three containers. There is a parent node, a choice of at least 2 interior nodes, and an exterior node. Not only is this not a problem; it’s one of the benefits of the GLOVER data structure. Those nodes $y$ and $w$ represent equidistant nodes from $v_{0}$ whose job is to ensure that $x$’s degree does not double. The arcs from $y$ to $u$ and $w$ to $u$ are exterior arcs, connecting these nodes to the next container. The purpose of these arcs is first to meet the degree requirements of each node and second to ensure connectivity. Once each node $y$ and $w$ has connected to the third container, it can take a role as a parent and form its own Seymour diamond.
xywu
Figure 13: Illustration of a Seymour diamond with parent node $x$, child nodes $y$ and $w$, and grandchild node $u$
Lemma 9.3.
Seymour Diamonds and Decreasing Path Algorithm Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP). We are attempting to prevent a degree-doubling node, where we are applying the DSP algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP algorithm.
Proof.
A Seymour diamond consists of two vertices, $x$ and $u$, connected by two parallel paths via $y$ and $w$.
The DSP algorithm can utilize one of these paths while disregarding the other. Both paths connect $x$ and $u$, thus maintaining degree constraints.
The presence of redundant paths does not diminish the outgoing neighbors of any node. This ensures that the minimum degree condition holds throughout the process.
Consequently, these additional connections offer alternative routes without disrupting the algorithm’s operation.
None of these cases are a deterrent to the execution of the DSP algorithm.
Lemma 9.3 (Seymour Diamonds and Decreasing Path Algorithm) shows how we have been proceeding through the GLOVER data structure through our Seymour diamonds. These diamonds are constructed to be presented in our containers. As a result, we simply select an interior node within the container and continue. This is what we saw in our examples in section 4, where Seymour diamonds were also present, for example 4.3. This is also what we saw in our social media example 5.5 in section 5. This is also what we saw in our regular graph example in 8.1 in section 8.
Lemma 9.4.
Interaction of Minimum Degree Nodes and Transitive Triangles
Suppose $G$ is an oriented graph represented by a GLOVER. Let $v_{0}$ be a minimum degree node with degree $\delta$. Consider a transitive triangle formed by nodes $u$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(u)=0$.
Then the following cases are possible:
  • $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$
  • $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$
  • $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$
  • $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$:
Proof.
  • Case 1: $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$: Both nodes $y$ and $u$ out-degree $\delta$. The triangle has the arc $(y,u)$. Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$. This means that $u$ has $\delta$ out-neighbors that are not in the triangle. Since we assumed that $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$, we know that $y$ cannot connect to all not $u$’s neighbors since $y$ has a neighbor in the triangle ($u$). This node that is $u$’s neighbor, but not $y$’s neighbor, is a candidate to extend the path.
  • Case 2: $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree $\delta$. The node $u$ has out-degree greater than $\delta$. Again, the triangle has the arc $(y,u)$. Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$. This means that $u$ has more than $\delta$ out-neighbors that are not in the triangle. We assumed that $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$. We know that $y$ cannot connect to all not $u$’s neighbors since $d^{+}(u)>d^{+}(y)$. We can then select one of these nodes that is a neighbor of $u$, but not a neighbor of $y$ to be a candidate to extend the path.
  • Case 3: If $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$: The node $y$ has out-degree greater than $\delta$. The node $u$ has out-degree $\delta$. The triangle has the arc $(y,u)$. Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$. The node $y$ has more than $\delta$ out-neighbors that are not in the triangle. The node $u$ has exactly $\delta$ out-neighbors that are not in the triangle. The maximum number of neighbors that $y$ and $u$ can share is $\delta$, that being $u$’s out-degree. Then we can select one of the nodes that is a neighbor of $y$, but not a neighbor of $u$ to be a candidate to extend the path.
  • Case 4: If $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree greater than $\delta$. The node $u$ has out-degree greater than $\delta$. We have a few cases because we have to remember that in this data structure the purpose of our containers and the interior nodes was to prevent parent nodes from having their degrees double. The case where both nodes in a transitive triangle have more than $\delta$ out-neighbors means that $x$ has up to $2*\delta$ out-neighbors, which would cause its degree to double. Thus, to prevent $x$’s degree from doubling, there must be some overlap between the out-neighbors of $y$ and the out-neighbors of $u$. These two sets cannot be equal to one another, though. The most overlap we can have is $N^{+}(u)\subsetneq N^{+}(y)$ or $N^{+}(y)\subsetneq N^{+}(u)$. Both of these situations leave room for a node in the larger set that is not in the smaller set as a candidate to extend the path. As we allow for less overlap, the number of candidate nodes to extend the path grows. We can see that there will always be a node in the larger set to extend the path.
Now that we have a better understanding of transitive triangles and Seymour diamonds and how they interact with our data structure. We can see that they are not a problem for the development of our algorithm. We will continue our investigation of oriented graphs. Next, we will blend everything together into our main results.
10 Main Theorem: Seymour’s Conjecture is True
10.1 Introduction
Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some results.
So far, this battle has been a battle between a catalyst, the chosen minimum degree node, and the bottleneck, the Decreasing Sequence Property (DSP). We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
There is a limit, though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
We first restate the Lemma 6.5 (Non-Existence of Back-Arcs), from section 6, but slightly rephrased.
Lemma 10.1.
Let $G$ be an oriented graph represented by a GLOVER, with all nodes satisfying the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node in $G$. We are attempting to prevent a degree-doubling node. If there are back arcs in $G$, then a node’s degree will double.
10.2 Theorem Statement
Degree Doubling by Exterior Decreasing
Theorem 10.1.
Let $G$ be an oriented graph represented by a GLOVER with all nodes having the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node where we set up our partition. Then Seymour’s conjecture holds true.
Proof.
We have set up the containers using a GLOVER data structure. There are no back arcs; else we will have a degree-doubling node. Next is lemma 7.2 (Interior Degree Doubling), which states that in every container we will have a node whose interior degree doubles.
These lemmas are what allowed us to set up and use lemma 5.6 (Weak Decreasing Exteriors Lemma). That lemma gives us a decreasing set of positive numbers. This set is bound to reach the size of two. When it does, one of the nodes in that set will have all of its arcs to the next container (thereby being exterior arcs). This will cause a parent of that node’s degree to double.
Algorithm 1 Algorithm for Analyzing Degree Doubling in Containers
1:Initialization:
2:Choose the minimum degree node $v_{0}$.
3:Divide nodes into containers $C_{1},C_{2},\ldots$ based on their degree and proximity to $v_{0}$.
4:for each container $C_{i}$ in $G$ do:
5:     Locate a node $x_{i}$ in $C_{i}$ whose interior degree doubles.
6:     Move to a neighboring container $C_{i+1}$.
7:     if there is a back arc from $C_{i+1}$ to a prior container then
8:         Locate child of $x_{i}$ whose exterior degree doubles, causing $x_{i}$’s overall degree to double.
9:         Halt.
10:     else
11:         Since there is no back arc, the container $C_{i+1}$ will be limited in size to be 1 less than the size of container $C_{i}$.
12:         Repeat from Step 4.
13:     end if
14:end for
Start with graph size $n$Select a minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots C_{k}$dentify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to a prior containerLocate the child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
Figure 14: Flowchart depicting the degree-doubling algorithm applied within containers, demonstrating the process of identifying nodes with the Decreasing Sequence Property.
Corollary 10.1.
Degree Doubling by Container Density In an oriented graph $G$ with a minimum degree node $v_{0}$ and containers $C_{1},\allowbreak C_{2},\allowbreak\ldots\allowbreak C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
There are two possibilities:
  • Case 1: The containers continue to shrink in size. Simultaneously, the interior degrees continue to grow. We reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
  • Case 2: For small minimum degree nodes, the container’s shrinkage is not fast enough to cause this collision.
Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and an increased number of nodes. This was represented by the yin and yang effects early on. We see it come full circle here. Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbors represented the size of the containers.
In small graphs like 4.3 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur. To understand the reasoning behind this, one must understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1. The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$. So, for small graphs, we are just not going out far enough to catch up. For larger graphs, though, we do start to see this.
Corollary 10.2.
Occurrence of Last Dense Container Let $G$ be an oriented graph represented by a GLOVER with minimum degree $k$. Then there will be a container that reaches maximal density. Further, if that container is $C_{i}$, it will be $i$-out-regular.
Proof.
We begin by stating lemma 6.1 (Regular Interior Degrees). While these are called interior degrees, these are the out-degrees of the nodes in a container. This lemma indicates that every container must have regular out-degrees.
The second part of this proof is lemma 8.2 (Strong Exterior Decreasing), which says that in container $i$, the size of the exterior is at most $d^{+}(v_{0})-i$. Lemma 8.3 (Interior Degree Lower Bound) tells us that each node must have degree $i$ in this container.
Thus, we have a series of regular containers that are getting smaller as we get further from the minimum degree none. At the same time, these containers are expected to hold fewer nodes of higher degree, i.e., a graph of more arcs. This is progressively getting to a denser container. As described in corollary 10.1 (Degree Doubling by Container Density), this cannot continue forever. There must be a last, most dense container. ∎
The SSNC is made more aesthetic by the appearance of these orientations of out-regular graphs as containers get smaller. Not only has the conjecture been proven true, but with an inherent symmetry and order. There is a visual pattern emerging inside these containers that is not just important to graph theorists but also possesses a structure that is easily appealing to non-mathematicians.
Corollary 10.3.
Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $\delta$, and let the most dense graph be in container $C_{i}$. Then every node in container $C_{i}$ has its degree doubled in the representation.
The discovery of multiple degree-doubling nodes elevates the SSNC from a mere question of existence to one in abundance. There were many who doubted that one even existed. Now there are many.
Second is the way we find it. It’s the presence of a last dense container with an oriented regular graph as these containers decrease in size. We see this occur as the interior degrees are forced to increase as the distance from $v_{0}$ is increased. These interior degrees are set up to prevent a parent node’s degree from doubling. Ultimately though, they are unable to do that, as we see the collision of the container shrinkage and interior degree requirements.
Then third we can look at the applications. We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group. Dense regular graphs have many applications in themselves, from areas like expander graphs to network topology design to coding theory.
Theorem 10.2.
Suppose we have a graph $G$. The Decreasing Sequence Property Algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ represents the number of vertices in $G$ and $|E|$ represents the number of edges in $G$.
Proof.
We will first show that the complexity of building the partitions has a worst-case complexity of $O(|V|+|E|)$.
Given a graph $G$, the search for a minimum degree node that’s $O(|V|)$ nodes. If a priority queue is used, this can be optimized further. Then, to evaluate the degrees of each of those nodes is a constant factor. This makes the search for the minimum degree node $O(|V|)$.
Once we have this value, we will place every node and every edge of $G$ into a pre-processing array ready to be partitioned. Each edge is examined, contributing $O(|E|)$ complexity, while nodes contribute $O(|V|)$. Each node will individually be processed based on its distance from the minimum degree node, along with the connected outgoing arcs. This will require $O(|V|)$ comparisons and $O(|V|+|E|)$ stores into containers. This means that the partitioning has a complexity of $O(|V|+|E|)$.
Next, we will look into the complexity of the run-time analysis of the Decreasing Sequence Algorithm.
First, remember that we proved in lemma 7.2 (Interior Degree Doubles) that every container has a node whose interior degree doubles. Secondly, we also note that the lemma 6.1 (Regular Interior Degrees) implies that all the nodes are regular inside a container. Third, lemma 7.1 (Indistinguishable Regular Interior Nodes) lets us know that nodes within a container are indistinguishable from one another. The combination of these three lemmas lets us know that every node inside a container has their interior degree doubled. This gives the complexity of finding a node whose interior degree has doubled $O(1)$.
What remains is to consider the complexity necessary to combine these interior degree nodes to solve the overall problem. Solving the full problem requires determining the number of containers necessary for this partition, which is bounded above by $O(|V|)$. The total number of containers cannot exceed the number of nodes, $|V|$, since containers are based on partitions. Processing each container’s exterior arcs and remaining possible nodes is bounded by $O(|V|+|E|)$. This gives the run time a complexity of $O(|V|+|E|)$. ∎
This complexity analysis shows that the Decreasing Sequence Property Algorithm can effectively find a degree-doubling node in polynomial time. Moreover, this will work for any oriented graph. Complexity analysis shows that it is not hindered by larger graphs.
A further point is to consider the concept in theorem 10.2 of all nodes inside a container having their degrees double. It was not spoken about in the paper because the focus was entirely set on proving the SSNC, but this is another example of multiple degree-doubling nodes, just a localized instance.
We opened by showing several applications of the SSNC, including social media, social network modeling and network analysis. This highlights the necessity for an algorithm that can work efficiently in practice. This proof that the algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ is the number of vertices and $|E|$ is the number of edges in the graph is important. This algorithm’s linear time complexity shows that even enormous graphs may be processed effectively. That makes it potentially applicable to real-world networks with millions or billions of nodes and edges. This suggests that the algorithm is not only theoretically sound but also practically feasible for implementation and use in real-world scenarios, such as analyzing social networks, identifying influential nodes in communication networks, or understanding the spread of information.
11 Conclusion and Future Work
In this work, we explored the Seymour Second Neighborhood Conjecture (SSNC) in oriented graphs. We did so through the lenses of graph theory, combinatorics, algorithms, and data structures. This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide-and-conquer paradigm.
One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph. This effectively solves the SSNC with an answer in the affirmative. The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree-doubling node was in the graph. Instead, the algorithm selects a minimum degree node and can partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
This approach introduced a new perspective on the SSNC. It was no longer framed as merely a graph theory problem. By utilizing the divide-and-conquer paradigm and solving different parts independently, this algorithm shows that the SSNC could effectively be treated with computer science techniques. This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists. Alternatively, it may be that there is not enough interdisciplinary communication about these problems and the potential techniques to solve them.
The SSNC started as a question in 1990. Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers. It has been talked about in books. on message boards, in classrooms, and at conferences. This problem has a long history.
With a solution now in hand, it’s time to begin looking forward. The SSNC can be represented as a social media problem instead of an oriented graph. Here, the nodes are people, where no one follows back. Now instead of wondering if there is a person whose sphere of influence doubles, we would simply be searching for that person. We just need to implement the algorithm to find the person. We could have the graph represent epidemiology. Here the nodes will be infected people, and the arcs would be one-way infections. The question of interest would be, can we implement the algorithm to find the person who is a super-spreader and do so quickly? There are many more real-life examples where we can apply this problem and this algorithm. There are fields like security, telecommunications, urban planning, and neural networks. There are too many to name in this short paper.
We have presented a linear time algorithm that can work on large graphs in these problems. By constructing the algorithm for the SSNC and proving that the algorithm will find a degree-doubling node, we have proved that the algorithm is theoretically sound.
Secondly, we introduced a data structure, Graph Level Order-GLOVER. This data structure allowed us to get a lot done on the conjecture. Many of the things inherent to SSNC made it easier for GLOVER—the well ordering and the dual metrics of distance and degrees.
Future research and development should explore questions of complexity and scalability. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
Finally, this research offered an alternative method for the SSNC. We focused on identifying a node whose out-degree doubles in the graph’s square rather than just showing that it exists. To do this, we developed an effective algorithm and proved its complexity. Other problems in graph theory and related domains may benefit from this change in emphasis from existential proofs to algorithmic solutions. This may encourages researchers to explore constructive approaches with a focus on practical applications. Hopefully, this work will inspire further research on constructive approaches to other graph theory problems. Also, we encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
12 Acknowledgments
I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:6 has been a guiding light throughout this journey. I am profoundly grateful to my family—my mother, father, and wife—for their unwavering support and encouragement.
Special thanks to the late Dr. Nate Dean for his contributions to this work and to Dr. Michael Ball for his support. I want to sincerely thank everyone who has taken the time to pray for me, share encouraging words, and offer constructive feedback. Your support has meant so much to me, and I’m truly grateful for the kindness and thoughtfulness you’ve shown.
Finally, I extend heartfelt gratitude to my family and friends for their patience and belief in me throughout this journey.
The GLOVER data structure is protected by a provisional patent application filed with the U.S. Patent and Trademark Office.
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Ranked Structure: The elements of $S$ are partitioned into a ranking of ”containers” or ”levels,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Definition 13.4.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the interior neighbors of $y$ with respect to $x$, denoted as $int(x,y)$, are the out-neighbors of $y$ that are also out-neighbors of $x$. Formally,
$int(x,y)=N^{+}(x)\cap N^{+}(y),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$ in the graph. The number of interior neighbors of $y$ with respect to $x$ is called the interior degree of $y$.
Definition 13.5.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the exterior neighbors of $y$ with respect to $x$, denoted as $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$. Formally,
$ext(x,y)=N^{+}(y)\cap N^{++}(x),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$, and $N^{++}(x)$ denotes the set of out-neighbors of the out-neighbors of $x$ (i.e., the second neighbors of $x$). The number of exterior neighbors of $y$ with respect to $x$ is called the exterior degree of $y$.
Definition 13.6.
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$. Definition: Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
Definition 13.7.
Definition: Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
14 Appendix 2: Applicability of the Graph Level Order
The deciding factor that held the algorithm together and helped solve the SSNC was the Graph Level Order (GLOVER) data structure. This research did not begin as an investigation into data structures. Instead, a series of operations were conducted on the oriented graphs—partitioning, ordering those partitions, and adding interior and exterior arcs—until what was left was an ordering of both the partition and nodes within the partition. Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric. Unlike the SSNC, exterior arcs do not always have to only be defined to the next container. That was a problem-specific definition.
Moving forward, the GLOVER data structure should be well suited to tackle problems in graph theory and beyond. By introducing containers, another way of doing induction on graphs has been introduced. The concept of divide-and-conquer, which has dominated much of computer science, shows up in this paper as well. We can treat these containers independently of one another and solve their problems locally before bringing them back to the global problem. This shows how versatile the GLOVER data structure is moving forward.
What makes GLOVER special are things like total order on containers. Now there is a two-way ranking of nodes, as opposed to traditional one-way rankings like lexicographical sorting. This agrees with many real-world systems where we have dual metrics that are often competing, like price and performance. This is hard to measure on a single scale.
Next, these nodes are not just treated like clusters inside these containers but like nodes in graph theory. The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth-first search, or centrality to be called. What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
This is still a new data structure, so we are learning a lot about it. There are many applications that have not been tried yet. However, since it brings the strengths of a total order and a graph into one unified structure, it should be able to model complex relationships. This should include domains of machine learning like clustering, dynamic scheduling, resource management, project management, and natural language processing.
15 Appendix 3: Visualization and Accessibility of Oriented Graphs
When we take a course on graph theory, oriented graphs are probably not on the course load. Second neighbors or squares of graphs are probably not either. To fully grasp this problem, though, we need to be able to delve into these definitions through visualizations.
For example, research on the SSNC led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs. The paper addressed how these terms were intertwined with each other. An example is where the paper says that there is a monotonically decreasing sequence.
James Robert Brown’s book ”Philosophy of Mathematics” speaks of the importance of visualizations in mathematics. Proofs are still the foundation, but sometimes visualizations can serve as inspiration or a spark for those proofs.
This realization motivated the creation of a complimentary website. The site serves as an interactive tool to understand the research. It also has things like MathJax installed to read lemmas, definitions, and proofs from this paper. Where possible, things like HTML canvas and JavaScript D3.js are used to visualize these concepts and help improve understanding.
Readers are encouraged to explore the website for additional information and examples at https://glovermethod.com. This is an evolving project. Additional examples and applications will continue to be added. Readers are invited to explore the website. Please provide feedback on any aspects of the site or content for improvement. Our goal here was to help users connect with the SSNC. Hopefully, this tool not only serves as support for this paper but also as an educational resource in the fields of graph theory and computer science. Thank you for your engagement.